### v1

# How LLMs Learn to Reason: A Complex Network Perspective

Abstract
Training large language models with Reinforcement Learning from Verifiable Rewards (RLVR) exhibits a set of distinctive and puzzling behaviors that remain poorly understood, including a two-stage learning curve, V-shaped response-length trajectories, and a pronounced vulnerability to catastrophic forgetting. In this work, we propose that these seemingly disparate phenomena can be explained using a single unifying theory: the model’s reasoning process maps to the self-organization of a semantic complex network whose topology remains persistently sparse, with the average degree pinned close to two. This topology imposes a fundamental mechanism for forgetting and learning: it first drives the system into a maximally frustrated state where “skill islands” form, slow-learning happens, and forgetting is induced; then it enters a sharp growth phase where the new skills are “bolted on”, driven by phase-transition-like learning at the web’s frontier. Equipped with the theory, we propose Annealed-RLVR, a principled algorithm that introduces an SFT-based “heating” step at the point of maximal frustration to resolve the competitive bottleneck and enhance the reasoning capability of the model. Experiments on a 1.5B-parameter model demonstrate that the approach outperforms standard RLVR on both in-distribution and out-of-distribution benchmarks. By recasting RLVR from black-box optimization into a predictable process of structural self-organization, our work provides a new physical intuition for engineering the emergent reasoning capabilities of future AI systems.
1 Introduction
Large language models (LLMs) trained with Reinforcement Learning from Verifiable Rewards (RLVR) (Liu et al., 2025; Guo et al., 2025; Team et al., 2025; Yang et al., 2025) are increasingly capable of multi-step, “System-2” reasoning. Yet, this training paradigm consistently exhibits a set of puzzling macroscopic behaviors across diverse architectures and benchmarks. These include (i) rapid initial gains that give way to a prolonged plateau; (ii) a V-shaped trajectory in the length of correct solutions, which first shortens before steadily lengthening (Luo et al., 2025; He et al., 2025; Fatemi et al., 2025); and (iii) a pronounced vulnerability to catastrophic forgetting when interleaved with supervised fine-tuning (SFT) (Li & Hoiem, 2017; Fernando et al., 2025; Ding & Wang, 2025). Although these phenomena are widely observed, the reasons for their co-occurrence remain elusive. A clear understanding of the underlying mechanism would be essential for developing more efficient approaches to training reasoning-capable models.
A promising theoretical framework analyzes these phenomena by modeling the LLM’s reasoning process on an implicit reasoning graph — a network of semantic states and logical transitions(Wang et al., 2024; Dutta et al., 2024; Xiang et al., 2025; Cabannes et al., 2024; Minegishi et al., 2025). However, this perspective faces a critical roadblock: a direct, microscopic analysis is currently intractable. Constructing this entire graph from a model’s high-dimensional latent space remains a formidable open problem, preventing a direct investigation into the structural origins of the observed dynamics.
Here, we argue that the very universality of the RLVR phenomena provides a path forward. In complex systems, behaviors that are independent of microscopic details are often emergent properties, governed by the system’s large-scale organization (Anderson, 1972; Watts & Strogatz, 1998; Barabási & Albert, 1999). This insight motivates an approach rooted in the core logic of the Renormalization Group (RG)(Wilson & Kogut, 1974): instead of getting bogged down by intractable local details, one can leap directly to hypothesizing a core principle about the system’s large-scale structure to explain its macroscopic behavior. Applying this strategy, we shift our focus from the full reasoning graph to its coarse-grained backbone, which emerges by abstracting away fine-grained skills encoded in local network patterns like motifs or community structures. This leads us to our central hypothesis regarding the emergent structure that forms under RLVR training for general-purpose System-2 reasoning:
Hypothesis 1.
The concept web, defined as the coarse-grained reasoning graph, is a sparse network whose effective average degree is pinned to $\left<k\right>\approx 2$.
The implications of this hypothesis are non-trivial: an average degree of $2$ implies a predominantly tree-like structure, which is efficient for generalization but inherently fragile. This topology can still support rare, large-scale cycles critical for complex operations like self-correction (as detailed in Appendix C), without significantly altering the global average degree.
(a) CoNet
(b) LLM
Figure 1: CoNet Faithfully Reproduces Core LLM Training Dynamics. The minimal CoNet model (a) remarkably reproduces the two core empirical signatures of RLVR training observed in the DeepScaleR-1.5B LLM (b). These signatures are: (i) a two-stage reward dynamic consisting of a fast-learning phase followed by a slow-learning plateau (top panels), and (ii) a non-monotonic, V-shaped evolution of the correct response length (bottom panels). This striking correspondence validates CoNet as a tractable minimal model for our theoretical analysis. The inset in (a) schematically depicts the CoNet abstraction. Lines in (b) are smoothed with a 60-step moving average.
To validate this hypothesis, we combine two approaches: a theoretical study using the Concept Network Model (CoNet) (Cai et al., 2025), a minimal “computational microscope” that reproduces the LLM’s core training dynamics, and numerical experiments on a 1.5B-parameter LLM(Luo et al., 2025). This dual approach reveals a clear structural narrative for the observed puzzles. The learning plateau is shown to be the arduous consolidation of disconnected “skill islands”, a transition that directly explains the V-shaped length trajectory. Furthermore, the resulting sparse, tree-like concept web accounts for the model’s fragility and its vulnerability to catastrophic forgetting with SFT.
This sparse topology imposes a microscopic dynamic of learning and forgetting that is a hallmark of the RLVR-driven self-organization process. We find the degree-2 constraint first forces the system into a maximally frustrated state at the onset of the slow-learning plateau, a phase dominated by frustration-induced forgetting as skills are competitively suppressed. Subsequently, the very same structure enables the growth mechanism that governs the entire slow-learning phase: a steady, sequential process where new skills are individually “bolted on” at the web’s expanding frontier via a phase-transition-like learning dynamic.
Equipped with the sparse-web theory, we propose Annealed-RLVR, a principled algorithm that introduces an SFT-based “heating” step at the point of maximal frustration to resolve the competitive bottleneck and enhance exploration and reasoning efficacy of the subsequent RLVR phase. Experiments on both CoNet and the 1.5B-parameter LLM confirm the efficacy of this strategy: Annealed-RLVR surpasses standard RLVR on both in-distribution and out-of-distribution benchmarks, transforming our physical intuition into a practical tool for engineering emergent reasoning.
(a) Step 20
(b) Step 50
(c) Step 800
Figure 2: Visualizing the Structural Evolution from Skill Islands to the Concept Web. These network snapshots provide a direct, qualitative narrative of the concept web’s formation, corresponding to the different phases of training. (a) In the early fast-learning phase (Step 20), the model discovers a few short, disjointed reasoning paths, representing the first nascent “skill islands”. (b) At the onset of slow learning (Step 50), the system has proliferated into a maximal collection of disconnected islands. (c) Deep in the slow-learning phase (Step 800), these previously separate islands have coalesced into a single, giant connected component, forming the unified and expansive concept web.
2 The V-Shaped Signature as Evidence for a Sparse Concept Web
The V-shaped evolution of response length is a well-demonstrated, quantifiable hallmark of deep structural reorganization during RLVR training(He et al., 2025; Luo et al., 2025). To anchor our investigation, we first reproduce this signature by applying RLVR to the DeepSeek-R1-Distill-Qwen-1.5B model following the DeepScaleR protocol (Luo et al., 2025). As shown in Fig. 1, our experiment recovers the characteristic V-shaped curve (Fig. 1). We contend that this is not a mere by-product but the primary empirical evidence for our central claim: the emergence of a sparse, unified concept web.
To dissect the mechanism underlying this signature—a task intractable within the LLM itself—we deploy an extended, multi-task variant of the CoNet as a “computational microscope” (See Appendix D for details). This minimal model reduces the learning process to a tractable graph-traversal problem and, remarkably, reproduces the V-shaped curve observed in our LLM experiment (Fig. 1), validating it as a faithful proxy for uncovering the underlying mechanism.
Armed with this confirmation, we now expose the topological origin of the V-shape. The V-shaped curve arises from a distinct two-stage process. During the initial fast-learning phase where response length decreases, the model undergoes local optimization: it rapidly discovers short, efficient solutions to individual problems, giving rise to a proliferation of disconnected “skill islands” (as shown in Fig. 2(a)). This burst is quantified by the number of distinct solution clusters, which peaks right before the onset of slow learning (Fig. 2(b)). Stage one thus constitutes a parallel search for local efficiency that naturally drives down the average response length.
(a) Correct Response Histogram
(b) Step 50
(c) Step 800
Figure 3: A Sparse Web Structure Necessitates Longer Reasoning Chains. This figure links the emergent sparse topology of the concept web to the observed increase in response length. (b, c) The largest connected component (edges with probability > 0.95) remains sparse (average degree $\approx 2$) even as it grows from step 50 to 800. (a) Consequently, the distribution of solution path lengths shifts decisively to the right during this period, confirming that navigating the sparse backbone requires longer reasoning chains.
The subsequent slow-learning phase, where response length steadily increases, marks a fundamental shift to global integration. The primary task is no longer discovering new islands but weaving them into a single, expansive concept web (see Fig. 2(c)). As the largest connected component relentlessly grows, a critical structural property emerges: the web remains persistently sparse, with its average degree stabilizing near two (as shown in Fig. 3(c)). This provides direct, quantitative validation for our central Hypothesis (Hypothesis 1).
This emergent sparsity is the topological key to the V-shape’s rising slope. A network with such a low average degree is dominated by chain and tree-like structures, inherently lacking the shortcuts and redundant pathways found in denser graphs. Consequently, connecting two previously distant concepts requires traversing long, intermediate paths. The histograms in Fig. 3(a) provide dispositive proof of this effect: As the sparse web expands, the distribution of solution lengths undergoes a decisive and sustained rightward shift, confirming that navigation on this sparse backbone necessitates longer reasoning chains. Thus, the V-shaped curve furnishes a robust macroscopic signature of an underlying topological transition: from parallel, local optimization on disconnected components to serial, global navigation across an expanding, sparse backbone.
Yet can this simple, largely tree-like model capture the full complexity of LLM reasoning? After all, an LLM’s autoregressive memory enables sophisticated behaviors such as self-correction or reflection (e.g., “Wait, let me rethink that…”)—that would manifest as cycles in the reasoning graph. We argue they are fully compatible(Minegishi et al., 2025). These reasoning cycles are not the dense, ubiquitous shortcuts of a highly connected graph. Instead, they are likely the cognitive manifestation of sparse, local loops—occasional, high-cost revisions on a much larger, fundamentally chain-like scaffold. The existence of a limited number of such cycles does not significantly alter the global average degree of the web, which remains close to two. With the sparse web model established, this link between a macroscopic dynamic and a microscopic topology forms the foundation for the analysis that follows.
(a) CoNet
(b) LLM
(c) Step 400 (Before SFT)
(d) Step 400 (After SFT)
(e) Step 450 (Continued RLVR)
Figure 4: The Fragility of a Sparse Web: Catastrophic Forgetting and Fast Recovery. This figure demonstrates a key prediction of our sparse web hypothesis: that the concept web is fragile, relying on critical bridge-like connections. (c-e) The microscopic view in CoNet shows how a lightweight supervised fine-tuning (SFT) on a converged web (c) severs these critical bridges, causing the structure to fragment (d). Subsequent RLVR rapidly repairs these links (e). (a, b) This microscopic severing manifests as macroscopic catastrophic forgetting in both CoNet and the 1.5B LLM, where performance collapses upon initiating SFT. The subsequent fast recovery once RLVR resumes highlights the localized nature of the damage, underscoring the web’s structural fragility.
3 The Double-Edged Sword of Sparsity: from catastrophic Forgetting to Simulated Annealing scheme
From the graph theoretic point of view, the sparsity in the conceptual network implies low redundancy, fragility, and critical vulnerability. It also reflects a structural hypothesis for a well-known operational challenge: why models that have formed a concept web via RLVR are so susceptible to catastrophic forgetting when subsequently fine-tuned with (Li et al., 2024). Our theory posits that this is not a widespread erasure of knowledge, but a specific topological failure: the surgical severing of critical, bridge-like connections.
To test this mechanistic hypothesis, we subjected models in the deep slow-learning phase to a lightweight in-distribution SFT (see Appendix H and I). As expected, the intervention triggered a precipitous collapse in overall performance in both CoNet and the 1.5B LLM (Fig. 44(a), 4(b)). Having reproduced this known macroscopic phenomenon, we leveraged CoNet’s transparency to inspect the microscopic aftermath.
The result is a direct validation of our structural explanation. The microscopic view reveals that the SFT catastrophic forgetting was not a global degradation of the web. Instead, SFT’s aggressive optimization of a few paths overwrote the policy at critical branching nodes, surgically severing the single connection points for vast downstream subgraphs. Fig. 44(c)4(d) provides a stark visualization of this: a unified web is instantly fragmented, offering a clear, structural cause for the catastrophic performance drop.
However, this is where the sword’s second edge is revealed. The damage, while profound in its effect, is topologically localized. SFT does not erase the knowledge encoded in the now-disconnected subgraphs; it merely renders them unreachable. This localization explains the remarkable speed of recovery, which can be observed in Fig. 44(a)4(b). When RLVR training resumes, it triggers a quick rebound because the task is simply to “re-solder” a few broken connections to restore access to the latent scaffold as shown in Fig. 44(e).
This two-step process of localized disruption followed by rapid repair is strongly analogous to simulated annealing Kirkpatrick et al. (1983). The brief SFT phase acts as a high-temperature “heating” shock, breaking connections and allowing the system to escape a potentially suboptimal state. The subsequent RLVR phase then acts as a controlled “cooling”, guiding the system as it re-forms connections, ideally settling into a more robust configuration (Lai et al., 2025). This insight—that a known vulnerability can be repurposed as a deliberate annealing step—forms the foundation for the Annealed-RLVR algorithm, which we introduce in Sec.5.
4 Microscopic Mechanisms of RLVR: Forgetting by Frustration, Learning by Phase Transition
Our sparse web theory reveals that the smooth, aggregate reward curve of RLVR training masks a dramatic microscopic behavior. Individual problem trajectories in Fig. 5 reveal a fundamental tension born from the network’s sparse topology: a subtractive process of frustration-induced forgetting, peaking at the onset of slow learning, and an opposing, additive process of phase-transition-like learning that drives knowledge integration throughout the slow-learning stage.
(a) CoNet
(b) LLM
Figure 5: Microscopic Mechanisms of RLVR: Forgetting by Frustration, Learning by Phase Transition The learning trajectories for individual problems in CoNet (a) and the 1.5B LLM (b) reveal a fundamental duality. (i) Frustration-Induced Forgetting: At the transition to slow learning (see inset in (a)), the intense competition for connections on a sparse web manifests as volatile, non-monotonic accuracy curves, where some skills are competitively suppressed. (ii) Phase-Transition-Like Learning: Subsequently, the web’s sparse frontier enables new skills to be acquired in punctuated, accelerated jumps (orange and green). The smoother gradients in the LLM (vs. the clean steps in CoNet) can be attributed to a classic finite-size effect, as explained in the main text.
4.1 Frustration-Induced Forgetting: The Price of Integration
The first dynamic is driven by frustration-induced forgetting: a competitive process where the advancement of some skills forces the deterministic regression of others as the model integrates its “skill islands”. This process manifests as volatile, non-monotonic learning curves in the learning dynamics (Fig. 5) and represents an internal “push-and-pull” fundamentally different from the external shock of SFT-induced catastrophic forgetting (Shenfeld et al., 2025; Jin et al., 2025).
This process peaks in the maximally frustrated state at the onset of the slow-learning stage, a critical juncture where skill islands compete for limited connections under the $\left<k\right>\approx 2$ constraint. Paradoxically, this state maximizes exploration: the model’s ability to find diverse, correct out-of-distribution solutions peaks, far exceeding the base model or the same model after prolonged training (Fig. 6(b), blue). This subsequent degradation in performance is consistent with the phenomenon of policy collapse, which has been recently reported in several studies to occur after extended RLVR, leading to a loss of exploratory power (Chen et al., 2025; Yue et al., 2025; Cui et al., 2025).
4.2 Phase-Transition-Like Learning: The Engine of Growth
The opposing dynamic, phase-transition-like learning, is an important engine of new knowledge acquisition (Power et al., 2022; Lin et al., 2025). The evidence for this is clearly shown in CoNet’s sharp, step-like accuracy jumps (Fig. 5(a)), and the steep accelerated learning gradients observed in the 1.5B LLM (Fig. 5(b), more details in Appendix G). This phenomenon is a multi-task extension of the "Learning at Criticality" (LaC) mechanism, first identified in a single-task context (Cai et al., 2025). The LaC framework models this process as a continuous phase transition(Stanley & Ahlers, 1972; Wilson, 1983) where the policy collapses from exploring diverse reasoning paths to exploiting a single optimal one. Crucially, analysis revealed that near this critical point, the system’s exploratory power is maximized, and the reasoning path lengths become scale-invariant—a hallmark of continuous phase transitions. This context explains the smoother gradient in the LLM. Compared to the step-like jumps in CoNet, the LLM’s learning curve is a classic manifestation of a finite-size effect(Fisher & Barber, 1972): it has a more constrained set of distinct reasoning paths for any given task, which rounds out the idealized sharp transition into a steep but continuous curve.
The puzzle is why this delicate phenomenon survives in a chaotic multi-task setting. Our sparse-web model explains this through topological isolation: a node on the expanding frontier of a $\left<k\right>\approx 2$ network is statistically isolated, decoupling it from system-wide competition. Shielded from interference, a new skill can be efficiently “bolted on” via a clean phase transition. This perspective also provides a microscopic explanation for aggregate policy collapse in the standard RLVR (Chen et al., 2025; Yue et al., 2025; Cui et al., 2025). We hypothesize this is not a monolithic failure but the statistical sum of asynchronous, microscopic collapses at the frontier, a hypothesis supported by the LaC finding that training just beyond criticality risks overfitting-induced collapse.
5 Annealed-RLVR: An Algorithm Born from the Sparse Web Theory
(a) Randomly Selected 512 Training Problems
(b) Minerva
Figure 6: Annealed-RLVR Outperforms Standard RLVR. This figure compares our Annealed-RLVR (“Annealed”) with a standard RLVR baseline (“RLVR”) on (a) in-distribution problems and (b) the out-of-distribution Minerva dataset. (Right panels) The best@k accuracy curves show that Annealed-RLVR achieves superior performance in both settings, confirming improved generalization. (Left panels) The histograms reveal the underlying mechanism: the annealing intervention systematically reduces the population of unsolved problems (low-accuracy bins) and improves the mastered ones (high-accuracy bins).
Our sparse-web theory reveals a core narrative of RLVR training: the system drives itself into a maximally frustrated state that is both a perilous bottleneck and a moment of peak exploratory potential (Fig. 6(b), blue curve). This insight motivates Annealed-RLVR, a theory-driven intervention that leverages the high-entropy frustrated state to guide the model toward a more globally optimal configuration (Cheng et al., 2025; Chen et al., 2025).
The algorithm operationalizes this concept with a precisely timed intervention, triggered by macroscopic signatures of the maximally frustrated state. At this juncture, a brief SFT “heating” phase begins. The timing is critical: because the “skill islands” have not yet solidified into a single web, the system is uniquely robust to a targeted shock that would otherwise cause catastrophic forgetting—a theoretical claim confirmed by our experiments (Fig. 4). The SFT thus functions not as a disruptive overwrite but as a corrective jolt to break the policy’s premature commitment to suboptimal routes. We implement this by surgically applying SFT only to problems with low accuracy (e.g., < 0.1) for which correct solutions exist. Immediately following this targeted heating, the “cooling” stage resumes standard RLVR, allowing the now more pliable policy to re-settle into a more robust and integrated final state.
The results validate this theory-driven approach (Fig. 6). Best@k accuracy curves demonstrate that Annealed-RLVR outperforms standard RLVR on both in-distribution problems and, crucially, the out-of-distribution Minerva dataset, confirming the intervention fosters true generalization. The histograms reveal the underlying mechanism: the SFT “heating” provides a crucial performance boost to the most difficult problems, lifting their success rates from near-zero and making them “visible” to subsequent RLVR exploration. The RLVR “cooling” phase then efficiently guides these newly accessible problems toward mastery. This results in a systematic reduction of unsolved problems and a corresponding population shift towards solved ones, confirming that our intervention guides the model past the inherent bottlenecks of standard RLVR to achieve superior reasoning. More details are given in Appendix H and I.
6 Conclusion and Outlook
We have introduced a sparse-concept-net theory to explain the puzzling behaviors of the RLVR training, including the two-stage learning curve, V-shaped response curves, and vulnerability to catastrophic forgetting. Our theory proposes that these diverse phenomena are consequences of one underlying process: the formation of a sparse concept web, with the averaged degree pinned close to two. This physical model directly explains the V-shaped curve as a shift from local optimization to global integration and recasts catastrophic forgetting as the severing of the web’s critical, bridge-like connections.
A natural consequence of this sparse structure is a fundamental RLVR mechanism for forgetting and learning within the microscopic training process. We find that the degree-2 constraint first forces the system into a maximally frustrated state at the onset of slow learning, a phase dominated by frustration-induced forgetting as skills are competitively suppressed. Subsequently, the same sparse structure enables a cleaner growth mechanism: phase-transition-like learning, where new knowledge is efficiently “bolted on” at the web’s expanding frontier.
This framework is not just descriptive; it is prescriptive. It identifies the maximally frustrated state—the peak of competitive forgetting—as the ideal moment for SFT intervention. Our Annealed-RLVR algorithm is born from this insight. By applying a timed, corrective SFT shock, Annealed-RLVR mitigates the policy collapse and, as our experiments confirm, unlocks a superior final reasoning capability. Ultimately, this work provides a new lens for understanding RLVR: not as black-box optimization, but as a predictable process of structural self-organization, paving the way for more principled methods for engineering the advanced reasoning capabilities of future AI.
Looking forward, a central long-term goal is the empirical mapping of the complete, microscopic reasoning graph from large-scale foundation models. This remains a significant undertaking, requiring substantial computational resources and novel algorithms for semantic chunking and clustering. The payoff for this effort, however, would be transformative, producing a powerful new analytical object: a detailed map of an AI’s internal world, composed of the reasoning graph and its emergent concept web. While the immediate application of this map would be to empirically test our sparse-web hypothesis, its long-term impact is far broader: charting this internal world is a critical step toward genuinely explainable LLMs and the robust safety of future artificial intelligence.
7 Related Work
Our work intersects with three domains: training paradigms for LLM reasoning, the modeling of reasoning as a complex network, and the study of RLVR’s failure modes. While building on the standard RLVR paradigm (Guo et al., 2025; Team et al., 2025; Yang et al., 2025), our goal is to provide a theoretical-physics-motivated theory for its perplexing macroscopic dynamics. To this end, we engage with the view of reasoning as a complex network; however, our approach differs from prior work that empirically extracts static reasoning graphs post-hoc (Xiang et al., 2025; Cabannes et al., 2024; Minegishi et al., 2025). Inspired by the core logic of the RG, we bypass the intractable graph-construction challenge and instead hypothesize a simple organizational principle for the emergent, coarse-grained concept web with $\left<k\right>\approx 2$, enabling us to study the universal behaviros of RLVR. This sparse-web perspective provides a unified origin for seemingly distinct failure modes like catastrophic forgetting (Li & Hoiem, 2017; Ding & Wang, 2025) and policy collapse (Yue et al., 2025; Cui et al., 2025), framing the former as the severing of the network’s fragile bridges, and the latter as a consequence of a phase-transition-like learning dynamics at its frontier(Power et al., 2022; Lin et al., 2025; Cai et al., 2025). An important line of research has made valuable progress on these failures using algorithmic solutions (Chen et al., 2025). Our work contributes a complementary, theory-driven perspective: we leverage the sparse-web model to design Annealed-RLVR, an intervention that targets the problem’s underlying structural root. See Appendix B for extended related work.
8 ACKNOWLEDGMENTS
The authors thank Haijun Zhou, Lei Wang, Hong Zhao and Ganqu Cui for inspiring discussions. K. C. and X. C. are supported by the National Key Research and Development Program of China under Grant No. 2024YFA1408604 and the National Natural Science Foundation of China under Grants No. 12047503 and No. 12447103. P. Z. is supported by Project 12325501 of the National Natural Science Foundation of China. Y. D. and S. H. are supported by the National Natural Science Foundation of China under Grants No. 12275263, the Innovation Program for Quantum Science and Technology under Grant No. 2021ZD0301900, and the Natural Science Foundation of Fujian Province of China under Grant No. 2023J02032. Z. Y. is supported by the National Natural Science Foundation of China under Grants No. 12304288 and No. 12247101.
9 Reproducibility Statement
To ensure full reproducibility, we provide a complete experimental blueprint in the Appendix: Sec. I details the LLM training pipeline—models, hyperparameters, and SFT dataset construction, while Sec. J describes our evaluation protocol. Our code repository of CoNet can be available at: https://anonymous.4open.science/r/CoNet-83A4. The repository includes a README.md file that details the functionality of each key Python script. The training process is described in Appendix D and Appendix H.
Appendix A The Use of Large Language Models (LLMs)
This work utilized Large Language Models (LLMs) to assist in the preparation of the manuscript, primarily for improving language clarity and readability. For the accompanying CoNet repository, we also employed LLMs to generate Python codes, such as the training algorithm, measuring network properties and visualizing the concept web and its largest connected component. All LLM-generated outputs, including text and code, were meticulously reviewed, edited, and verified by the authors to ensure their accuracy and suitability for this work.
Appendix B Extended Related Work
System-2 Reasoning in LLMs. The dual-process theory, distinguishing between fast, intuitive System-1 thinking and slow, deliberate System-2 reasoning (Kahneman, 2011), has become a valuable framework for analyzing LLM behavior. While standard LLMs often exhibit System-1-like responses, significant research has focused on eliciting more robust, System-2 deliberation. Prompting strategies such as Chain-of-Thought (CoT) (Wei et al., 2022) and its variants like Self-Consistency (Wang et al., 2022) and Tree of Thoughts (ToT) (Yao et al., 2023) encourage models to generate explicit reasoning steps. Other works aim to build this deliberative capacity into the model’s architecture or training process through self-correction and reflection (Shinn et al., 2023; Madaan et al., 2023). Our work complements these efforts by shifting the focus from eliciting reasoning to understanding the emergent mechanism of how this capability develops internally during training.
Complex Network Analysis of Neural Models. Applying network science to understand neural networks is a long-standing idea, previously used to study the topology and robustness of smaller models (Estrada & Hatano, 2008). With the advent of Transformers, researchers have begun analyzing attention patterns as graphs (Vig, 2019; Ferrando et al., 2023) to better understand information flow. More recently, studies have explored how network structures emerge during supervised learning, identifying the formation of essential subnetworks (Li et al., 2023). A recent study by (Minegishi et al., 2025) shares our goal of analyzing the network structures that underpin System-2 reasoning, though they do so by empirically extracting a graph from hidden-state representations. In contrast, our CoNet framework utilizes a fixed, abstract concept space, which allows us to specifically isolate and analyze the dynamic mechanism of how this reasoning structure emerges during training.
Appendix C The Reasoning Graph: From Empirical Extraction to a Unified Concept Web
The idea of a reasoning graph has emerged in the literature as a powerful framework for understanding the chain-of-thought process in LLMs for a given task. This graph is not explicitly encoded in the Transformer’s static parameters but is an emergent property of its inference dynamics. The standard model for its construction is based on a key observation: an LLM’s generative process is guided by a few decision points of high Shannon entropy, or “forking tokens” (Fig. 7). This allows for a deconstruction of a chain-of-thought: stable, low-entropy text chunks form the graph’s nodes (concepts), while the probabilistic choices at these high-entropy forks form the weighted, directed edges.
Figure 7: Illustration of the LLM reasoning process. A chain-of-thought is composed of stable, low-entropy text chunks (concepts), connected by high-entropy ’forking tokens’ that represent decision points. The LLM reasoning process can be deconstructed into stable “concepts” (e.g., step 3) linked by high-entropy decision points. The probabilities at these junctures (e.g., “next” vs. “tens”) represent transition weights on an underlying reasoning graph, a view supported by convergent findings on the role of high-entropy tokens.
Recent work, notably by (Minegishi et al., 2025), has developed detailed methods to empirically construct these task-specific reasoning graphs. Their process involves generating a large corpus of reasoning chains, segmenting them into individual steps, and then computing a vector representation for each step by averaging the hidden states of its tokens at a specific Transformer layer. Finally, a clustering algorithm (e.g., K-means) is used on the full set of step vectors, with the resulting cluster centroids being defined as the graph’s nodes. By analyzing graphs constructed this way, they found that more advanced models exhibit distinct topological properties, such as larger diameters, which correlate with improved performance.
While analyzing these static, task-specific graphs is insightful, it leads to a deeper question: what is the underlying, universal knowledge structure that allows an LLM to generate so many different reasoning graphs across a multitude of tasks? It follows that these empirically observed reasoning graphs are best understood as small projections of a much vaster, latent concept web. This web represents the entirety of the model’s abstract knowledge, and the ultimate goal is to understand its structure.
Figure 8: From Intractable LLM Dynamics to a Minimal CoNet Model. The reasoning process in an LLM (left) can be viewed as a traversal on a high-dimensional latent graph, where path probabilities are determined by the complex, autoregressive Transformer. Directly analyzing this graph’s evolution is intractable. CoNet (right) provides a minimal abstraction by replacing the intractable latent graph with a fixed K-regular random graph and simplifying the generation policy to a learnable Markov chain. This creates a tractable “computational microscope” to precisely analyze how the transition policy evolves under reinforcement, isolating the core dynamics of reasoning path formation.
However, attempting to reverse-engineer this universal concept web from its sparse, task-specific projections presents formidable obstacles. The difficulties are multi-fold. First is a problem of scale: the true concept web is of astronomical size, and any extracted graph is an infinitesimal sample. More fundamentally is the ambiguity of the concept mapping: the mapping from a low-level token sequence to a high-level semantic concept is inherently many-to-one. Finally, this already-ambiguous mapping is itself dynamic, as it can shift during RLVR training. These challenges collectively lead to the observable node stability problem in practice, making it nearly impossible with current methods to track the web’s true topological evolution.
It is precisely this observational challenge that our paper bypasses with a two-pronged, indirect approach. First, we form a theoretical hypothesis about the consequences of a specific topological structure (i.e., sparsity). Second, we use a minimal model (CoNet) that captures the essential physics of the system to study these consequences in a controlled environment. Thus, while the direct construction of the concept web remains an open problem, our work proceeds from a central hypothesis about its fundamental structure. This Hypothesis—that the web is profoundly sparse with an average degree near two—serves as the theoretical foundation for our paper. From it, we first demonstrate how this assumption of sparsity explains the known macroscopic dynamics of RLVR. We then use it to deduce previously overlooked microscopic mechanisms, namely frustration-induced forgetting and phase-transition-like learning. Finally, armed with these new insights, we explore how this deeper understanding can be leveraged to design a novel, improved RLVR algorithm.
Appendix D CoNet: A Minimal Model for RLVR dynamics
To bypass the intractable node stability problem, we leverage the Concept Network Model (CoNet). CoNet was originally introduced in (Cai et al., 2025) to study the dynamics of single-task RLVR, but its core design makes it perfectly suited for our investigation. By positing a fixed space of concept nodes (a static graph structure) and isolating all learning dynamics to the evolving policy that navigates it, CoNet allows us to sidestep the challenges of direct graph extraction and precisely measure how the reinforcement signal sculpts reasoning paths. The model represents the LLM’s concept space as an abstract K-regular random graph, with the reasoning process simplified to a policy-guided Markovian random walk, as shown in Fig.8. The transition policy, $\pi_{\theta}(j|i)\propto\theta_{ij}$, is governed by a set of learnable parameters $\theta_{ij}$ that represent the strength of the connection from node i to node j. This policy is updated via RLVR on successful paths from a question (Q) to an answer (A).
This design offers a clear cognitive analogy. A single-step hop on the graph represents “System 1” thinking—fast, intuitive associations. A multi-step traversal, guided by the learned policy, represents “System 2” reasoning—a deliberate, compositional chain of thought. It is crucial to note that CoNet is not a model of the Transformer’s parameters themselves; rather, it models the emergent semantic network induced by inference—the structure that governs how questions are connected to answers through chains of concepts.
CoNet was originally introduced to model the phenomenon of Learning at Criticality (LaC), which was identified in single-task RLVR training (Cai et al., 2025). The central finding of LaC is that for a model trained on extremely sparse data (e.g., a single exemplar), generalization performance on unseen problems peaks precisely at a sharp, sigmoidal learning transition, before declining due to overfitting. The original work used CoNet to establish that this dynamic is a learning phase transition, providing a physical model for this peak in generalization.
The central methodological advance of this paper is to deploy this existing model in a richer, multi-task setting with numerous concurrent Q-A pairs. This extension is crucial: it transforms CoNet from a model of isolated skill acquisition into an ideal theoretical laboratory for studying skill integration. In this regime, the model is compelled by competing reinforcement signals to discover a shared, reusable structure. As Fig. 1 shows, this multi-task CoNet stunningly reproduces the key macroscopic signatures of the 1.5B LLM. This correspondence validates our use of CoNet as a minimal model for exploring the network-level mechanisms of frustration and web formation.
CoNet was originally introduced to model the phenomenon of Learning at Criticality (LaC), which was identified in single-task RLVR training (Cai et al., 2025). The central finding of LaC is that for a model trained on extremely sparse data, generalization performance peaks precisely at a sharp learning transition. The key methodological advance of this paper is to deploy this model in a richer, multi-task setting. This extension is crucial: it transforms CoNet from a model of isolated skill acquisition into an ideal theoretical laboratory for studying the emergent phenomena of skill integration, frustration, and concept web formation. As Fig. 1 shows, this multi-task CoNet stunningly reproduces the key macroscopic signatures of the 1.5B LLM, validating its use for exploring the network-level mechanisms that drive the emergence of reasoning.
Appendix E CoNet Implementation Details
We configured CoNet with a specific set of structural and learning parameters. The underlying concept space was instantiated as a directed random regular graph with $N=800$ nodes, where each node has a uniform out-degree of $k=40$. In this graph, we established a multi-task learning environment consisting of $128$ randomly generated question-answer (Q-A) node pairs. The initial transition parameters $\theta_{ij}$ for all edges were assigned randomly. The use of a multi-task setting with 128 concurrent Q-A pairs is the fundamental mechanism that induces structural competition. While a single task would only require finding one efficient path, the presence of numerous tasks provides competing reinforcement signals. When multiple Q-A paths need to traverse a common node, they may “disagree” on the optimal subsequent step, leading to conflicting policy updates. This microscopic conflict is the origin of the macroscopic “frustration” discussed in the main text, compelling the system to find a shared, globally coherent structure that can arbitrate these conflicts, rather than settling for 128 disconnected “skill islands”.
The policy was optimized using a variant of the Group Relative Policy Optimization (GRPO) algorithm ((Shao et al., 2024; Liu et al., 2025; Yu et al., 2025)). For a given Q-A pair, $n_{\text{rollout}}=128$ reasoning paths (indexed by $m$) were sampled, with each path’s exploration capped at a maximum length of 20 steps. Each path received a reward based on its success, and its advantage $A_{m}$ (relative to the average reward over all 128 rollouts) was used to guide the update of the policy parameters $\theta_{ij}$. The update rule is given by:
$\Delta\theta_{ij}\propto\sum_{m}A_{m}\nabla_{\theta_{ij}}\log\pi_{\theta}(j|i)$
This advantage-based update functions as a competitive reinforcement mechanism. Only paths performing better than the current average are strengthened, while underperforming paths are implicitly suppressed. This “winner-take-more” dynamic acts as an algorithmic sculpting tool, systematically pruning the vast majority of the initial $k=40$ edges from each node to carve out the sparse, efficient backbone of the concept web. The learning rate was set to $0.04$ to control the speed and stability of this training process.
Crucially, the two-stage learning dynamic is not an artifact of fine-tuned parameters but a robust emergent property of the model. Repeated experiments show that this behavior—a fast-learning phase followed by a slow integration plateau, and the corresponding V-shaped response length—consistently emerges under a broad condition: when the number of Q-A pairs, $N_{QA}$, is in the regime of $1\ll N_{QA}\ll N$. This dynamic is qualitatively different from the single Q-A pair ($N_{QA}=1$) case, which produces a simple sigmoidal learning transition (the LaC phenomenon). The condition is essential: if $N_{QA}$ is too small, there is no systemic competition to drive the formation of a unified web; if it is too large, the conflicting reinforcement signals prevent any coherent structure from forming. The robust emergence of these LLM-like signatures in this intermediate regime confirms that our CoNet setup successfully isolates the core principles of structural self-organization under multi-task reinforcement.
Appendix F Structural Evolution of the Concept Web in CoNet
This section provides direct, quantitative evidence from the CoNet model for the two-stage learning dynamic—a fast acquisition phase followed by a slow integration phase—that underlies our theory. Fig. 9 serves as the central exhibit for this analysis, tracking two key structural properties over the course of training: the total number of disconnected solution clusters (the “skill islands”) and the size of the single largest cluster, namely the concept web.
Figure 9: The Formation of the Concept Web in CoNet. This figure illustrates the evolution of network clusters during training, showing a structural transition from isolated skills to a unified conceptual web. In the initial phase, the Cluster Number (orange curve) spikes, indicating the rapid discovery of numerous disconnected “skill islands”. The peak marks a critical transition point. Following this peak, the Cluster Number steadily declines as these islands begin to coalesce. This merging process is confirmed by the corresponding monotonic growth in the Max Cluster Size (blue curve), which represents the formation of a single giant component. By the end of training, the network is dominated by this large, interconnected structure.
The first stage of training ( $\sim$ 0-50 steps), corresponding to the fast-learning phase, is characterized by the rapid proliferation of “skill islands”. The quantitative evidence for this is the steep, initial rise of the “Cluster Number” (orange curve) in Fig. 9. Each new cluster represents the discovery of a distinct, high-confidence reasoning path for a specific problem. During this phase, the system operates in a greedy, parallel fashion, finding many efficient, local solutions without regard for their interconnection.
The transition to the slow-learning phase occurs at the critical juncture marked by the peak of the orange curve around step 20. At this point, the system has exhausted the “low-hanging fruit” of simple, isolated solutions and has reached a state of maximal fragmentation. This peak signals the onset of the “maximally frustrated state” discussed in the main text. The subsequent, gradual decline of the “Cluster Number” curve is the direct signature of this frustrated state, representing the arduous and competitive process of integrating existing islands rather than discovering new ones. The “Max Cluster Size” (blue curve), which begins its steep, sustained ascent at precisely this moment, confirms this shift from local discovery to a global integration objective.
The slow-learning phase is thus defined by the coalescence of these islands into a unified concept web. The plot reveals the two complementary trends that define this stage: a steady decline in the total number of clusters (orange curve) and a simultaneous, relentless growth in the size of the largest cluster (blue curve). The causal link is direct: the largest component grows precisely by absorbing smaller, independent clusters, which necessarily reduces the total cluster count. This structural evolution culminates in a state where a single, giant component—the functional concept web—dominates the network. This process provides a direct, mechanical explanation for the V-shaped response length signature: Stage 1 shortens paths by finding local solutions, while Stage 2 lengthens them by forcing the policy to traverse longer, connective paths between previously distant concepts.
Appendix G Learning by Phase Transition in the Slow-Learning Stage
Figure 10: Per-Problem Critical Dynamics During the Slow-Learning Stage. These plots reveal the microscopic learning dynamics for a representative subset of problems that successfully learn during the slow-learning stage, providing evidence for localized phase transitions. (Left) Accuracy trajectories for these problems. All problems shown exhibit a sharp, sigmoidal-like increase in accuracy late in the training process, characteristic of a critical learning event. (Right) Corresponding variance of the solution path length. A sharp, pronounced peak in variance directly coincides with the rapid accuracy gain for each problem. This dual signature is the hallmark of a continuous phase transition, suggesting that the slow-learning stage in CoNet is composed of discrete, sharp learning events rather than a uniform, gradual process.
In the main text, we argue that new knowledge acquisition during the slow-learning stage is not gradual but occurs via discrete, sharp events. This appendix provides detailed, per-problem evidence for this learning by phase transition mechanism, drawing a direct analogy to critical phenomena in statistical physics.
Fig. 10 isolates a representative subset of problems mastered late in the training process. The first signature of a phase transition is visible in the left panel: each problem exhibits a sharp, sigmoidal-like jump in accuracy. In the language of statistical physics, accuracy acts as an order parameter, signaling the system’s transition from a disordered (unsolved) state to an ordered (solved) state. An abrupt change in the order parameter is a primary indicator of a phase transition.
The second, more telling signature is the pronounced, transient peak in the variance of the solution path length, shown in the right panel. This provides a deeper link to the physics of critical phenomena. In physical systems, the variance (or fluctuations) of an order parameter is related to the system’s susceptibility—its response sensitivity to external perturbations. At the critical point of a continuous phase transition, this susceptibility is known to diverge, creating a characteristic sharp peak. A famous example of this is the lambda ($\lambda$) peak observed in the specific heat of liquid helium at the normal-to-superfluid transition. The variance peaks in our figure are the direct analogue of this phenomenon. This “dual signature”—an abrupt rise in the order parameter (accuracy) and a lambda-like peak in its fluctuations (variance)—is the classic hallmark of a critical learning transition. This same analysis was performed in the original work on LaC, where the variance peak in the single-task CoNet was identified as a key hallmark of a learning phase transition (Cai et al., 2025).
The intuitive reason for this variance peak is straightforward. When accuracy is very low, nearly all sampled paths fail, resulting in low variance. When accuracy is very high, the model has converged on a dominant, efficient reasoning path, again leading to low variance. It is only during the narrow transition window that a rich diversity of paths coexists—some still failing, while others explore newly viable routes to the solution. This temporary coexistence of competing strategies is what produces the characteristic peak in variance.
Taken together, these results provide strong microscopic validation for our claims. They show that the slow-learning stage in CoNet is not a process of uniform, gradual improvement. Instead, it is punctuated by a series of discrete, critical learning events, where individual skills are integrated into the concept web one by one via localized phase transitions at the frontier.
Appendix H The SFT and Annealed-RLVR Algorithm in CoNet
This appendix details the implementation and validation of the Annealed-RLVR algorithm in CoNet, our “computational microscope” for LLM reasoning dynamics. The algorithm is designed to intervene at the point of “maximum frustration” with SFT to resolve the competitive bottlenecks that arise during the formation of the sparse concept web.
H.1 SFT Implementation in CoNet
CoNet’s minimalistic policy model significantly simplifies the standard SFT process. This process involves first using the original model’s policy to sample a successful reasoning path for a given problem. Then, for this selected path, we adjust its transition probabilities. For any transition on the path with a probability below a certain threshold, we raise it to that threshold value. To ensure that the probability distribution of each node remains valid, the probabilities of other outgoing transitions are proportionally reduced. The value of this target transition probability is a tunable parameter that determines the intervention’s effect. For our annealing strategy, we set this value to 0.1. To induce catastrophic forgetting, we use a more aggressive value of 0.5.
H.2 Expeimental Results
Figure 11: Validation of Annealed-RLVR in CoNet. (Left) The histogram of per-problem accuracy shows that the annealed model (SFT-CKPT750) significantly reduces the number of completely unsolved problems (Accuracy = 0.0) and increases the number of mastered problems (Accuracy = 1.0) compared to the standard model (CKPT800). (Right) The line plot of training dynamics shows that immediately after the SFT intervention at step 50, best@1 accuracy dips while best@16 accuracy sharply increases, indicating a successful trade of exploitation for exploration. The final annealed policy (“Annealed”) ultimately achieves superior performance over the baseline (“RLVR”).
The algorithm consists of a strategically timed Supervised Fine-Tuning (SFT) phase, followed by a resumption of standard Reinforcement Learning from Verifiable Rewards (RLVR) training. The SFT intervention is applied at training step 50, the empirically determined point of maximum frustration where the system transitions from forming disconnected “skill islands” to integrating them. At this stage, we first identify all problems with an accuracy below 0.1, select a total of 50 low-accuracy problems, and sample a known successful reasoning path for each. This targeted adjustment serves as the “heating” step in our simulated annealing analogy; it gently increases the discoverability of latent correct paths to restore the system’s exploratory capacity. Following this brief SFT, standard RLVR training is resumed, which functions as the “cooling” phase, allowing the now more exploratory policy to settle into a new, more globally optimal configuration.
The success of this SFT-RLVR cycle is confirmed by the results in Fig. 11. The line plot on the right shows the immediate effect of the intervention: best@1 accuracy (a proxy for exploitation) temporarily dips, while best@16 accuracy (a proxy for exploration) surges. This demonstrates a successful trade of greedy exploitation for enhanced exploration, and the long-term benefit is clear, as the final annealed policy significantly surpasses the baseline RLVR policy. The histogram on the left reveals the underlying mechanism for this improvement. The SFT-RLVR cycle successfully resolves the frustration bottleneck by dramatically reducing the population of completely unsolved problems (Accuracy = 0.0) and correspondingly increasing the number of fully mastered problems (Accuracy = 1.0) compared to the standard model.
In summary, the CoNet experiment provides strong evidence for our theory. By applying a targeted and conservative SFT intervention at the point of maximum frustration, the Annealed-RLVR algorithm effectively resolves the system’s competitive bottlenecks, enabling it to discover a more robust and better reasoning policy.
Appendix I LLM Training Details
I.1 Base Model and GRPO Setup
We used the DeepSeek-R1-Distill-Qwen-1.5B model (Guo et al., 2025), trained on the DeepScaleR-Preview-Dataset (Luo et al., 2025) with their modified version of the veRL open-source engine (Sheng et al., 2024; ByteDance Seed Team and verl community, ). Our Generative Reward Policy Optimization (GRPO) hyperparameters followed the first-stage settings, which specified a response length of 8192, as outlined in the DeepScaleR paper (Luo et al., 2025).
I.2 SFT Dataset for Annealed-RLVR
The Supervised Fine-Tuning (SFT) dataset for the annealed-RLVR experiment (labeled “Annealed” in Fig. 6) was constructed to focus on difficult problems. First, we trained the base model using GRPO for 100 steps to create a checkpoint (CKPT100), which is in the vicinity of the maximally frustrated state. We then used this checkpoint to generate 8 rollouts for each of the approximately 40,000 problems in the training set. For the roughly 10,000 problems that yielded no correct responses, we generated an additional 52 rollouts, bringing the total to 60 per problem. From this data, we created the SFT dataset by selecting all correct responses for problems where the initial model accuracy was below 10%, which yielded 3,933 trajectories. Finally, we fine-tuned the model on these trajectories for two epochs with a learning rate of $3\times 10^{-5}$. The resulting model served as the new starting point for continuing GRPO training up to step 1,700.
I.3 SFT Dataset for Catastrophic Forgetting
The SFT dataset for the catastrophic forgetting experiment (labeled “SFT+RLVR” in Fig. 4(b)) was designed to specifically target problems forgotten by a more saturated model, using a data generation process that involved multiple checkpoints. The process began with problem filtering: we used a checkpoint from step 600 (CKPT600, deep within the slow-learning stage) to generate 8 rollouts for all training problems. We then identified the subset of problems with no correct responses and generated another 32 rollouts for them, again using CKPT600. Problems that still had no correct responses were selected as the final “forgotten” set. For the second stage of targeted data generation, each problem in this forgotten set was used to generate a total of 96 rollouts: 32 from the base model, 32 from CKPT100, and 32 from CKPT600. We then collected correct responses from problems that had at least 3 correct answers among these 96 rollouts, resulting in 798 trajectories. This SFT dataset was used to fine-tune the model for two epochs with a learning rate of $5\times 10^{-5}$ before resuming GRPO training.
Appendix J LLM Evaluation Details
We evaluated checkpoints from our RLVR (“GRPO”) and annealed-RLVR (“Annealed”) training runs on two datasets: 512 randomly selected training problems and the Minerva math dataset (Hendrycks et al., 2021; Lewkowycz et al., 2022). The evaluation parameters were kept consistent with the training configuration (temperature: $0.6$, top-p: $1.0$, response length: $8192$, etc.). From the $512$ rollout results, we calculated the best@k accuracy and the standard error. Fig. 6 illustrates the full performance curves, while Table 1 highlights the results from representative checkpoints.
Table 1: Best@k Accuracy (%) with Standard Error at Key Training Checkpoints
Checkpoints (Steps)
Metric Method 0 100 200 600 1000 1400 1800
Dataset: Randomly Selected 512 Training Problems
best@1 GRPO 43.66(6) 50.17(6) 52.45(6) 56.49(6) 58.00(6) 58.90(6) 59.55(6)
Annealed 50.00(6) 52.27(6) 55.80(6) 58.10(6) 59.46(6) 60.31(6)
best@4 GRPO 61.32(8) 68.46(8) 70.38(8) 72.99(7) 73.46(7) 73.71(7) 74.20(7)
Annealed 69.49(8) 71.51(8) 73.75(7) 75.10(7) 75.91(7) 76.77(7)
best@16 GRPO 72.97(12) 78.60(11) 80.12(11) 81.60(10) 81.61(10) 81.62(10) 81.86(10)
Annealed 80.62(11) 82.22(11) 83.15(10) 83.77(10) 84.48(10) 85.28(10)
best@128 GRPO 82.16(24) 86.98(23) 87.81(20) 87.47(19) 88.23(21) 87.77(19) 87.99(22)
Annealed 88.00(19) 89.35(20) 89.37(19) 89.98(18) 90.23(18) 91.22(18)
Dataset: Minerva
best@1 GRPO 23.85(7) 26.08(7) 26.83(7) 27.63(7) 28.53(7) 28.45(7) 28.77(7)
Annealed 26.09(7) 26.62(7) 27.73(7) 28.28(7) 28.75(7) 28.88(7)
best@4 GRPO 37.84(12) 40.77(12) 41.33(12) 41.12(12) 41.46(11) 41.04(11) 41.32(11)
Annealed 40.63(12) 40.90(12) 41.66(12) 42.04(11) 42.53(11) 42.39(11)
best@16 GRPO 51.39(21) 54.29(21) 54.20(21) 53.05(20) 52.38(19) 52.12(20) 51.95(19)
Annealed 53.80(21) 53.71(21) 53.74(20) 53.86(21) 54.24(21) 53.50(20)
best@128 GRPO 66.6(5) 68.8(4) 69.8(5) 66.2(4) 65.8(5) 64.9(4) 64.4(4)
Annealed 68.2(5) 68.3(4) 68.4(5) 68.0(4) 69.4(5) 67.8(5)

### v2

@@
 Abstract
-Training large language models with Reinforcement Learning from Verifiable Rewards (RLVR) exhibits a set of distinctive and puzzling behaviors that remain poorly understood, including a two-stage learning curve, V-shaped response-length trajectories, and a pronounced vulnerability to catastrophic forgetting.
-In this work, we propose that these seemingly disparate phenomena can be explained using a single unifying theory: the model’s reasoning process maps to the self-organization of a semantic complex network whose topology remains persistently sparse, with the average degree pinned close to two.
-This topology imposes a fundamental mechanism for forgetting and learning: it first drives the system into a maximally frustrated state where “skill islands” form, slow-learning happens, and forgetting is induced; then it enters a sharp growth phase where the new skills are “bolted on”, driven by phase-transition-like learning at the web’s frontier.
-Equipped with the theory, we propose Annealed-RLVR, a principled algorithm that introduces an SFT-based “heating” step at the point of maximal frustration to resolve the competitive bottleneck and enhance the reasoning capability of the model.
-Experiments on a 1.5B-parameter model demonstrate that the approach outperforms standard RLVR on both in-distribution and out-of-distribution benchmarks.
+Training large language models with Reinforcement Learning with Verifiable Rewards (RLVR) exhibits a set of distinctive and puzzling behaviors that remain poorly understood, including a two-stage learning curve, a V-shaped response-length trajectory, and a pronounced vulnerability to catastrophic forgetting.
+In this work, we propose that these behaviors are emergent collective phenomena governed not by neural implementation details, but by the topological evolution of the latent reasoning graph in semantic space.
+By demonstrating a dynamical isomorphism between a 1.5B-parameter LLM and a minimal Concept Network Model (CoNet), we trace the causal source to the self-organization of a sparse concept web pinned to an average degree of two.
+This geometric perspective provides a unified physical explanation for the observed anomalies: the V-shaped trajectory tracks the evolution from parallel local skill optimization to global network integration; catastrophic forgetting stems from the topological disconnection of critical “trunk” edges; and policy collapse arises from the accumulation of sequential transitions at the web’s leaf nodes, where broad exploration abruptly freezes into rigid, high-reward trajectories.
+Identifying a “maximally frustrated state” at the transition between learning stages, we propose Annealed-RLVR, a principled algorithm that injects a targeted SFT “heating” step to resolve this topological bottleneck.
+Experiments confirm that this theory-driven intervention outperforms standard RLVR on both in-distribution and out-of-distribution benchmarks (including Minerva and AIME).
 By recasting RLVR from black-box optimization into a predictable process of structural self-organization, our work provides a new physical intuition for engineering the emergent reasoning capabilities of future AI systems.
 1 Introduction
-Large language models (LLMs) trained with Reinforcement Learning from Verifiable Rewards (RLVR) (Liu et al., 2025; Guo et al., 2025; Team et al., 2025; Yang et al., 2025) are increasingly capable of multi-step, “System-2” reasoning.
-Yet, this training paradigm consistently exhibits a set of puzzling macroscopic behaviors across diverse architectures and benchmarks.
-These include (i) rapid initial gains that give way to a prolonged plateau; (ii) a V-shaped trajectory in the length of correct solutions, which first shortens before steadily lengthening (Luo et al., 2025; He et al., 2025; Fatemi et al., 2025); and (iii) a pronounced vulnerability to catastrophic forgetting when interleaved with supervised fine-tuning (SFT) (Li & Hoiem, 2017; Fernando et al., 2025; Ding & Wang, 2025).
-Although these phenomena are widely observed, the reasons for their co-occurrence remain elusive.
-A clear understanding of the underlying mechanism would be essential for developing more efficient approaches to training reasoning-capable models.
-A promising theoretical framework analyzes these phenomena by modeling the LLM’s reasoning process on an implicit reasoning graph — a network of semantic states and logical transitions(Wang et al., 2024; Dutta et al., 2024; Xiang et al., 2025; Cabannes et al., 2024; Minegishi et al., 2025).
-However, this perspective faces a critical roadblock: a direct, microscopic analysis is currently intractable.
-Constructing this entire graph from a model’s high-dimensional latent space remains a formidable open problem, preventing a direct investigation into the structural origins of the observed dynamics.
-Here, we argue that the very universality of the RLVR phenomena provides a path forward.
-In complex systems, behaviors that are independent of microscopic details are often emergent properties, governed by the system’s large-scale organization (Anderson, 1972; Watts & Strogatz, 1998; Barabási & Albert, 1999).
-This insight motivates an approach rooted in the core logic of the Renormalization Group (RG)(Wilson & Kogut, 1974): instead of getting bogged down by intractable local details, one can leap directly to hypothesizing a core principle about the system’s large-scale structure to explain its macroscopic behavior.
-Applying this strategy, we shift our focus from the full reasoning graph to its coarse-grained backbone, which emerges by abstracting away fine-grained skills encoded in local network patterns like motifs or community structures.
+Training large language models with reinforcement learning with verifiable rewards (RLVR) (Liu et al., 2025; Guo et al., 2025; Team et al., 2025; Yang et al., 2025) has revealed four striking and still poorly understood phenomena.
+On math and coding benchmarks, RLVR runs typically show (i) rapid initial gains followed by a long performance plateau, even as training continues (Luo et al., 2025a; He et al., 2025; Jin et al., 2025b); (ii) a V-shaped trajectory of the length of correct solutions, where answers first become markedly shorter before later growing longer again (Luo et al., 2025a; He et al., 2025; Fatemi et al., 2025; Wang et al., 2025); (iii) pronounced vulnerability to catastrophic forgetting when RLVR-trained policies are interleaved or followed by supervised fine-tuning (SFT), with earlier reasoning skills abruptly degraded (Li & Hoiem, 2017; Luo et al., 2025b; Ding & Wang, 2025); and (iv) a collapse of solution diversity, with the policy converging onto a narrow and brittle solution manifold (He et al., 2025; Cui et al., 2025; Hao et al., 2025; Jiang et al., 2025).
+Existing work offers only partial, largely phenomenological explanations for each phenomenon in isolation: plateaus and the collapse of solution diversity are linked to entropy exhaustion and over-exploitation of a fixed reward signal (Cui et al., 2025; Hao et al., 2025; Jiang et al., 2025); the initial shortening of solutions is attributed to reward-driven pruning of redundant reasoning, whereas the subsequent lengthening has been ascribed either to a shift toward solving harder instances that demand longer chains of thought, or to the emergence of multi-pass, self-reflective reasoning (“aha”-like behavior) under RLVR(Fatemi et al., 2025; Wang et al., 2025); and catastrophic forgetting is analyzed as an objective-mismatch problem between RLVR and subsequent SFT or other post-training stages (Luo et al., 2025b; Ding & Wang, 2025).
+However, these accounts remain fragmented: we still lack a unifying framework that ties all four behaviors to a common underlying structure or dynamical mechanism of RLVR-trained policies, and explains why they so often arise together in practice.
+A growing body of work tries to explain these phenomena by modeling LLM’s reasoning process as an implicit reasoning graph—a network of semantic states and logical transitions (Wang et al., 2024; Dutta et al., 2024; Xiang et al., 2025; Cabannes et al., 2024; Minegishi et al., 2025).
+However, translating this idea into a fully microscopic analysis has proven extremely difficult: constructing the reasoning graph from a model’s high-dimensional latent space is a formidable open problem, preventing a direct investigation into the structural origins of the observed dynamics (See Appendix C for details).
+Here, we argue that the very universality of the above RLVR dynamics provides a path forward.
+In complex systems, collective emergent behaviors are largely independent of microscopic details and governed by the system’s large-scale organization (Anderson, 1972; Watts & Strogatz, 1998; Barabási & Albert, 1999).
+This insight motivates an approach rooted in the core logic of the Renormalization Group (RG) (Wilson & Kogut, 1974): instead of directly studying the system at the full microscopic level, one can model and approach the system at a coarse-grained level to explain its macroscopic behavior.
+Applying this strategy, we shift our focus from the full reasoning graph at the token level to its coarse-grained backbone at the semantic level, which emerges by abstracting away fine-grained skills encoded in local network patterns like motifs or community structures (Milo et al., 2002; Fortunato, 2010; Zhang & Moore, 2014).
 This leads us to our central hypothesis regarding the emergent structure that forms under RLVR training for general-purpose System-2 reasoning:
@@
 (b) LLM
-Figure 1: CoNet Faithfully Reproduces Core LLM Training Dynamics.
-The minimal CoNet model (a) remarkably reproduces the two core empirical signatures of RLVR training observed in the DeepScaleR-1.5B LLM (b).
-These signatures are: (i) a two-stage reward dynamic consisting of a fast-learning phase followed by a slow-learning plateau (top panels), and (ii) a non-monotonic, V-shaped evolution of the correct response length (bottom panels).
-This striking correspondence validates CoNet as a tractable minimal model for our theoretical analysis.
+Figure 1: CoNet Reproduces Core LLM Training Dynamics.
+The minimal CoNet model (a) reproduces the two core empirical signatures of RLVR training observed in the DeepScaleR-1.5B LLM (b).
+These signatures are: (i) a two-stage reward dynamic consisting of a fast-learning stage followed by a slow-learning phase (top panels), and (ii) a non-monotonic, V-shaped evolution of the correct response length (bottom panels).
+This striking correspondence validates CoNet as a valuable minimal model for our theoretical analysis.
 The inset in (a) schematically depicts the CoNet abstraction.
 Lines in (b) are smoothed with a 60-step moving average.
-To validate this hypothesis, we combine two approaches: a theoretical study using the Concept Network Model (CoNet) (Cai et al., 2025), a minimal “computational microscope” that reproduces the LLM’s core training dynamics, and numerical experiments on a 1.5B-parameter LLM(Luo et al., 2025).
-This dual approach reveals a clear structural narrative for the observed puzzles.
-The learning plateau is shown to be the arduous consolidation of disconnected “skill islands”, a transition that directly explains the V-shaped length trajectory.
-Furthermore, the resulting sparse, tree-like concept web accounts for the model’s fragility and its vulnerability to catastrophic forgetting with SFT.
-This sparse topology imposes a microscopic dynamic of learning and forgetting that is a hallmark of the RLVR-driven self-organization process.
-We find the degree-2 constraint first forces the system into a maximally frustrated state at the onset of the slow-learning plateau, a phase dominated by frustration-induced forgetting as skills are competitively suppressed.
-Subsequently, the very same structure enables the growth mechanism that governs the entire slow-learning phase: a steady, sequential process where new skills are individually “bolted on” at the web’s expanding frontier via a phase-transition-like learning dynamic.
-Equipped with the sparse-web theory, we propose Annealed-RLVR, a principled algorithm that introduces an SFT-based “heating” step at the point of maximal frustration to resolve the competitive bottleneck and enhance exploration and reasoning efficacy of the subsequent RLVR phase.
-Experiments on both CoNet and the 1.5B-parameter LLM confirm the efficacy of this strategy: Annealed-RLVR surpasses standard RLVR on both in-distribution and out-of-distribution benchmarks, transforming our physical intuition into a practical tool for engineering emergent reasoning.
+To investigate this hypothesis, we adopt a comparative approach connecting macroscopic phenomenology to graph topology.
+We first confirm that the DeepSeek-R1-Distill-Qwen-1.5B model (Luo et al., 2025a) exhibits the aforementioned RLVR signatures: a two-stage learning curve, a V-shaped response trajectory, catastrophic forgetting, and policy collapse.
+We then show that these complex behaviors are qualitatively reproduced by the Concept Network Model (CoNet) (Cai et al., 2025), a minimal framework abstracting away neural implementation details.
+This striking resemblance suggests that these phenomena are emergent collective behaviors rather than artifacts of specific neural architectures, allowing them to be studied via a coarse-grained proxy.
+Employing CoNet, we identify the emergence of a sparse concept web that orchestrates these dynamics.
+This phenomenological alignment supports the hypothesis that similar topological self-organization underlies LLM reasoning, enabling us to derive the following mechanisms from the principles of sparse graph growth:
+First, the V-shaped response length implies a structural evolution from local “skill islands” to a global concept web.
+The initial decrease results from the parallel optimization of independent skills, where the policy rapidly discards inefficient exploratory detours to converge on short, optimal paths for simple tasks (Fig.
+2).
+The subsequent rise, however, marks the onset of global integration, where a tree-like web emerges (Fig.
+3).
+On such a sparse substrate, the average geodesic distance is inherently correlated with network size.
+Thus, as the web expands to bridge distant islands during the slow-learning phase, the policy is geometrically forced to traverse longer paths.
+The rising V-curve is therefore not a change in reasoning style, but the result of the concept web’s continuous growth.
+Second, the sparse topology explains the model’s specific vulnerability to catastrophic forgetting (Fig.
+4).
+Unlike dense networks, a sparse web relies on critical “trunk” edges to support vast downstream branches.
+We show that SFT-induced forgetting acts as a topological disconnection: gradient updates overwrite weights at these high-traffic bottlenecks, severing bridges to entire sub-trees.
+This offers a geometric rationale for the “easy-to-break, easy-to-fix” nature of RLVR models; since the disconnected downstream knowledge remains intact, a brief RLVR resumption leads to rapid recovery by merely “re-soldering” the broken trunks rather than relearning the web.
+Third, we trace policy collapse to the dynamics of learning at the web’s frontier (the leaf nodes).
+Extending the critical learning theory of Cai et al.
+(2025), our experiments suggest that during the slow-learning phase, a significant portion of tasks sequentially undergo continuous-phase-transition-like learning dynamics (Fig.
+5).
+For these tasks, the policy shifts from broad exploration to locking onto specific solution paths.
+While this collapse is necessary for mastering individual skills, standard RLVR accumulates these local entropy reductions across the leaf nodes.
+This aggregation ultimately results in a global loss of diversity, manifesting as the policy collapse observed in reasoning benchmarks (Fig.
+6).
+Finally, identifying a “maximally frustrated state” at the onset of slow learning—where competition between disconnected islands peaks—motivates our Annealed-RLVR algorithm.
+By introducing a targeted SFT “heating” step at this topological bottleneck, we temporarily relax competitive tension, helping the system navigate past local optima before cooling into a refined structure.
+Our experiments confirm that this theory-driven approach resolves the bottleneck and achieves superior reasoning performance on in-distribution and out-of-distribution benchmarks (including Minerva and AIME), surpassing standard RLVR baselines by engineering the topology of reasoning.
 (a) Step 20
@@
 (c) Step 800
-Figure 2: Visualizing the Structural Evolution from Skill Islands to the Concept Web.
+Figure 2: Visualizing the Structural Evolution from “Skill Islands" to the Concept Web.
 These network snapshots provide a direct, qualitative narrative of the concept web’s formation, corresponding to the different phases of training.
+In each subfigure, green, red, and blue dots indicate the question, answer, and intermediate nodes of the CoNet, respectively.
+The color of each directed edge represents the transition probability from the head to the tail and is drawn according to the colorbar in the rightmost column.
+Here, only edges with transition probabilities greater than $0.95$ are shown.
 (a) In the early fast-learning phase (Step 20), the model discovers a few short, disjointed reasoning paths, representing the first nascent “skill islands”.
@@
 2 The V-Shaped Signature as Evidence for a Sparse Concept Web
-The V-shaped evolution of response length is a well-demonstrated, quantifiable hallmark of deep structural reorganization during RLVR training(He et al., 2025; Luo et al., 2025).
-To anchor our investigation, we first reproduce this signature by applying RLVR to the DeepSeek-R1-Distill-Qwen-1.5B model following the DeepScaleR protocol (Luo et al., 2025).
+The V-shaped evolution of the response length is a well-demonstrated, quantifiable hallmark of deep structural reorganization during RLVR training (He et al., 2025; Luo et al., 2025a).
+To anchor our investigation, we first reproduce this signature by applying RLVR to the DeepSeek-R1-Distill-Qwen-1.5B model following the DeepScaleR protocol (Luo et al., 2025a).
 As shown in Fig.
@@
 1).
-We contend that this is not a mere by-product but the primary empirical evidence for our central claim: the emergence of a sparse, unified concept web.
 To dissect the mechanism underlying this signature—a task intractable within the LLM itself—we deploy an extended, multi-task variant of the CoNet as a “computational microscope” (See Appendix D for details).
+CoNet instantiates the implicit reasoning graph discussed in the Introduction as a concrete, tractable model: it maps the LLM “semantic states” to abstract nodes in a fixed random graph and LLM “logical transitions” to learnable, probabilistic edges.
 This minimal model reduces the learning process to a tractable graph-traversal problem and, remarkably, reproduces the V-shaped curve observed in our LLM experiment (Fig.
 1), validating it as a faithful proxy for uncovering the underlying mechanism.
-Armed with this confirmation, we now expose the topological origin of the V-shape.
-The V-shaped curve arises from a distinct two-stage process.
-During the initial fast-learning phase where response length decreases, the model undergoes local optimization: it rapidly discovers short, efficient solutions to individual problems, giving rise to a proliferation of disconnected “skill islands” (as shown in Fig.
+(a) Correct Response Histogram
+(b) Step 50
+(c) Step 800
+Figure 3: A Sparse-Web Structure Necessitates Longer Reasoning Chains.
+This figure links the emergent sparse topology of the concept web to the observed increase in response length.
+(b, c) The color and marker scheme is identical to that used in Fig.
+2, and only edges with transition probability greater than $0.95$ are retained.
+The largest connected component (a local region of which is shown here) remains sparse with average degree around 22 even as it grows from step $50$ to $800$.
+(a) Consequently, the distribution of solution path lengths, shown in both raw counts (histogram) and probability density (inset), shifts decisively to the right during this period, confirming that navigating the sparse backbone requires longer reasoning chains.
+Armed with this confirmation, we now demonstrate that the V-shaped evolution of the response length arises from a two-stage process.
+The decreases of the response length in the initial fast-learning stage is dominated by local optimization.
+Given the random initialization of transition probabilities, shorter paths inherently have a higher probability of being sampled.
+Consequently, the model rapidly discovers short, efficient solutions to individual problems, giving rise to a proliferation of disconnected “skill islands” (as shown in Fig.
 2(a)).
@@
 2(b)).
-Stage one thus constitutes a parallel search for local efficiency that naturally drives down the average response length.
-(a) Correct Response Histogram
-(b) Step 50
-(c) Step 800
-Figure 3: A Sparse Web Structure Necessitates Longer Reasoning Chains.
-This figure links the emergent sparse topology of the concept web to the observed increase in response length.
-(b, c) The largest connected component (edges with probability > 0.95) remains sparse (average degree $\approx 2$) even as it grows from step 50 to 800.
-(a) Consequently, the distribution of solution path lengths shifts decisively to the right during this period, confirming that navigating the sparse backbone requires longer reasoning chains.
+Stage one thus constitutes a parallel search for local efficiency that naturally shortens the average response length.
 The subsequent slow-learning phase, where response length steadily increases, marks a fundamental shift to global integration.
 The primary task is no longer discovering new islands but weaving them into a single, expansive concept web (see Fig.
-2(c)).
-As the largest connected component relentlessly grows, a critical structural property emerges: the web remains persistently sparse, with its average degree stabilizing near two (as shown in Fig.
+2(c)), defined to be the largest connected component in which every edge has a transition probability above $0.95$.
+As the concept web relentlessly grows, a critical structural property emerges: the web remains persistently sparse, with its average degree stabilizing near two (as shown in Fig.
 3(c)).
@@
 The histograms in Fig.
-3(a) provide dispositive proof of this effect: As the sparse web expands, the distribution of solution lengths undergoes a decisive and sustained rightward shift, confirming that navigation on this sparse backbone necessitates longer reasoning chains.
-Thus, the V-shaped curve furnishes a robust macroscopic signature of an underlying topological transition: from parallel, local optimization on disconnected components to serial, global navigation across an expanding, sparse backbone.
+3(a) provide clear evidence of this effect: As the sparse web expands, the distribution of solution lengths undergoes a decisive and sustained rightward shift, confirming that navigation on this sparse backbone necessitates longer reasoning chains.
+Thus, the V-shaped curve furnishes a robust macroscopic signature of an underlying topological shift: from parallel, local optimization on disconnected components to serial, global navigation across an expanding, sparse backbone.
 Yet can this simple, largely tree-like model capture the full complexity of LLM reasoning?
@@
 The existence of a limited number of such cycles does not significantly alter the global average degree of the web, which remains close to two.
-With the sparse web model established, this link between a macroscopic dynamic and a microscopic topology forms the foundation for the analysis that follows.
+With the sparse-web model established, this link between a macroscopic dynamic and a microscopic topology forms the foundation for the analysis that follows.
 (a) CoNet
@@
 Figure 4: The Fragility of a Sparse Web: Catastrophic Forgetting and Fast Recovery.
-This figure demonstrates a key prediction of our sparse web hypothesis: that the concept web is fragile, relying on critical bridge-like connections.
-(c-e) The microscopic view in CoNet shows how a lightweight supervised fine-tuning (SFT) on a converged web (c) severs these critical bridges, causing the structure to fragment (d).
+This figure demonstrates a key prediction of our sparse-web hypothesis: that the concept web is fragile, relying on critical bridge-like connections.
+(c-e) The color and shape conventions follow those of Fig.
+2, and only a local view of the concept web is shown.
+Orange crosses on the edges indicate decreases in transition weights, with thicker crosses denoting more substantial decreases.
+The microscopic view in CoNet shows how a lightweight supervised fine-tuning (SFT) on a converged web (c) severs these critical bridges, causing the structure to fragment (d).
 Subsequent RLVR rapidly repairs these links (e).
-(a, b) This microscopic severing manifests as macroscopic catastrophic forgetting in both CoNet and the 1.5B LLM, where performance collapses upon initiating SFT.
+(a, b) This microscopic severing manifests as macroscopic catastrophic forgetting in both CoNet and the 1.5B LLM, where performance plummets upon initiating SFT.
 The subsequent fast recovery once RLVR resumes highlights the localized nature of the damage, underscoring the web’s structural fragility.
@@
 From the graph theoretic point of view, the sparsity in the conceptual network implies low redundancy, fragility, and critical vulnerability.
-It also reflects a structural hypothesis for a well-known operational challenge: why models that have formed a concept web via RLVR are so susceptible to catastrophic forgetting when subsequently fine-tuned with (Li et al., 2024).
+It also reflects a structural hypothesis for a well-known operational challenge: why models that have formed a concept web via RLVR are so susceptible to catastrophic forgetting when subsequently fine-tuned with supervised learning (Li et al., 2024).
 Our theory posits that this is not a widespread erasure of knowledge, but a specific topological failure: the surgical severing of critical, bridge-like connections.
 To test this mechanistic hypothesis, we subjected models in the deep slow-learning phase to a lightweight in-distribution SFT (see Appendix H and I).
-As expected, the intervention triggered a precipitous collapse in overall performance in both CoNet and the 1.5B LLM (Fig.
+As expected, the intervention triggered a sharp drop in overall performance in both CoNet and the 1.5B LLM (Fig.
 44(a), 4(b)).
@@
 The microscopic view reveals that the SFT catastrophic forgetting was not a global degradation of the web.
-Instead, SFT’s aggressive optimization of a few paths overwrote the policy at critical branching nodes, surgically severing the single connection points for vast downstream subgraphs.
+Instead, SFT’s aggressive optimization of a few paths overwrote the policy at critical branching nodes, precisely breaking the single connection points for vast downstream subgraphs.
 Fig.
@@
 The subsequent RLVR phase then acts as a controlled “cooling”, guiding the system as it re-forms connections, ideally settling into a more robust configuration (Lai et al., 2025).
-This insight—that a known vulnerability can be repurposed as a deliberate annealing step—forms the foundation for the Annealed-RLVR algorithm, which we introduce in Sec.5.
+This insight—that a known vulnerability can be repurposed as a deliberate annealing step—forms the foundation for the Annealed-RLVR algorithm, which we introduce in Sec.
+5.
 4 Microscopic Mechanisms of RLVR: Forgetting by Frustration, Learning by Phase Transition
-Our sparse web theory reveals that the smooth, aggregate reward curve of RLVR training masks a dramatic microscopic behavior.
-Individual problem trajectories in Fig.
-5 reveal a fundamental tension born from the network’s sparse topology: a subtractive process of frustration-induced forgetting, peaking at the onset of slow learning, and an opposing, additive process of phase-transition-like learning that drives knowledge integration throughout the slow-learning stage.
+Although the overall reward curve in RLVR training looks smooth, our sparse-web theory shows that much more is happening underneath.
+As shown in the individual trajectories in Fig.
+5, the network’s sparse structure creates two competing effects: (i) a forgetting process caused by frustration, which is strongest at the start of the slow-learning phase, and (ii) an opposite, phase-transition-like learning process that gradually builds and integrates knowledge during slow learning.
 (a) CoNet
@@
 Figure 5: Microscopic Mechanisms of RLVR: Forgetting by Frustration, Learning by Phase Transition The learning trajectories for individual problems in CoNet (a) and the 1.5B LLM (b) reveal a fundamental duality.
-(i) Frustration-Induced Forgetting: At the transition to slow learning (see inset in (a)), the intense competition for connections on a sparse web manifests as volatile, non-monotonic accuracy curves, where some skills are competitively suppressed.
+(i) Frustration-Induced Forgetting: At the onset of slow learning [see inset in (a)], the intense competition for connections on a sparse web manifests as volatile, non-monotonic accuracy curves, where some skills are competitively suppressed.
 (ii) Phase-Transition-Like Learning: Subsequently, the web’s sparse frontier enables new skills to be acquired in punctuated, accelerated jumps (orange and green).
 The smoother gradients in the LLM (vs.
-the clean steps in CoNet) can be attributed to a classic finite-size effect, as explained in the main text.
+the clean steps in CoNet) can be attributed to finite-size effects, as explained in the main text.
 4.1 Frustration-Induced Forgetting: The Price of Integration
@@
 This process manifests as volatile, non-monotonic learning curves in the learning dynamics (Fig.
-5) and represents an internal “push-and-pull” fundamentally different from the external shock of SFT-induced catastrophic forgetting (Shenfeld et al., 2025; Jin et al., 2025).
-This process peaks in the maximally frustrated state at the onset of the slow-learning stage, a critical juncture where skill islands compete for limited connections under the $\left<k\right>\approx 2$ constraint.
+5) and represents an internal “push-and-pull” fundamentally different from the external shock of SFT-induced catastrophic forgetting (Shenfeld et al., 2025; Jin et al., 2025a).
+This process peaks in the maximally frustrated state at the onset of the slow-learning stage, a critical juncture where “skill islands” compete for limited connections under the $\left<k\right>\approx 2$ constraint.
 Paradoxically, this state maximizes exploration: the model’s ability to find diverse, correct out-of-distribution solutions peaks, far exceeding the base model or the same model after prolonged training (Fig.
@@
 5(b), more details in Appendix G).
-This phenomenon is a multi-task extension of the "Learning at Criticality" (LaC) mechanism, first identified in a single-task context (Cai et al., 2025).
+This phenomenon is a multi-task extension of the “Learning at Criticality” (LaC) mechanism, first identified in a single-task context (Cai et al., 2025).
 The LaC framework models this process as a continuous phase transition (Stanley & Ahlers, 1972; Wilson, 1983) where the policy collapses from exploring diverse reasoning paths to exploiting a single optimal one.
 Crucially, analysis revealed that near this critical point, the system’s exploratory power is maximized, and the reasoning path lengths become scale-invariant—a hallmark of continuous phase transitions.
+Beyond this transition, the policy collapses onto fixed optimal paths; the associated generalization drop is possibly linked to overfitting within the RLVR process (Dong et al., 2025; Gui & Cheng, 2025; Shypula et al., 2025; Zheng et al., 2025).
 This context explains the smoother gradient in the LLM.
-Compared to the step-like jumps in CoNet, the LLM’s learning curve is a classic manifestation of a finite-size effect(Fisher & Barber, 1972): it has a more constrained set of distinct reasoning paths for any given task, which rounds out the idealized sharp transition into a steep but continuous curve.
+Compared to the step-like jumps in CoNet, the LLM’s learning curve is a classic manifestation of finite-size effects (Fisher & Barber, 1972): it has a more constrained set of distinct reasoning paths for any given task, which rounds out the idealized sharp transition into a steep but continuous curve.
 The puzzle is why this delicate phenomenon survives in a chaotic multi-task setting.
 Our sparse-web model explains this through topological isolation: a node on the expanding frontier of a $\left<k\right>\approx 2$ network is statistically isolated, decoupling it from system-wide competition.
-Shielded from interference, a new skill can be efficiently “bolted on” via a clean phase transition.
+Shielded from interference, a new skill can be efficiently integrated via a clean phase transition.
 This perspective also provides a microscopic explanation for aggregate policy collapse in the standard RLVR (Chen et al., 2025; Yue et al., 2025; Cui et al., 2025).
 We hypothesize this is not a monolithic failure but the statistical sum of asynchronous, microscopic collapses at the frontier, a hypothesis supported by the LaC finding that training just beyond criticality risks overfitting-induced collapse.
-5 Annealed-RLVR: An Algorithm Born from the Sparse Web Theory
-(a) Randomly Selected 512 Training Problems
+5 Annealed-RLVR: An Algorithm Born from the Sparse-Web Theory
+(a) Training Sample (512 Problems)
 (b) Minerva
+(c) AIME 2024/2025
 Figure 6: Annealed-RLVR Outperforms Standard RLVR.
-This figure compares our Annealed-RLVR (“Annealed”) with a standard RLVR baseline (“RLVR”) on (a) in-distribution problems and (b) the out-of-distribution Minerva dataset.
-(Right panels) The best@k accuracy curves show that Annealed-RLVR achieves superior performance in both settings, confirming improved generalization.
-(Left panels) The histograms reveal the underlying mechanism: the annealing intervention systematically reduces the population of unsolved problems (low-accuracy bins) and improves the mastered ones (high-accuracy bins).
+This figure compares the best@k accuracy curves of our Annealed-RLVR ("Annealed") with the standard RLVR baseline ("RLVR").
+The evaluation is performed on (a) an in-distribution set (Randomly Selected 512 Training Problems) and two out-of-distribution (OOD) sets: (b) the Minerva dataset and (c) the AIME 2024/2025 datasets.
+The results show that Annealed-RLVR consistently achieves superior performance across all in-distribution and OOD benchmarks, confirming that the intervention fosters improved generalization.
 Our sparse-web theory reveals a core narrative of RLVR training: the system drives itself into a maximally frustrated state that is both a perilous bottleneck and a moment of peak exploratory potential (Fig.
@@
 This insight motivates Annealed-RLVR, a theory-driven intervention that leverages the high-entropy frustrated state to guide the model toward a more globally optimal configuration (Cheng et al., 2025; Chen et al., 2025).
-The algorithm operationalizes this concept with a precisely timed intervention, triggered by macroscopic signatures of the maximally frustrated state.
+The algorithm operationalizes this concept with a precisely timed intervention.
+This intervention is triggered by monitoring the macroscopic signatures of the maximally frustrated state, which, as shown in Fig.
+1, correspond to the crossover regime characterized by the “knee” of the reward curve (when accuracy gains plateau) and the bottom of the V-shaped response length curve.
 At this juncture, a brief SFT “heating” phase begins.
@@
 The SFT thus functions not as a disruptive overwrite but as a corrective jolt to break the policy’s premature commitment to suboptimal routes.
-We implement this by surgically applying SFT only to problems with low accuracy (e.g., < 0.1) for which correct solutions exist.
+We implement this by surgically applying SFT only to problems with low accuracy (e.g., <0.1<0.1) for which correct solutions exist.
 Immediately following this targeted heating, the “cooling” stage resumes standard RLVR, allowing the now more pliable policy to re-settle into a more robust and integrated final state.
-The results validate this theory-driven approach (Fig.
-6).
-Best@k accuracy curves demonstrate that Annealed-RLVR outperforms standard RLVR on both in-distribution problems and, crucially, the out-of-distribution Minerva dataset, confirming the intervention fosters true generalization.
-The histograms reveal the underlying mechanism: the SFT “heating” provides a crucial performance boost to the most difficult problems, lifting their success rates from near-zero and making them “visible” to subsequent RLVR exploration.
+We validated this approach on the DeepSeek-R1-Distill-Qwen-1.5B model, trained on the DeepScaleR-Preview-Dataset .
+Our method, Annealed-RLVR, was compared against the standard RLVR baseline where the policy is optimized using a variant of the Group Relative Policy Optimization (GRPO) algorithm (Shao et al., 2024; Liu et al., 2025; Yu et al., 2025).
+The results (Fig.
+6) confirm our theory.
+Best@k accuracy curves demonstrate that Annealed-RLVR outperforms standard RLVR on both in-distribution evaluation set (512 randomly selected training problems) and on out-of-distribution (OOD) benchmarks, including the Minerva and AIME 2024/2025 dataset.
+The histograms (see Fig.
+13) reveal the underlying mechanism: the SFT “heating” provides a crucial performance boost to the most difficult problems, lifting their success rates from near-zero and making them “visible” to subsequent RLVR exploration.
 The RLVR “cooling” phase then efficiently guides these newly accessible problems toward mastery.
@@
 We find that the degree-2 constraint first forces the system into a maximally frustrated state at the onset of slow learning, a phase dominated by frustration-induced forgetting as skills are competitively suppressed.
-Subsequently, the same sparse structure enables a cleaner growth mechanism: phase-transition-like learning, where new knowledge is efficiently “bolted on” at the web’s expanding frontier.
+Subsequently, the same sparse structure enables a cleaner growth mechanism: phase-transition-like learning, where new knowledge is efficiently integrated at the web’s expanding frontier , yet this very per-skill collapse onto a single high-reward path also accumulates over time into a macroscopic policy collapse and a global loss of solution diversity when training proceeds too far beyond criticality.
 This framework is not just descriptive; it is prescriptive.
@@
 See Appendix B for extended related work.
-8 ACKNOWLEDGMENTS
-The authors thank Haijun Zhou, Lei Wang, Hong Zhao and Ganqu Cui for inspiring discussions.
-K.
-C.
-and X.
-C.
-are supported by the National Key Research and Development Program of China under Grant No.
-2024YFA1408604 and the National Natural Science Foundation of China under Grants No.
-12047503 and No.
-12447103.
-P.
-Z.
-is supported by Project 12325501 of the National Natural Science Foundation of China.
-Y.
-D.
-and S.
-H.
-are supported by the National Natural Science Foundation of China under Grants No.
-12275263, the Innovation Program for Quantum Science and Technology under Grant No.
-2021ZD0301900, and the Natural Science Foundation of Fujian Province of China under Grant No.
-2023J02032.
-Z.
-Y.
-is supported by the National Natural Science Foundation of China under Grants No.
-12304288 and No.
-12247101.
 8 Reproducibility Statement
@@
 The dual-process theory, distinguishing between fast, intuitive System-1 thinking and slow, deliberate System-2 reasoning (Kahneman, 2011), has become a valuable framework for analyzing LLM behavior.
-While standard LLMs often exhibit System-1-like responses, significant research has focused on eliciting more robust, System-2 deliberation.
+While standard LLMs often exhibit System-1-like responses, significant research has focused on eliciting deliberate System-2 reasoning.
 Prompting strategies such as Chain-of-Thought (CoT) (Wei et al., 2022) and its variants like Self-Consistency (Wang et al., 2022) and Tree of Thoughts (ToT) (Yao et al., 2023) encourage models to generate explicit reasoning steps.
@@
 Complex Network Analysis of Neural Models.
-Applying network science to understand neural networks is a long-standing idea, previously used to study the topology and robustness of smaller models (Estrada & Hatano, 2008).
-With the advent of Transformers, researchers have begun analyzing attention patterns as graphs (Vig, 2019; Ferrando et al., 2023) to better understand information flow.
-More recently, studies have explored how network structures emerge during supervised learning, identifying the formation of essential subnetworks (Li et al., 2023).
-A recent study by (Minegishi et al., 2025) shares our goal of analyzing the network structures that underpin System-2 reasoning, though they do so by empirically extracting a graph from hidden-state representations.
+Applying network science to understand neural networks is a long-standing idea, previously used to study the topology and robustness of smaller models (Simard et al., 2005; Estrada & Hatano, 2008).
+With the advent of Transformers, researchers have begun analyzing attention patterns as graphs (Vig, 2019) to better understand information flow (Abnar & Zuidema, 2020; Chefer et al., 2021).
+More recently, studies have explored the topological properties of neural networks, identifying optimal graph structures (You et al., 2020) and the formation of essential sparse subnetworks (Frankle & Carbin, 2019; Filan et al., 2021).
+Complementing these static analyses, recent work has modeled the training process itself as the evolution of complex networks (Danovski et al., 2024; Rocchetti, 2025).
+A recent study by (Wang et al., 2024; Minegishi et al., 2025; Matsutani et al., 2025) shares our goal of analyzing the network structures that underpin System-2 reasoning, though they do so by empirically extracting a graph from hidden-state representations.
 In contrast, our CoNet framework utilizes a fixed, abstract concept space, which allows us to specifically isolate and analyze the dynamic mechanism of how this reasoning structure emerges during training.
@@
 “tens”) represent transition weights on an underlying reasoning graph, a view supported by convergent findings on the role of high-entropy tokens.
+This figure is adapted from Ref.
+Cai et al.
+(2025).
 Recent work, notably by (Minegishi et al., 2025), has developed detailed methods to empirically construct these task-specific reasoning graphs.
@@
 By analyzing graphs constructed this way, they found that more advanced models exhibit distinct topological properties, such as larger diameters, which correlate with improved performance.
-While analyzing these static, task-specific graphs is insightful, it leads to a deeper question: what is the underlying, universal knowledge structure that allows an LLM to generate so many different reasoning graphs across a multitude of tasks?
+However, the cluster centroids derived from K-means offer limited interpretability as stable reasoning concepts, making their correspondence to the model’s internal processes unclear.
+Moreover, the induced graph’s structure is unstable, varying significantly with the choice of model, dataset, segmentation, and representation layer.
+Hence, this graph cannot track evolving representational dynamics during RLVR training process.
+How to construct a principled model of an LLM’s internal reasoning process remains a formidable open problem.
+While analyzing these static, task-specific graphs is useful, it leads to a deeper question: what is the underlying, universal knowledge structure that allows an LLM to generate so many different reasoning graphs across a multitude of tasks?
 It follows that these empirically observed reasoning graphs are best understood as small projections of a much vaster, latent concept web.
@@
 This creates a tractable “computational microscope” to precisely analyze how the transition policy evolves under reinforcement, isolating the core dynamics of reasoning path formation.
+This figure is adapted from Ref.
+Cai et al.
+(2025).
 However, attempting to reverse-engineer this universal concept web from its sparse, task-specific projections presents formidable obstacles.
-The difficulties are multi-fold.
-First is a problem of scale: the true concept web is of astronomical size, and any extracted graph is an infinitesimal sample.
-More fundamentally is the ambiguity of the concept mapping: the mapping from a low-level token sequence to a high-level semantic concept is inherently many-to-one.
-Finally, this already-ambiguous mapping is itself dynamic, as it can shift during RLVR training.
+The difficulties are multifold.
+First, there is an issue of scale: the true concept web is of astronomical size, and any extracted graph is a small sample.
+Second, the mapping itself is ambiguous: the mapping from a low-level token sequence to a high-level semantic concept is inherently many-to-one.
+Third, this already-ambiguous mapping is itself dynamic, as it can shift during RLVR training.
 These challenges collectively lead to the observable node stability problem in practice, making it nearly impossible with current methods to track the web’s true topological evolution.
@@
 By positing a fixed space of concept nodes (a static graph structure) and isolating all learning dynamics to the evolving policy that navigates it, CoNet allows us to sidestep the challenges of direct graph extraction and precisely measure how the reinforcement signal sculpts reasoning paths.
-The model represents the LLM’s concept space as an abstract K-regular random graph, with the reasoning process simplified to a policy-guided Markovian random walk, as shown in Fig.8.
+The model represents the LLM’s concept space as an abstract K-regular random graph, with the reasoning process simplified to a policy-guided Markovian random walk, as shown in Fig.
+8.
 The transition policy, $\pi_{\theta}(j|i)\propto\theta_{ij}$, is governed by a set of learnable parameters $\theta_{ij}$ that represent the strength of the connection from node i to node j.
@@
 This correspondence validates our use of CoNet as a minimal model for exploring the network-level mechanisms of frustration and web formation.
-CoNet was originally introduced to model the phenomenon of Learning at Criticality (LaC), which was identified in single-task RLVR training (Cai et al., 2025).
-The central finding of LaC is that for a model trained on extremely sparse data, generalization performance peaks precisely at a sharp learning transition.
-The key methodological advance of this paper is to deploy this model in a richer, multi-task setting.
-This extension is crucial: it transforms CoNet from a model of isolated skill acquisition into an ideal theoretical laboratory for studying the emergent phenomena of skill integration, frustration, and concept web formation.
-As Fig.
-1 shows, this multi-task CoNet stunningly reproduces the key macroscopic signatures of the 1.5B LLM, validating its use for exploring the network-level mechanisms that drive the emergence of reasoning.
 Appendix E CoNet Implementation Details
@@
 The underlying concept space was instantiated as a directed random regular graph with $N=800$ nodes, where each node has a uniform out-degree of $k=40$.
-In this graph, we established a multi-task learning environment consisting of $128$ randomly generated question-answer (Q-A) node pairs.
-The initial transition parameters $\theta_{ij}$ for all edges were assigned randomly.
+In this graph, we established a multi-task learning environment consisting of $128$ randomly generated Q-A node pairs.
+The transition parameters $\theta_{ij}$ for all edges were initialized randomly.
 The use of a multi-task setting with 128 concurrent Q-A pairs is the fundamental mechanism that induces structural competition.
@@
 $\Delta\theta_{ij}\propto\sum_{m}A_{m}\nabla_{\theta_{ij}}\log\pi_{\theta}(j|i)$
-This advantage-based update functions as a competitive reinforcement mechanism.
-Only paths performing better than the current average are strengthened, while underperforming paths are implicitly suppressed.
+This advantage-based update functions as a competitive reinforcement mechanism, strengthening above-average paths and suppressing underperforming ones.
 This “winner-take-more” dynamic acts as an algorithmic sculpting tool, systematically pruning the vast majority of the initial $k=40$ edges from each node to carve out the sparse, efficient backbone of the concept web.
@@
 The condition is essential: if $N_{QA}$ is too small, there is no systemic competition to drive the formation of a unified web; if it is too large, the conflicting reinforcement signals prevent any coherent structure from forming.
-The robust emergence of these LLM-like signatures in this intermediate regime confirms that our CoNet setup successfully isolates the core principles of structural self-organization under multi-task reinforcement.
+The robust emergence of these LLM-like signatures in this intermediate regime confirms that our CoNet setup successfully reproduces the essential dynamics of structural self-organization under multi-task reinforcement.
 Appendix F Structural Evolution of the Concept Web in CoNet
-This section provides direct, quantitative evidence from the CoNet model for the two-stage learning dynamic—a fast acquisition phase followed by a slow integration phase—that underlies our theory.
+This section provides direct, quantitative evidence from the CoNet model for the two-stage learning dynamic—a fast acquisition process followed by a slow integration phase—that underlies our theory.
 Fig.
@@
 Figure 9: The Formation of the Concept Web in CoNet.
-This figure illustrates the evolution of network clusters during training, showing a structural transition from isolated skills to a unified conceptual web.
+This figure illustrates the evolution of network clusters during training, showing a structural reorganization from isolated skills to a unified conceptual web.
 In the initial phase, the Cluster Number (orange curve) spikes, indicating the rapid discovery of numerous disconnected “skill islands”.
-The peak marks a critical transition point.
+The peak marks a critical structural juncture.
 Following this peak, the Cluster Number steadily declines as these islands begin to coalesce.
@@
 During this phase, the system operates in a greedy, parallel fashion, finding many efficient, local solutions without regard for their interconnection.
-The transition to the slow-learning phase occurs at the critical juncture marked by the peak of the orange curve around step 20.
+The shift to the slow-learning phase occurs at the critical juncture marked by the peak of the orange curve around step 20.
 At this point, the system has exhausted the “low-hanging fruit” of simple, isolated solutions and has reached a state of maximal fragmentation.
@@
 (Left) Accuracy trajectories for these problems.
-All problems shown exhibit a sharp, sigmoidal-like increase in accuracy late in the training process, characteristic of a critical learning event.
+All problems shown exhibit a sharp, sigmoidal-like increase in accuracy rate in the training process, characteristic of a critical learning event.
 (Right) Corresponding variance of the solution path length.
@@
 H.2 Expeimental Results
+Figure 11: Structural Impact of Annealed-RLVR on the Concept Web.
+Evolution of concept web for standard RLVR (blue) versus Annealed-RLVR (orange).
+The SFT intervention at step 50 (dashed line) induces an immediate drop in the Annealed model’s cluster size.
+Subsequently, the Annealed model recovers and surpasses the standard RLVR baseline, which exhibits slower growth, ultimately forming a larger final concept web.
 Figure 12: Validation of Annealed-RLVR in CoNet.
@@
 The final annealed policy (“Annealed”) ultimately achieves superior performance over the baseline (“RLVR”).
-The algorithm consists of a strategically timed Supervised Fine-Tuning (SFT) phase, followed by a resumption of standard Reinforcement Learning from Verifiable Rewards (RLVR) training.
-The SFT intervention is applied at training step 50, the empirically determined point of maximum frustration where the system transitions from forming disconnected “skill islands” to integrating them.
+The algorithm consists of a strategically timed SFT phase, followed by a resumption of standard RLVR training.
+The SFT intervention is applied at training step 50, the empirically determined state of maximum frustration, marking the crossover from the formation of disconnected “skill islands” to their subsequent integration.
 At this stage, we first identify all problems with an accuracy below 0.1, select a total of 50 low-accuracy problems, and sample a known successful reasoning path for each.
@@
 Following this brief SFT, standard RLVR training is resumed, which functions as the “cooling” phase, allowing the now more exploratory policy to settle into a new, more globally optimal configuration.
-The success of this SFT-RLVR cycle is confirmed by the results in Fig.
-11.
+Fig.
+11 provides the direct structural evidence for this annealing process.
+As shown by the orange “Annealed” curve, the SFT intervention at step 50 locally disrupted network connections, resulting in a decline in the concept web size.
+However, this disruption is precisely what allows the system to overcome the integration bottleneck.
+While the standard RLVR model (blue curve) is shown to grow slowly, trapped in the maximally frustrated state, the Annealed model’s cluster size rapidly recovers and grows to a larger final scale.
+This demonstrates that the SFT intervention, by resolving the competitive frustration, enables the subsequent RLVR “cooling” phase to successfully expand the network and integrate a greater number of “skill islands”.
+The success of this SFT-RLVR cycle is also confirmed by the results in Fig.
+12.
 The line plot on the right shows the immediate effect of the intervention: best@1 accuracy (a proxy for exploitation) temporarily dips, while best@16 accuracy (a proxy for exploration) surges.
@@
 The histogram on the left reveals the underlying mechanism for this improvement.
-The SFT-RLVR cycle successfully resolves the frustration bottleneck by dramatically reducing the population of completely unsolved problems (Accuracy = 0.0) and correspondingly increasing the number of fully mastered problems (Accuracy = 1.0) compared to the standard model.
+The SFT-RLVR cycle successfully resolves the frustration bottleneck by reducing the population of completely unsolved problems (Accuracy = 0.0) and correspondingly increasing the number of fully mastered problems (Accuracy = 1.0) compared to the standard model.
 In summary, the CoNet experiment provides strong evidence for our theory.
@@
 I.1 Base Model and GRPO Setup
-We used the DeepSeek-R1-Distill-Qwen-1.5B model (Guo et al., 2025), trained on the DeepScaleR-Preview-Dataset (Luo et al., 2025) with their modified version of the veRL open-source engine (Sheng et al., 2024; ByteDance Seed Team and verl community, ).
-Our Generative Reward Policy Optimization (GRPO) hyperparameters followed the first-stage settings, which specified a response length of 8192, as outlined in the DeepScaleR paper (Luo et al., 2025).
+We used the DeepSeek-R1-Distill-Qwen-1.5B model (Guo et al., 2025), trained on the DeepScaleR-Preview-Dataset (Luo et al., 2025a) with their modified version of the veRL open-source engine (Sheng et al., 2024; ByteDance Seed Team and verl community, ).
+Our Group Relative Policy Optimization (GRPO) hyperparameters followed the first-stage settings, which specified a response length of 8192, as outlined in the DeepScaleR paper (Luo et al., 2025a).
 I.2 SFT Dataset for Annealed-RLVR
-The Supervised Fine-Tuning (SFT) dataset for the annealed-RLVR experiment (labeled “Annealed” in Fig.
+The SFT dataset for the annealed-RLVR experiment (labeled “Annealed” in Fig.
 6) was constructed to focus on difficult problems.
@@
 This SFT dataset was used to fine-tune the model for two epochs with a learning rate of $5\times 10^{-5}$ before resuming GRPO training.
+I.4 Computational Cost Analysis
+The overall computational complexity of Annealed-RLVR is dominated by the iterative GRPO training, making it nearly identical to the standard RLVR baseline.
+The only addition is a one-time cost for the SFT intervention, which is applied once at the point of maximal frustration (around step 100 in our case).
+This one-time cost consists of two phases:
+  1. 1.
+    Evaluation Pass: An evaluation pass over the $N$ problems in the training set (where $N\approx 40,000$) is required to identify the low-accuracy problems.
+As described in Section I.2, this involved generating 8 to 60 rollouts per problem.
+This cost can be formally described as $O(N\cdot k)$ for a fixed $k$ rollouts, or $O(N\cdot\log(k))$ if using an adaptive, binary-like rollout strategy to find problems below the accuracy threshold.
+  2. 2.
+    SFT Phase: The SFT itself is performed on the small subset of resulting correct trajectories.
+Letting $\eta$ be the accuracy threshold parameter (e.g., $\eta=0.1$ for our 10% threshold), the SFT cost is proportional to $O(\eta N)$.
+As detailed in Section I.2, this amounted to only 3,933 trajectories (approximately $\eta N$) in our experiment.
+This $O(\eta N)$ SFT cost is significantly smaller than the evaluation cost.
+Both of these marginal, one-time costs are much smaller than the total iterative cost of the full GRPO training run (e.g., 1,700 steps).
 Appendix J LLM Evaluation Details
@@
 6 illustrates the full performance curves, while Table 1 highlights the results from representative checkpoints.
+(a) Training Sample (512 Problems)
+(b) Minerva
+(c) AIME 2024/2025
+Figure 13: Mechanism of Annealed-RLVR: Accuracy Distributions.
+This figure complements the best@k curves in the main text by revealing the underlying mechanism of performance improvement.
+It shows the per-problem accuracy histograms for the checkpoints of Annealed-RLVR ("Annealed") versus the standard RLVR ("RLVR") baseline, evaluated on (a) the in-distribution set (Random 512), (b) the OOD Minerva dataset, and (c) the OOD AIME 2024/2025 datasets.
+The histograms demonstrate that the Annealed-RLVR intervention systematically reduces the population of unsolved problems (low-accuracy bins) and increases the population of mastered problems (high-accuracy bins).
 Table 1: Best@k Accuracy (%) with Standard Error at Key Training Checkpoints
@@
 Annealed 68.2(5) 68.3(4) 68.4(5) 68.0(4) 69.4(5) 67.8(5)
+Dataset: AIME 2024 & 2025
+best@1 GRPO 20.11(0.10) 23.08(0.11) 24.04(0.12) 26.98(0.12) 27.77(0.11) 28.11(0.11) 28.10(0.11)
+Annealed 23.27(0.11 23.86(0.12 26.77(0.12 27.67(0.11 28.26(0.11 27.93(0.11
+best@4 GRPO 33.31(0.17) 38.44(0.19) 40.33(0.19) 42.74(0.18) 43.36(0.19) 43.51(0.18) 43.88(0.18)
+Annealed 39.36(0.19) 40.21(0.19) 43.38(0.19) 43.76(0.19) 43.79(0.19) 43.42(0.19)
+best@16 GRPO 45.3(0.3) 52.4(0.3) 54.4(0.3) 55.5(0.3) 56.9(0.3) 56.1(0.3) 57.1(0.3)
+Annealed 54.1(0.3) 54.7(0.3) 57.6(0.3) 57.7(0.3) 58.1(0.3) 57.8(0.3)
+best@64 GRPO 55.9(0.5) 62.4(0.5) 64.2(0.5) 64.5(0.6) 65.9(0.6) 64.8(0.6) 66.2(0.6)
+Annealed 63.8(0.5) 64.2(0.5) 67.0(0.5) 66.7(0.5) 67.5(0.5) 67.7(0.5)

### v3

@@
-How LLMs Learn to Reason: A Complex Network Perspective
+Emergent Slow Thinking in LLMs as Inverse Tree Freezing
 Abstract
-Training large language models with Reinforcement Learning with Verifiable Rewards (RLVR) exhibits a set of distinctive and puzzling behaviors that remain poorly understood, including a two-stage learning curve, a V-shaped response-length trajectory, and a pronounced vulnerability to catastrophic forgetting.
-In this work, we propose that these behaviors are emergent collective phenomena governed not by neural implementation details, but by the topological evolution of the latent reasoning graph in semantic space.
-By demonstrating a dynamical isomorphism between a 1.5B-parameter LLM and a minimal Concept Network Model (CoNet), we trace the causal source to the self-organization of a sparse concept web pinned to an average degree of two.
-This geometric perspective provides a unified physical explanation for the observed anomalies: the V-shaped trajectory tracks the evolution from parallel local skill optimization to global network integration; catastrophic forgetting stems from the topological disconnection of critical “trunk” edges; and policy collapse arises from the accumulation of sequential transitions at the web’s leaf nodes, where broad exploration abruptly freezes into rigid, high-reward trajectories.
-Identifying a “maximally frustrated state” at the transition between learning stages, we propose Annealed-RLVR, a principled algorithm that injects a targeted SFT “heating” step to resolve this topological bottleneck.
-Experiments confirm that this theory-driven intervention outperforms standard RLVR on both in-distribution and out-of-distribution benchmarks (including Minerva and AIME).
-By recasting RLVR from black-box optimization into a predictable process of structural self-organization, our work provides a new physical intuition for engineering the emergent reasoning capabilities of future AI systems.
+Reinforcement learning with verifiable rewards (RLVR) enables large language models to acquire slow, multi-step reasoning from sparse final-answer signals.
+We provide a statistical-physics picture of this emergence.
+We show that an autoregressive model’s finite capacity forces it to compress its exponentially large prefix space into a Markov network of predictive states, on which slow thinking unfolds as a random walk—the Concept Network (CoNet) picture.
+Within CoNet, RLVR dynamics are governed by two mechanisms: merging of compatible paths and frustrated competition among incompatible ones.
+Together they drive the network through nucleation, growth, and freezing into multi-input, single-output directed inverse trees.
+The picture reproduces the training dynamics of a 1.5-billion-parameter LLM and yields three predictions: reasoning chains lengthen as a geometric necessity of sparse topology; SFT induces catastrophic forgetting through bridge-node rupture; and frustration drives policy collapse.
+Building on the structural timing inherent in inverse-tree freezing, we propose Annealed-RLVR—a brief SFT intervention at the moment of maximum frustration.
+It outperforms standard RLVR on both in- and out-of-distribution benchmarks, with the largest gains at high sampling budgets where standard RLVR collapses.
+The same SFT applied after the trees freeze instead triggers catastrophic forgetting, isolating timing as the active ingredient.
 1 Introduction
-Training large language models with reinforcement learning with verifiable rewards (RLVR) (Liu et al., 2025; Guo et al., 2025; Team et al., 2025; Yang et al., 2025) has revealed four striking and still poorly understood phenomena.
-On math and coding benchmarks, RLVR runs typically show (i) rapid initial gains followed by a long performance plateau, even as training continues (Luo et al., 2025a; He et al., 2025; Jin et al., 2025b); (ii) a V-shaped trajectory of the length of correct solutions, where answers first become markedly shorter before later growing longer again (Luo et al., 2025a; He et al., 2025; Fatemi et al., 2025; Wang et al., 2025); (iii) pronounced vulnerability to catastrophic forgetting when RLVR-trained policies are interleaved or followed by supervised fine-tuning (SFT), with earlier reasoning skills abruptly degraded (Li & Hoiem, 2017; Luo et al., 2025b; Ding & Wang, 2025); and (iv) a collapse of solution diversity, with the policy converging onto a narrow and brittle solution manifold (He et al., 2025; Cui et al., 2025; Hao et al., 2025; Jiang et al., 2025).
-Existing work offers only partial, largely phenomenological explanations for each phenomenon in isolation: plateaus and the collapse of solution diversity are linked to entropy exhaustion and over-exploitation of a fixed reward signal (Cui et al., 2025; Hao et al., 2025; Jiang et al., 2025); the initial shortening of solutions is attributed to reward-driven pruning of redundant reasoning, whereas the subsequent lengthening has been ascribed either to a shift toward solving harder instances that demand longer chains of thought, or to the emergence of multi-pass, self-reflective reasoning (“aha”-like behavior) under RLVR(Fatemi et al., 2025; Wang et al., 2025); and catastrophic forgetting is analyzed as an objective-mismatch problem between RLVR and subsequent SFT or other post-training stages (Luo et al., 2025b; Ding & Wang, 2025).
-However, these accounts remain fragmented: we still lack a unifying framework that ties all four behaviors to a common underlying structure or dynamical mechanism of RLVR-trained policies, and explains why they so often arise together in practice.
-A growing body of work tries to explain these phenomena by modeling LLM’s reasoning process as an implicit reasoning graph—a network of semantic states and logical transitions (Wang et al., 2024; Dutta et al., 2024; Xiang et al., 2025; Cabannes et al., 2024; Minegishi et al., 2025).
-However, translating this idea into a fully microscopic analysis has proven extremely difficult: constructing the reasoning graph from a model’s high-dimensional latent space is a formidable open problem, preventing a direct investigation into the structural origins of the observed dynamics (See Appendix C for details).
-Here, we argue that the very universality of the above RLVR dynamics provides a path forward.
-In complex systems, collective emergent behaviors are largely independent of microscopic details and governed by the system’s large-scale organization (Anderson, 1972; Watts & Strogatz, 1998; Barabási & Albert, 1999).
-This insight motivates an approach rooted in the core logic of the Renormalization Group (RG) (Wilson & Kogut, 1974): instead of directly studying the system at the full microscopic level, one can model and approach the system at a coarse-grained level to explain its macroscopic behavior.
-Applying this strategy, we shift our focus from the full reasoning graph at the token level to its coarse-grained backbone at the semantic level, which emerges by abstracting away fine-grained skills encoded in local network patterns like motifs or community structures (Milo et al., 2002; Fortunato, 2010; Zhang & Moore, 2014).
-This leads us to our central hypothesis regarding the emergent structure that forms under RLVR training for general-purpose System-2 reasoning:
-Hypothesis 1.
-The concept web, defined as the coarse-grained reasoning graph, is a sparse network whose effective average degree is pinned to $\left<k\right>\approx 2$.
-The implications of this hypothesis are non-trivial: an average degree of $2$ implies a predominantly tree-like structure, which is efficient for generalization but inherently fragile.
-This topology can still support rare, large-scale cycles critical for complex operations like self-correction (as detailed in Appendix C), without significantly altering the global average degree.
+Cognitive science draws a sharp distinction between fast thinking (System 1) and slow thinking (System 2) (Kahneman, 2011; Yang et al., 2026): the former is rapid, automatic, pattern-driven intuition; the latter is slow, deliberate, step-by-step rational reasoning.
+A large language model (LLM)—an autoregressive sequence generator built on the Transformer architecture (Vaswani et al., 2017)—is, by construction, a System 1 machine (Hua & Zhang, 2022): given a token sequence $(a_{1},a_{2},\ldots,a_{t})$, the model uses its parameters $\theta$ to emit, in a single forward pass, the conditional probability $\pi_{\theta}(a_{t+1}\,|\,a_{1},\ldots,a_{t})$ of the next token; the sampled $a_{t+1}$ is appended to the sequence and generation continues, with no explicit search or planning at any point.
+A recent finding has overturned this picture: an extraordinarily simple training paradigm—reinforcement learning with verifiable rewards (RLVR) (Liu et al., 2025; Guo et al., 2025; Team et al., 2025; Yang et al., 2025)—suffices for LLMs to spontaneously acquire slow thinking (Yang et al., 2026).
+After the model produces a complete response, RLVR issues a sparse $\{0,1\}$ reward based solely on the correctness of the final answer, with no supervision of intermediate steps, and the parameters $\theta$ are then updated by policy gradient.
+From GPT-o1 (Jaech et al., 2024) to DeepSeek-R1 (Guo et al., 2025), RLVR-trained models spontaneously unfold reasoning chains hundreds of thousands of tokens long—decomposing problems step by step, checking intermediate results, correcting errors—reaching unprecedented levels in mathematical proof and program generation (Team et al., 2026; Zeng et al., 2026; Singh et al., 2025; DeepSeek-AI, 2026).
+The crux is that the model is never taught how to reason—it is told only whether the answer is right—and slow thinking emerges spontaneously from the fast-thinking substrate.
+From the standpoint of statistical physics, the emergence of slow thinking on top of a fast-thinking substrate carries the classical hallmark of an emergent phenomenon (Anderson, 1972; Cai et al., 2025): a separation of timescales, with microscopic degrees of freedom (individual tokens and their local statistics) unchanged yet a wholly new collective behaviour (multi-step logical reasoning) appearing at the macroscopic level.
+Wang et al.
+(Wang et al., 2025) provide direct empirical evidence for this separation through token-by-token information-entropy measurements (the model’s uncertainty over the next token) of generated sequences before and after RLVR training: about 80% of tokens lie within local computation chunks (multi-token phrases with near-deterministic transitions) where the entropy is very low (<0.5<0.5 nats) and the probability distribution barely changes between pre- and post-RLVR—these are the fast computation steps already mastered during pretraining; what RLVR genuinely reshapes is the remaining ${\sim}20\%$ of high-entropy tokens located at the transitions between chunks, which dictate the macroscopic trajectory of the reasoning chain—exactly where slow thinking takes place.
+Even more telling, updating only these 20% high-entropy tokens reproduces the performance gain of full training (Wang et al., 2025).
+Fast thinking (within chunks) stays put; slow thinking (the joints between chunks) emerges—a clean separation of timescales between the two.
+Not every system trained with reinforcement learning develops emergent reasoning.
+AlphaGo (Silver et al., 2016), trained over the vast search space of Go ($\sim\!10^{170}$ legal positions), never spontaneously acquired a search strategy of its own—search had to be hard-coded into the algorithm via Monte Carlo tree search (MCTS) (Coulom, 2006; Browne et al., 2012).
+The difference lies in two structural features that LLMs possess and AlphaGo does not.
+The first is a domain-agnostic action-sequence representation.
+AlphaGo’s policy takes the current board (a state representation) as input; the move history is discarded, and a search strategy cannot be expressed in the network’s input/output format and must be bolted on externally as MCTS.
+An LLM’s policy, by contrast, takes the full token-generation history $(a_{1},\ldots,a_{t})$ as input (an action-sequence representation).
+Under this representation, fast and slow thinking share a single format—both are token sequences differing only in length—and a reasoning strategy is itself a pattern expressible and learnable within that sequence, requiring no external search module.
+The second is self-reference: every output $a_{t+1}$ is immediately fed back as part of the input for the next step, contributing to the prediction of $a_{t+2}$.
+Self-reference equips the system with an additional axis along which it can scale—inference time.
+Beyond the parameter count $|\theta|$ and the size of training data, it opens a third scaling channel, namely test-time scaling (Jaech et al., 2024; Snell et al., 2024; Brown et al., 2024), allowing a policy in the unified action space to deepen itself through stepwise iteration.
+These two structural features explain why LLMs can develop emergent slow thinking—they possess a domain-general action space and a scalable temporal axis.
+They have not yet answered a more fundamental question: what does RLVR actually change at the microscopic level so that macroscopic reasoning can emerge?
+The first contribution of this work is to introduce, for Transformer-class autoregressive models, an effective Markov representation built on the idea of predictive state variables (PSVs) (Littman et al., 2001).
+Under this representation, the exponentially large token-sequence space is compressed onto a sparse network of effective states, and reasoning is mapped to a random walk on that network.
+This framework also resolves a practical puzzle: why does reasoning training in LLMs require a Group Relative Policy Optimization (GRPO) based RLVR (Shao et al., 2024) while Monte Carlo tree search (MCTS) is unable to do the job?
+The reason is that MCTS relies on an explicit state representation (such as a board position) to construct its search tree, whereas the effective states of an action-sequence representation only emerge implicitly through training and cannot be enumerated in advance.
+Building on this effective Markov representation, we propose that the training dynamics of RLVR can be mapped onto a freezing problem on a complex network.
+Multi-task RLVR with shared parameters merges paths and carves the high-probability edges into a sparse, multi-input, single-output directed tree—an inverse tree.
+The inverse tree forms spontaneously through processes analogous to nucleation and growth in a supercooled liquid (Avrami, 1939; Kelton & Greer, 2010).
+Almost every node on the tree has a single, near-deterministic outgoing edge, ensuring that even though reasoning may span many steps, each individual choice is nearly certain and unlikely to err—precisely the structural guarantee that allows long-range reasoning to proceed reliably.
+At the same time, the reasoning paths of different tasks compete on shared nodes, producing a frustration mechanism: when two reasoning trajectories share an intermediate node but demand different outgoing edges, competition at that node eventually freezes the dominant path into a deterministic transition while almost entirely erasing the subordinate one.
+This theoretical framework gives a unified account of several core phenomena in RLVR training: the lengthening of reasoning chains with growing capability (Luo et al., 2025; He et al., 2025) follows naturally from the geometry of a sparse tree (Inverse Tree Picture section); the catastrophic forgetting triggered by interleaving SFT into RL (Li & Hoiem, 2017; Luo et al., 2023; Ding & Wang, 2025) stems from the rupture of critical branching nodes on the inverse tree (Catastrophic Forgetting section); and a proper understanding of the forgetting-by-frustration mechanism (Two Core Mechanisms section) points to a favourable timing window for SFT intervention—the Annealed-RLVR scheme (Annealed-RLVR section), which suppresses policy collapse (Cui et al., 2025; Yue et al., 2025) while preserving high exploration capacity.
+2 Effective Markov Representation and State Compression
+The autoregressive generation of a Transformer is a textbook non-Markov process: the output at step $t$ depends on the full history $(a_{1},\ldots,a_{t})$ rather than on a finite-dimensional hidden state.
+This dependence on the entire history is usually counted among the strengths of Transformers over models with explicit state representations such as hidden Markov models—it grants greater expressive power.
+By the same token, however, the dynamics of an autoregressive model is hard to analyse theoretically: treating each distinct history as a separate state yields a state space of size $V^{t}$ ($V$ the vocabulary size, $t$ the sequence length), which explodes exponentially in $t$.
+The work of Littman, Sutton and Singh on predictive state representations (Littman et al., 2001) provides a key insight: for any observable dynamical system, no matter how intricate its internal mechanism, the state can be characterised equivalently by a set of predictions about future observables—the number of effective states is determined by the number of distinguishable future behaviours, not by the size of the history space.
+For an autoregressive model with $d=|\theta|$ parameters this means that the parameter space can encode at most $O(e^{d})$ distinguishable conditional distributions, so a great many distinct histories must necessarily map to identical future behaviour.
+Following this line of thought, we define the predictive state variable (PSV): given a prompt $Q$ and an action history $h$, two pairs $(Q,h)$ and $(Q^{\prime},h^{\prime})$ belong to the same effective state if and only if they induce future conditional distributions that are indistinguishable at coarse-graining scale $\epsilon$,
+$(Q,h)\sim_{\epsilon}(Q^{\prime},h^{\prime})\;\Longleftrightarrow\;\|\pi_{\theta}(\cdot\,|\,Q,h)-\pi_{\theta}(\cdot\,|\,Q^{\prime},h^{\prime})\|<\epsilon,$ (1)
+where $\pi_{\theta}(\cdot\,|\,Q,h)$ denotes not the single-step distribution but the joint conditional distribution over the entire remaining token sequence.
+The threshold $\epsilon$ has a natural lower bound set by the finite context window and finite vocabulary.
+By construction, members of the same PSV class $s=[(Q,h)]_{\sim}$ produce identical future distributions up to $\epsilon$.
+The next token therefore depends only on $s$, not on the specific history that reached it—the defining property of a Markov state.
+A crucial property follows: histories from different problems can be merged into a single equivalence class through compression.
+This cross-problem state sharing will become the microscopic basis of path merging and inverse-tree formation in later sections.
+We emphasise that the argument above establishes the existence of an effective Markov representation as a logical consequence of finite parametric capacity, but the actual partition of histories into PSV classes depends on the trained parameters and can only be constructed numerically, not derived analytically.
+Direct enumeration of equivalence classes in real LLMs (with $10^{9}$$10^{11}$ parameters) is currently infeasible.
+We therefore turn to a model system whose state space is exhaustively enumerable: the $2\times 3$ sliding puzzle ($N=120$ legal positions; the three actions $V$, $L$, $R$ are compound permutations that keep the blank in the centre column; Fig.
+1a), on which we train a minimal autoregressive Transformer with GRPO (experimental details in the supplementary material).
+The decisive advantage of this system is that the entire branching subtree below every decision point can be expanded explicitly, so the equivalence relation (1) can be checked directly and the full time evolution of the compression ratio can be tracked.
+In this experiment, a prompt $Q$ is a start state and a history $h$ is a sequence of actions $(V,L,R)$ output so far.
+Each prefix $(Q,h)$ constitutes a decision point at which the model must choose the next action.
+Most of the ${\sim}120\times 3^{12}$ possible decision points are never visited during training because the policy assigns them negligible probability; including them would inflate $R_{c}$ with spurious classes.
+We therefore restrict the count to $\mathcal{S}_{\text{active}}$, the set of decision points whose ancestral chain of action probabilities is $\geq 0.05$, and define the compression ratio $R_{c}=1-N_{c}/|\mathcal{S}_{\text{active}}|$, where $N_{c}$ is the number of PSV equivalence classes on $\mathcal{S}_{\text{active}}$.
+$R_{c}\to 1$ means that distinct histories are highly equivalent in their future behaviour and the effective state space is heavily compressed.
+The model is small enough (${\sim}14$K parameters) that GRPO converges reliably from a random initialization, without pretraining.
+This keeps the training dynamics purely RLVR-driven, free of any confound introduced by a pretraining phase.
+The temporal evolution of $R_{c}$ (Fig.
+1b) traces a non-monotonic V-shape.
+Before training, the randomly initialized policy is near-uniform at every decision point—all nodes share essentially the same future, almost all states fall into a single PSV class, and $R_{c}$ is close to 1.
+Stage 1 (the rapid-learning window) is when the model begins to assign distinct action preferences to distinct states; equivalence classes split apart and $R_{c}$ plunges—this is the inevitable price of forming meaningful structure.
+In Stage 2, $R_{c}$ rebounds and stabilises in a high-compression regime: reasoning paths of different problems start sharing intermediate nodes, and many distinct histories collapse back onto a small number of classes—unlike the trivial initial compression, this one reflects genuine policy convergence.
+Throughout Stage 2, $R_{c}$ barely changes while the success rate continues to climb slowly.
+Visual inspection of the PSV graph (Fig.
+1c) suggests that the topological skeleton is largely laid down by the end of Stage 1, with later gains arising from refinement within that already-established structure rather than from further coarse-scale reorganisation.
+Figure 1: Direct measurement of PSV compression during training of a small Transformer.
+(a) State-graph representation of the $2\times 3$ sliding puzzle: $N=120$ legal positions form a $K=3$ regular graph; shown is a shortest solution path of distance 4.
+(b) Time evolution of success rate (top) and compression ratio $R_{c}$ (bottom).
+$R_{c}$ traces a V-shape: an initial drop driven by policy diversification, followed by a rebound and saturation in a high-compression regime (65–91%) that signals the freezing-in of the topological skeleton.
+(c) PSV effective state-transition graph at early (left, epoch 10) and late (right, epoch 1000) training.
+The number of states contracts by roughly an order of magnitude, revealing a sparse multi-input, single-output inverse-tree structure.
+As a scalar, $R_{c}$ only quantifies the degree of compression and says nothing about the topology of the compressed state space.
+Building the effective state-transition graph—PSV equivalence classes as nodes and high-probability transitions as directed edges—makes the post-training topology directly visible (Fig.
+1c): the dense, disordered initial graph contracts into a sparse multi-input, few-output directed structure—many starting paths converge into a small number of shared hub nodes and then flow, along almost unique channels, to the target.
+A widely accepted view in LLM research holds that “compression is intelligence” (Delétang et al., 2023; Huang et al., 2024)—the generalization capability of a model arises fundamentally from effective compression of the training data.
+The state-space compression observed in our exhaustively enumerable system is therefore very likely also at work during the RLVR training of LLMs built on the same Transformer architecture.
+Since the state space of a real LLM cannot be directly enumerated, the next sections develop the question from both theoretical and experimental sides: how is this sparse topology sculpted by the training dynamics of RLVR, and can its formation account for the macroscopic phenomena observed in real LLM training—two-stage dynamics, growth of reasoning chains, and catastrophic forgetting?
+3 The Concept Network: A Physical Picture of Slow Thinking
+(a) CoNet
+(b) Minimal Transformer
+(c) DeepScaleR-1.5B
+Figure 2: The same two-stage dynamics span three scales of model.
+(a) The minimal CoNet implementation: no neural network, transition probabilities as trainable parameters.
+(b) A 14K-parameter single-layer Transformer trained on the sliding puzzle.
+(c) DeepScaleR-1.5B, a 1.5-billion-parameter LLM.
+Top: reward; bottom: path/token length of correct responses.
+All three exhibit the universal Stage 1 (fast learning, shortening length) $\to$ Stage 2 (slow optimization, rebound in length) trajectory.
+The PSV analysis of the previous section shows that, after RLVR training, the LLM admits an effective Markov representation.
+Within this representation the reasoning process is naturally a random walk on a sparse directed graph: each PSV equivalence class is a node, transition probabilities are determined implicitly by the model parameters, a single transition is one local computation (fast thinking), and a multi-step path from a question node to an answer node constitutes the full reasoning chain (slow thinking).
+This is the core of the Concept Network (CoNet) picture (Cai et al., 2025): slow thinking is fast thinking unfolded over time on a sparse network.
+This provides a different representation for RLVR from the reasoning-graph analyses discussed in (Wang et al., 2024; Minegishi et al., 2025).
+In a real Transformer, every transition probability is jointly determined by self-attention, multilayer perceptrons and other intricate components; at the level of the effective Markov representation, however, all such microscopic detail is compressed into a single transition probability $p(s^{\prime}\,|\,s)$ between nodes.
+After pretraining, before any task-specific optimization, a great many outgoing edges at every node carry non-zero transition probability and the effective state-transition graph is a dense, high-entropy network—a random walker faces too many possible branches at every node to find a reliable path to the answer.
+RLVR’s role is precisely to prune these transition probabilities systematically: through GRPO and other policy-gradient methods, edges along successful paths are reinforced and the rest are suppressed; the outgoing-edge entropy of every node decreases, so that the network converges from a high-entropy uniform distribution to a low-entropy structure dominated by a few edges.
+The minimal implementation of the CoNet picture treats the transition probabilities themselves as trainable parameters.
+We start from a $K$-regular random graph $G(N,K)$ (NN nodes, each emitting exactly $K$ directed edges); each directed edge $s\to s^{\prime}$ carries a trainable probability $p(s^{\prime}\,|\,s)$ with $\sum_{s^{\prime}}p(s^{\prime}\,|\,s)=1$; reasoning is a random walk that starts at the question node $Q$ and samples transitions according to $p(s^{\prime}\,|\,s)$, receiving reward 1 if it reaches the answer node $A$ and 0 otherwise; transition probabilities are updated by GRPO.
+This minimal implementation strips away every Transformer component—self-attention, MLPs, embeddings, positional encoding—and retains only the three essential ingredients of RLVR: path search on a sparse graph, shared transition probabilities, and a sparse reward at the end of the sequence.
+The value of this minimal implementation is full observability: every transition probability at every node can be logged in full throughout training, and the subgraph of high-probability edges can be visualised directly.
+If the CoNet picture captures the core mechanism of RLVR, then the macroscopic dynamical signatures of the minimal implementation should also appear in any other system that satisfies the same conditions (sparse graph, shared parameters, sparse reward).
+The curve shapes depend on the initial policy and the number of training problems, and under different conditions—a pretrained model, a small problem set, or a single QA pair—they can look quite different (Power et al., 2022; Cai et al., 2025).
+What persists across all these regimes is cross-model consistency: the three systems track each other regardless.
+Figure 2 shows a representative case—a high-entropy initial policy with many training problems—for CoNet (no neural network), a 14K-parameter Transformer on the sliding puzzle, and DSR-1.5B trained under the DeepScaleR protocol (Luo et al., 2025) using the verl framework (Sheng et al., 2025; ByteDance Seed Team and verl community, ), covering five orders of magnitude in parameter count.
+All three display two-stage dynamics with shortening-then-lengthening responses.
+Among these shared features, the persistent growth of response length during Stage 2 is the most distinctive: unlike the two-stage reward profile, which is common to many optimization processes, the steady lengthening of reasoning chains is a geometric consequence of the inverse-tree picture (see below) and is not expected from generic convergence.
+4 Two Core Mechanisms: Merging and Frustration
+(a) Merging
+(b) Frustration
+(c) Step 20
+(d) Step 50
+(e) Step 800
+[…]
 (a) CoNet
 (b) LLM
-Figure 1: CoNet Reproduces Core LLM Training Dynamics.
-The minimal CoNet model (a) reproduces the two core empirical signatures of RLVR training observed in the DeepScaleR-1.5B LLM (b).
-These signatures are: (i) a two-stage reward dynamic consisting of a fast-learning stage followed by a slow-learning phase (top panels), and (ii) a non-monotonic, V-shaped evolution of the correct response length (bottom panels).
-This striking correspondence validates CoNet as a valuable minimal model for our theoretical analysis.
-The inset in (a) schematically depicts the CoNet abstraction.
-Lines in (b) are smoothed with a 60-step moving average.
-To investigate this hypothesis, we adopt a comparative approach connecting macroscopic phenomenology to graph topology.
-We first confirm that the DeepSeek-R1-Distill-Qwen-1.5B model (Luo et al., 2025a) exhibits the aforementioned RLVR signatures: a two-stage learning curve, a V-shaped response trajectory, catastrophic forgetting, and policy collapse.
-We then show that these complex behaviors are qualitatively reproduced by the Concept Network Model (CoNet) (Cai et al., 2025), a minimal framework abstracting away neural implementation details.
-This striking resemblance suggests that these phenomena are emergent collective behaviors rather than artifacts of specific neural architectures, allowing them to be studied via a coarse-grained proxy.
-Employing CoNet, we identify the emergence of a sparse concept web that orchestrates these dynamics.
-This phenomenological alignment supports the hypothesis that similar topological self-organization underlies LLM reasoning, enabling us to derive the following mechanisms from the principles of sparse graph growth:
-First, the V-shaped response length implies a structural evolution from local “skill islands” to a global concept web.
-The initial decrease results from the parallel optimization of independent skills, where the policy rapidly discards inefficient exploratory detours to converge on short, optimal paths for simple tasks (Fig.
-2).
-The subsequent rise, however, marks the onset of global integration, where a tree-like web emerges (Fig.
-3).
-On such a sparse substrate, the average geodesic distance is inherently correlated with network size.
-Thus, as the web expands to bridge distant islands during the slow-learning phase, the policy is geometrically forced to traverse longer paths.
-The rising V-curve is therefore not a change in reasoning style, but the result of the concept web’s continuous growth.
-Second, the sparse topology explains the model’s specific vulnerability to catastrophic forgetting (Fig.
-4).
-Unlike dense networks, a sparse web relies on critical “trunk” edges to support vast downstream branches.
-We show that SFT-induced forgetting acts as a topological disconnection: gradient updates overwrite weights at these high-traffic bottlenecks, severing bridges to entire sub-trees.
-This offers a geometric rationale for the “easy-to-break, easy-to-fix” nature of RLVR models; since the disconnected downstream knowledge remains intact, a brief RLVR resumption leads to rapid recovery by merely “re-soldering” the broken trunks rather than relearning the web.
-Third, we trace policy collapse to the dynamics of learning at the web’s frontier (the leaf nodes).
-Extending the critical learning theory of Cai et al.
-(2025), our experiments suggest that during the slow-learning phase, a significant portion of tasks sequentially undergo continuous-phase-transition-like learning dynamics (Fig.
-5).
-For these tasks, the policy shifts from broad exploration to locking onto specific solution paths.
-While this collapse is necessary for mastering individual skills, standard RLVR accumulates these local entropy reductions across the leaf nodes.
-This aggregation ultimately results in a global loss of diversity, manifesting as the policy collapse observed in reasoning benchmarks (Fig.
+Figure 4: The “learn-then-forget” signature of frustration-induced forgetting appears jointly in CoNet and the LLM.
+(a) Per-problem accuracy trajectories in CoNet; combined with the internal record of shared-node/edge conflicts, the non-monotonic “learn-then-forget” behaviour can be attributed directly to inter-task competition at shared nodes.
+(b) The same class of trajectories in DSR-1.5B, showing consistent non-monotonic patterns.
+The CoNet picture provides a mechanistic account of the PSV-graph contraction observed in Fig.
+1c: RLVR prunes a dense, high-entropy initial network into a sparse, low-entropy structure.
+Because the initial network is dense (Fig.
+1c, left), reasoning paths from different problems inevitably collide on shared nodes.
+Since multi-task RLVR forces all question–answer pairs to share a single set of transition probabilities $\{p(s^{\prime}\,|\,s)\}$, each shared node $s$ receives competing demands on its outgoing edges from different tasks.
+The central question is how these shared-node transition probabilities evolve under RLVR (Fig.
+3a,b).
+A collision admits only two outcomes.
+If the two tasks need the same outgoing edge from a shared node ss—their constraints on $p(s^{\prime}\,|\,s)$ are compatible—then GRPO’s positive feedback drives $p(s^{\prime}\,|\,s)\to 1$ and the two paths fuse at $s$.
+The cumulative effect of merging is to forge multiple independent paths into a shared chain-like structure: many entry points reach the same shared node along different prefixes and then flow towards the answer along a single high-probability channel.
+[…]
+Merging and frustration act jointly and point to a particular topology: a sparse multi-input, single-output directed tree—merging creates “multi-input,” frustration enforces “single-output.” We call this structure an inverse tree: multiple paths flow in and a single path flows out, the reverse of a conventional rooted tree.
+The argument is general: it requires only a sparse state space, shared transition probabilities across tasks, and positive feedback from policy gradient—no specific architecture or training algorithm.
+The minimal CoNet implementation makes this picture directly observable.
+By tracking every transition probability throughout training and thresholding at $p>0.95$, one can watch the two microscopic mechanisms reshape the network in real time.
+The result (Fig.
+3) confirms the predicted topology: out-degree close to 1, in-degree of merging nodes $\gg 1$, average degree $\langle k\rangle\approx 2$.
+Figure 3 shows this process at three training steps, marked on the CoNet training curve in Fig.
+2a.
+The formation mirrors the freezing of a supercooled liquid (Avrami, 1939; Kelton & Greer, 2010): disconnected subtrees nucleate independently (Step 20, Stage 1), then weld together at shared nodes as merging and frustration sculpt the network (Step 50), and the nucleation-to-merging crossover coincides with the Stage 1-to-2 inflection; by Step 800 the network has frozen into a single connected inverse tree.
+This sparse geometry has a direct consequence for response length.
+With out-degree close to 1 at almost every node, the graph has average degree $\langle k\rangle\approx 2$ and lacks short-range shortcuts.
+As the inverse tree extends to cover more question–answer pairs, geodesic distances grow monotonically—each edge is one coarse-grained reasoning step, so longer geodesics mean longer reasoning chains.
+When training begins from a high-entropy initial policy, the length trajectory is V-shaped: the early shortening reflects local optimization pruning redundant steps (Fatemi et al., 2025), while the subsequent rebound is the geometric signature of the sparse backbone having formed.
+The length curves across all three scales (Fig.
+2, bottom row) are consistent with this prediction.
+The frustration mechanism can be tested independently by tracking per-problem accuracy.
+In Stage 1, subtrees nucleate in isolation and do not interfere.
+Once they meet at shared nodes, competition prunes the subordinate task’s path, producing a characteristic non-monotonic “learn-then-forget” trajectory.
+The minimal CoNet makes this mechanism directly visible: by jointly recording per-problem accuracy and the internal conflict record at each shared node, the “learn-then-forget” curves can be attributed to specific inter-task competitions (Fig.
+4a).
+The same signature appears in DSR-1.5B (Fig.
+4b), providing experimental support in a real LLM.
+The Stage-1-to-2 inflection thus marks the maximally frustrated state—the moment of fiercest competition, when the inverse tree is most plastic.
+(a) CoNet
+(b) DeepScaleR-1.5B
+(c) Before SFT (Step 400)
+(d) After SFT (Step 400)
+(e) Resumed RLVR (Step 450)
+[…]
+5 Catastrophic Forgetting and the Inverse Tree Picture
+A widely studied empirical puzzle in RLVR training is catastrophic forgetting (Li & Hoiem, 2017; Luo et al., 2023; Ding & Wang, 2025): inserting a supervised fine-tuning (SFT) intervention into the RLVR process causes a sharp drop in reasoning capability.
+Unlike RLVR, in which the model discovers solution paths through autonomous exploration, SFT requires the model to memorize prepared reasoning trajectories directly—maximizing token-by-token the conditional likelihood of the given trajectory, with no trial-and-error of the model’s own.
+The phenomenon has been reported in multiple independent studies (Luo et al., 2023; Ding & Wang, 2025; Fernando et al., 2025; Matsutani et al., 2025).
+To amplify and study it under controlled conditions, we designed a dedicated SFT experiment (details in the supplementary material): the SFT trajectories are sourced from the model’s own rollouts on hard problems in the training set, filtered by rejection sampling to retain only correct solutions—a deliberate data-preparation strategy intended to minimize the distributional gap between SFT data and the current policy.
+Applied during Stage 2 of DSR-1.5B’s RLVR training (Fig.
+5b), this SFT triggers an immediate sharp reward drop; once SFT is stopped and RLVR resumes, performance rapidly returns to near its pre-SFT level.
+Even with SFT data drawn from the model’s own correct trajectories—and the distributional gap therefore minimized—catastrophic forgetting cannot be avoided.
+This means that the root of forgetting lies not in any data distribution mismatch but in a conflict between the SFT mode of training itself and the internal structure that RLVR has already established.
+The minimal CoNet implementation reproduces exactly the same response (Fig.
+5a) and additionally provides direct microscopic observation.
+Before SFT, the inverse tree is a single, fully connected structure (Fig.
+5c); after SFT, the tree fragments from the connected state into several disconnected subgraphs (Fig.
+5d); once RLVR resumes, the broken bridges are rapidly re-welded and the tree is restored (Fig.
+5e).
+The inverse-tree picture supplies a unified explanation.
+Because almost every node has out-degree close to 1, the global connectivity of the inverse tree depends concentratedly on a small number of bridge nodes (high-in-degree merging points where multiple paths converge).
+SFT overwrites the transition probabilities only at the nodes its supervised trajectories visit, with no regard for the global tree structure already in place.
+Such trajectories are highly likely to traverse the key bridge nodes of the inverse tree, and an SFT overwrite at such a position effectively cuts the bridge—even when the overwrite happens to agree with the inverse-tree direction, the global modification of the outgoing-edge probabilities at that node disrupts other paths passing through the same node.
+The severity of forgetting is therefore set by the number of bridge nodes the SFT trajectory traverses, not by the quality of the SFT data itself—which explains why even using the model’s own correct trajectories does not avert the destruction.
+The “break-but-not-erase” recovery, in turn, follows from the same topology: bridge nodes are few, hence fragile; but breakage does not destroy the subtrees themselves, so RLVR only needs to repair a small number of bridges to restore global connectivity.
+(a) In-distribution (512 problems)
+(b) OOD: Minerva
+(c) OOD: AIME 2024/25
+Figure 6: Annealed-RLVR alleviates policy collapse, with the advantage concentrated at large-$k$ sampling.
+(a) In-distribution 512 problems; (b, c) out-of-distribution Minerva and AIME 2024/25 benchmarks.
+Orange = Annealed-RLVR; blue = standard RLVR.
+The advantage of Annealed-RLVR grows monotonically with $k$, in agreement with the prediction that frustration-induced policy collapse is concentrated at large $k$.
+6 Annealed-RLVR: From Physical Picture to Practical Method
+Policy collapse is one of the central practical challenges of RLVR training (Yue et al., 2025; Dong et al., 2025; Zheng et al., 2025): as training proceeds, best@1 keeps improving while policy diversity is steadily lost—best@$k$, the probability that at least one of $k$ independent samples is correct, stagnates or even declines at large $k$ (blue curves in Fig.
 6).
-Finally, identifying a “maximally frustrated state” at the onset of slow learning—where competition between disconnected islands peaks—motivates our Annealed-RLVR algorithm.
-By introducing a targeted SFT “heating” step at this topological bottleneck, we temporarily relax competitive tension, helping the system navigate past local optima before cooling into a refined structure.
-Our experiments confirm that this theory-driven approach resolves the bottleneck and achieves superior reasoning performance on in-distribution and out-of-distribution benchmarks (including Minerva and AIME), surpassing standard RLVR baselines by engineering the topology of reasoning.
-(a) Step 20
-(b) Step 50
-(c) Step 800
-Figure 2: Visualizing the Structural Evolution from “Skill Islands" to the Concept Web.
-These network snapshots provide a direct, qualitative narrative of the concept web’s formation, corresponding to the different phases of training.
-In each subfigure, green, red, and blue dots indicate the question, answer, and intermediate nodes of the CoNet, respectively.
-The color of each directed edge represents the transition probability from the head to the tail and is drawn according to the colorbar in the rightmost column.
-Here, only edges with transition probabilities greater than $0.95$ are shown.
-(a) In the early fast-learning phase (Step 20), the model discovers a few short, disjointed reasoning paths, representing the first nascent “skill islands”.
-(b) At the onset of slow learning (Step 50), the system has proliferated into a maximal collection of disconnected islands.
-(c) Deep in the slow-learning phase (Step 800), these previously separate islands have coalesced into a single, giant connected component, forming the unified and expansive concept web.
-2 The V-Shaped Signature as Evidence for a Sparse Concept Web
-The V-shaped evolution of the response length is a well-demonstrated, quantifiable hallmark of deep structural reorganization during RLVR training (He et al., 2025; Luo et al., 2025a).
-To anchor our investigation, we first reproduce this signature by applying RLVR to the DeepSeek-R1-Distill-Qwen-1.5B model following the DeepScaleR protocol (Luo et al., 2025a).
-As shown in Fig.
-1, our experiment recovers the characteristic V-shaped curve (Fig.
-1).
-To dissect the mechanism underlying this signature—a task intractable within the LLM itself—we deploy an extended, multi-task variant of the CoNet as a “computational microscope” (See Appendix D for details).
-CoNet instantiates the implicit reasoning graph discussed in the Introduction as a concrete, tractable model: it maps the LLM “semantic states” to abstract nodes in a fixed random graph and LLM “logical transitions” to learnable, probabilistic edges.
-This minimal model reduces the learning process to a tractable graph-traversal problem and, remarkably, reproduces the V-shaped curve observed in our LLM experiment (Fig.
-1), validating it as a faithful proxy for uncovering the underlying mechanism.
-(a) Correct Response Histogram
-(b) Step 50
-(c) Step 800
-Figure 3: A Sparse-Web Structure Necessitates Longer Reasoning Chains.
-This figure links the emergent sparse topology of the concept web to the observed increase in response length.
-(b, c) The color and marker scheme is identical to that used in Fig.
-2, and only edges with transition probability greater than $0.95$ are retained.
-The largest connected component (a local region of which is shown here) remains sparse with average degree around 22 even as it grows from step $50$ to $800$.
-(a) Consequently, the distribution of solution path lengths, shown in both raw counts (histogram) and probability density (inset), shifts decisively to the right during this period, confirming that navigating the sparse backbone requires longer reasoning chains.
-Armed with this confirmation, we now demonstrate that the V-shaped evolution of the response length arises from a two-stage process.
-The decreases of the response length in the initial fast-learning stage is dominated by local optimization.
-Given the random initialization of transition probabilities, shorter paths inherently have a higher probability of being sampled.
-Consequently, the model rapidly discovers short, efficient solutions to individual problems, giving rise to a proliferation of disconnected “skill islands” (as shown in Fig.
-2(a)).
-This burst is quantified by the number of distinct solution clusters, which peaks right before the onset of slow learning (Fig.
-2(b)).
-Stage one thus constitutes a parallel search for local efficiency that naturally shortens the average response length.
-The subsequent slow-learning phase, where response length steadily increases, marks a fundamental shift to global integration.
-The primary task is no longer discovering new islands but weaving them into a single, expansive concept web (see Fig.
-2(c)), defined to be the largest connected component in which every edge has a transition probability above $0.95$.
-As the concept web relentlessly grows, a critical structural property emerges: the web remains persistently sparse, with its average degree stabilizing near two (as shown in Fig.
-3(c)).
-This provides direct, quantitative validation for our central Hypothesis (Hypothesis 1).
-This emergent sparsity is the topological key to the V-shape’s rising slope.
-A network with such a low average degree is dominated by chain and tree-like structures, inherently lacking the shortcuts and redundant pathways found in denser graphs.
-Consequently, connecting two previously distant concepts requires traversing long, intermediate paths.
-The histograms in Fig.
-3(a) provide clear evidence of this effect: As the sparse web expands, the distribution of solution lengths undergoes a decisive and sustained rightward shift, confirming that navigation on this sparse backbone necessitates longer reasoning chains.
-Thus, the V-shaped curve furnishes a robust macroscopic signature of an underlying topological shift: from parallel, local optimization on disconnected components to serial, global navigation across an expanding, sparse backbone.
-Yet can this simple, largely tree-like model capture the full complexity of LLM reasoning?
-After all, an LLM’s autoregressive memory enables sophisticated behaviors such as self-correction or reflection (e.g., “Wait, let me rethink that…”)—that would manifest as cycles in the reasoning graph.
-We argue they are fully compatible(Minegishi et al., 2025).
-These reasoning cycles are not the dense, ubiquitous shortcuts of a highly connected graph.
-Instead, they are likely the cognitive manifestation of sparse, local loops—occasional, high-cost revisions on a much larger, fundamentally chain-like scaffold.
-The existence of a limited number of such cycles does not significantly alter the global average degree of the web, which remains close to two.
-With the sparse-web model established, this link between a macroscopic dynamic and a microscopic topology forms the foundation for the analysis that follows.
-(a) CoNet
-(b) LLM
-(c) Step 400 (Before SFT)
-(d) Step 400 (After SFT)
-(e) Step 450 (Continued RLVR)
-Figure 4: The Fragility of a Sparse Web: Catastrophic Forgetting and Fast Recovery.
-This figure demonstrates a key prediction of our sparse-web hypothesis: that the concept web is fragile, relying on critical bridge-like connections.
-(c-e) The color and shape conventions follow those of Fig.
-2, and only a local view of the concept web is shown.
-Orange crosses on the edges indicate decreases in transition weights, with thicker crosses denoting more substantial decreases.
-The microscopic view in CoNet shows how a lightweight supervised fine-tuning (SFT) on a converged web (c) severs these critical bridges, causing the structure to fragment (d).
-Subsequent RLVR rapidly repairs these links (e).
-(a, b) This microscopic severing manifests as macroscopic catastrophic forgetting in both CoNet and the 1.5B LLM, where performance plummets upon initiating SFT.
-The subsequent fast recovery once RLVR resumes highlights the localized nature of the damage, underscoring the web’s structural fragility.
-3 The Double-Edged Sword of Sparsity: from catastrophic Forgetting to Simulated Annealing scheme
-From the graph theoretic point of view, the sparsity in the conceptual network implies low redundancy, fragility, and critical vulnerability.
-It also reflects a structural hypothesis for a well-known operational challenge: why models that have formed a concept web via RLVR are so susceptible to catastrophic forgetting when subsequently fine-tuned with supervised learning (Li et al., 2024).
-Our theory posits that this is not a widespread erasure of knowledge, but a specific topological failure: the surgical severing of critical, bridge-like connections.
-To test this mechanistic hypothesis, we subjected models in the deep slow-learning phase to a lightweight in-distribution SFT (see Appendix H and I).
-As expected, the intervention triggered a sharp drop in overall performance in both CoNet and the 1.5B LLM (Fig.
-44(a), 4(b)).
-Having reproduced this known macroscopic phenomenon, we leveraged CoNet’s transparency to inspect the microscopic aftermath.
-The result is a direct validation of our structural explanation.
-The microscopic view reveals that the SFT catastrophic forgetting was not a global degradation of the web.
-Instead, SFT’s aggressive optimization of a few paths overwrote the policy at critical branching nodes, precisely breaking the single connection points for vast downstream subgraphs.
+Dominant paths grow ever stronger while the covered solution space contracts.
+Existing remedies rely largely on entropy regularization (Cui et al., 2025; Hao et al., 2025; Jiang et al., 2025)—adding an entropy penalty to the loss to maintain policy stochasticity.
+Entropy regularization, however, is a global and persistent intervention: it cannot distinguish “beneficial growth of certainty” (the pruning of redundant paths in Stage 1) from “harmful loss of diversity” (the pruning of legitimate alternatives in Stage 2), and can only strike a compromise between the two.
+The inverse-tree picture offers a different perspective: policy collapse is precisely the macroscopic accumulation of frustration-induced forgetting discussed above.
+The problem is not that the system’s overall “temperature” is too low but that, throughout RLVR’s monotonic cooling, the system has no opportunity to escape the configurations frozen in by frustration.
+Borrowing the idea of simulated annealing from condensed-matter physics (Kirkpatrick et al., 1983)—reheating during cooling to restore ergodicity—the inverse-tree picture suggests an intervention dimension that has rarely been considered in current methodology: structural timing.
+The concrete scheme is to apply a brief SFT “heating” at the moment of maximum frustration (the Stage-1-to-Stage-2 inflection) and then resume RLVR to let the system cool back down.
+The timing is empirically located by the inflection of the training curve where the reward growth slows and the response length turns from shrinking to rebounding; the SFT data are the same self-rollout correct solutions used in the catastrophic-forgetting experiment (parameter values in the supplementary material).
+The central insight is not in SFT itself but in the fact that the same operation has opposite effects at different stages of the inverse tree’s growth—a notion that can be generalised to other forms of policy perturbation.
+Figure 6 shows the experimental results on DSR-1.5B.
+The three panels report best@$k$ on, respectively, the in-distribution 512 training problems (a) (Hendrycks et al., 2021), the out-of-distribution Minerva benchmark (b) (Lewkowycz et al., 2022), and the out-of-distribution AIME 2024/25 benchmark (c); the horizontal axis is the sample count $k$ and the vertical axis is best@$k$.
+On all three benchmarks, Annealed-RLVR (orange) outperforms standard RLVR (blue), with a striking pattern: the gap is small at $k=1$ but grows monotonically with $k$, peaking at $k=64$.
+This pattern carries a clear physical meaning: annealing restores the subordinate paths that frustration had pruned; the dominant path (which determines best@1) is itself unaffected, so the improvement is concentrated along the policy-diversity axis.
+An improvement of best@$k$ at large $k$ means the model has regained multiple independent solution strategies rather than producing all samples along a single path.
+Notably, the improvement holds and is even more pronounced on the out-of-distribution benchmarks (Fig.
+6b,c), indicating that the recovered diversity generalises and is not confined to the training set.
+A causal test of timing further supports the picture.
+Postponing the same SFT to Step 400, by which time the inverse tree has fully solidified, not only fails to improve performance but in fact triggers catastrophic forgetting; resumed RLVR only returns the model to its pre-SFT level (Fig.
+5b).
+The same SFT data, the same amount of training—and yet the effect reverses depending solely on timing.
+This rules out the trivial interpretation that “SFT supplied useful information” and shows that the improvement of Annealed-RLVR is attributable specifically to recovering ergodicity in the policy space at the right structural moment.
+7 Conclusion and Outlook
+This work began from a physics question—what does RLVR change at the microscopic level so that long-range reasoning emerges in LLMs?—and arrived at a statistical-physics answer: RLVR drives the freezing-in of an inverse-tree structure on the model’s effective state space.
+The formation process is dynamically akin to freezing, but the vast configurational degeneracy hints that the final state may sit closer to a glassy phase, awaiting further characterization.
+The supporting logic of this answer can be read at three levels.
+The first is representation: the self-referentiality of an autoregressive model and its finite parametric capacity force the effective state space to be compressed into a sparse Markov network; microscopic histories of order $V^{t}$ collapse into PSV equivalence classes, with compression ratios of 65–91% measured in an exhaustively enumerable system.
+The second is dynamics: on this sparse network, shared-parameter multi-task RLVR, through the merging of compatible paths and the frustrated competition of incompatible ones, freezes the dense initial network into a sparse, multi-input, single-output directed tree—an inverse tree—through three stages of nucleation $\to$ growth $\to$ merging.
+This dynamics displays qualitatively consistent macroscopic signatures across three systems whose parameter counts span five orders of magnitude.
+The third is prediction and intervention: the geometric properties of the inverse tree directly yield three falsifiable predictions—reasoning chains lengthen with capability (a geodesic-distance constraint of the sparse tree), the break-and-recover response under SFT perturbation (topological fragility at bridge nodes), and frustration-driven policy collapse—all supported by experiment.
+Annealed-RLVR, designed on the strength of this picture, alleviates policy collapse while preserving exploration capacity, and a controlled experiment shows that the improvement is tied to the timing the mechanism specifies.
+The significance of this picture extends beyond the explanation of specific empirical phenomena.
+It provides a methodology for translating LLM reasoning into a statistical-physics problem: extracting effective degrees of freedom as PSV equivalence classes (analogous to coarse-graining), reducing reasoning to path search on a sparse graph (analogous to a Markov process), translating training dynamics into freezing on the graph (analogous to nucleation and growth), and recasting multi-task competition as frustration (analogous to spin glasses).
+Once this translation dictionary is in place, the existing toolbox of statistical physics—phase-transition theory, critical phenomena, and the replica method for disordered systems (the standard tool for handling vast quasi-degenerate states)—may all be brought to bear on the theoretical analysis of LLM reasoning.
+Acknowledgements
+The authors thank Xianrui Ke, Daiqiu Mou, Yuansheng Cao, Weinan E, Haijun Zhou, Lei Wang, Hong Zhao and Ganqu Cui for inspiring discussions.
+K.
+C.
+and X.
+C.
+are supported by the National Key Research and Development Program of China under Grant No.
+2024YFA1408604 and the National Natural Science Foundation of China under Grants No.
+12474245 and No.
+12447103.
+P.
+Z.
+is supported by Project 12325501 of the National Natural Science Foundation of China.
+Y.
+D.
+and S.
+H.
+are supported by the National Natural Science Foundation of China under Grants No.
+12275263, the Quantum Science and Technology-National Science and Technology Major Project Grant No.
+2021ZD0301900, and the Natural Science Foundation of Fujian Province of China under Grant No.
+2023J02032.
+Z.
+Y.
+is supported by the National Natural Science Foundation of China under Grants No.
+12304288 and No.
+12247101.
+Appendix A Sliding-Puzzle Experiment: Architecture, Training, and Compression Measurement
+This section gives the full experimental specification for the sliding-puzzle Transformer used in §EFFECTIVE MARKOV REPRESENTATION of the main text, on which the PSV compression ratio $R_{c}$ is measured directly.
+A.1 Environment: $2\times 3$ Sliding Puzzle as a Reasoning Task
+The $2\times 3$ board contains five labelled tiles ($1,\ldots,5$) and one blank.
+To make the dynamics a regular graph we restrict the blank to the middle column (positions $1$ and $4$ in row-major indexing) at all times.
+Enumerating the $5!=120$ permutations of tiles with the blank at each of the two middle-column positions gives $240$ candidate configurations, of which exactly $120$ form a single connected component under the action set below (the other $120$ are unreachable due to permutation parity).
+The reachable set has $|\mathcal{S}|=N=120$ states.
+The action set $\mathcal{A}=\{\mathrm{V},\mathrm{L},\mathrm{R}\}$ consists of three composite moves that all preserve the “blank-in-middle-column” invariant:
+  • $\mathrm{V}$ (vertical swap): swap the blank between positions $1$ and $4$.
+  • $\mathrm{L}$ (left rotation): three-cycle the tiles in the left half of the board.
+  • $\mathrm{R}$ (right rotation): three-cycle the tiles in the right half of the board.
+Every legal state admits all three actions, so the resulting state graph is $K=3$-regular (and bipartite under this generating set; cf. Fig. 1a of the main text).
+A reasoning task is a question–answer pair $(Q,A)$ consisting of a starting state $Q$ and a target state $A$.
+For each fixed shortest-path distance $D\in\{4,6\}$ we draw $(Q,A)$ uniformly from all ordered state pairs satisfying $\mathrm{BFS}\text{-}\mathrm{dist}(Q,A)=D$; the first $B=500$ such pairs are used for training and the next $200$ for validation, so train and validation sets are disjoint subsets of the same candidate pool.
+The two reported settings (dist=4=4, dist=6=6; Fig.
+7) correspond to two independent training experiments, not a curriculum.
+The model interacts with the environment by emitting a sequence of action tokens $a_{1},a_{2},\ldots$; the state advances deterministically and the episode terminates either when the answer state $A$ is reached (reward $1$) or the maximum length $L_{\max}=12$ is hit (reward 0).
+A.2 Transformer Architecture (14K Parameters)
+The policy $\pi_{\theta}(a_{t+1}\mid Q,a_{1},\ldots,a_{t})$ is parameterised by a single-layer decoder-only Transformer with $\sim 14{,}000$ parameters in total.
+The configuration is summarised in Table 1.
+Table 1: Architecture of the sliding-puzzle Transformer.
+Component Value
+Number of decoder layers 1
+Embedding / hidden dimension $d_{\text{model}}$ 32
+Number of attention heads 2 (head dim $d_{h}=16$)
+Feed-forward dimension $d_{\text{ff}}$ 128
+Activation ReLU
+Dropout 0
+[…]
+Token-type embedding 2 types (state, action), additive
+Maximum sequence length $L_{\max}^{\text{seq}}$ 37
+Tile vocabulary 7 (six tile ids $0,\ldots,5$ plus a sep token)
+Action vocabulary 3 (independent of tile vocabulary)
+Output head Linear($d_{\text{model}}=32\to 3$)
+Total parameter count $\sim 14{,}000$
+The tile and action vocabularies use independent embedding tables; the model never tokenises an action with the tile vocabulary or vice versa. The prompt is encoded as
+$\textsc{prompt}\;=\;[\,Q_{1},Q_{2},\ldots,Q_{6},\ \textsc{sep},\ A_{1},A_{2},\ldots,A_{6}\,]\qquad(\text{length}\ P=13),$ (2)
+where $(Q_{1},\ldots,Q_{6})$ and $(A_{1},\ldots,A_{6})$ are the row-major tile sequences of the question and answer states.
+After the prompt, the model autoregressively appends action tokens (each of which contributes its action embedding plus the position embedding plus the action token-type embedding).
+A standard causal mask is applied throughout, and the next-action logits are read from the position of the most recently emitted action token.
+A.3 GRPO Training Protocol
+The policy is trained by Group Relative Policy Optimization (GRPO) starting from a randomly initialised network — no pretraining is used, since the entire reachable state space ($N=120$) is small enough that the policy can be learned from scratch.
+Hyperparameters are listed in Table 2.
+Table 2: Training hyperparameters for the sliding-puzzle Transformer.
+Setting Value
+Optimiser AdamW
+Learning rate $3\times 10^{-4}$
+Weight decay 0
+PPO clip $\varepsilon_{\text{clip}}$ $0.2$
+KL coefficient $\beta$ 0 (main $R_{c}$ experiments)
+Inner update epochs $K$ per GRPO step 44
+Mini-batch size (trajectories) 32,00032{,}000
+Sampling temperature $1.0$
+GRPO group size $G$ (rollouts per $(Q,A)$) $1024$
+Number of training $(Q,A)$ pairs $B$ $500$
+Number of validation $(Q,A)$ pairs $200$
+Maximum episode length $L_{\max}$ $12$
+Reward +1+1 if answer reached, 0 otherwise
+Total training epochs 1,0001{,}000
+Random seed $42$ (single seed; no multi-seed averaging)
+Each training epoch performs the same loop: for every one of the $B=500$ training pairs, $G=1024$ trajectories are rolled out under the current policy; rewards are normalised within each group (group baseline); $K=4$ inner GRPO updates are then performed on the resulting batch; finally the validation set is evaluated.
+The training curve is reported in Fig.
+1b of the main text.
+A.4 PSV Equivalence Relation: Operational Definition
+The PSV equivalence relation of Eq. (1) of the main text is implemented exactly via subtree hashing rather than via a continuous distance.
+For every active decision point $(Q,h)$ we expand its complete future tree, pruning any branch whose conditional next-action probability is below $p_{\text{prune}}=0.05$ (such branches contribute negligibly to behaviour).
+At every surviving node we compute the rounded next-action distribution
+$\widetilde{\pi}_{\theta}(\cdot\mid Q,h)\;=\;\big(\,\mathrm{round}_{r}(\pi_{\theta}(a_{1}\mid Q,h)),\ \ldots,\ \mathrm{round}_{r}(\pi_{\theta}(a_{K}\mid Q,h))\big),$ (3)
+where $\mathrm{round}_{r}(\cdot)$ rounds to $r$ decimal places. The signature of $(Q,h)$ is the recursively constructed tuple
+$\sigma(Q,h)\;=\;\big(\,\widetilde{\pi}_{\theta}(\cdot\mid Q,h),\ \sigma(Q,h\!\cdot\!a_{1}),\ \ldots,\ \sigma(Q,h\!\cdot\!a_{K})\big),$ (4)
+which is a Python tuple equipped with exact equality.
+Two histories are declared equivalent if and only if their signatures are byte-identical: $(Q,h)\sim_{r}(Q^{\prime},h^{\prime})\iff\sigma(Q,h)=\sigma(Q^{\prime},h^{\prime})$.
+Equivalently, this enforces $|\pi_{\theta}(\cdot\mid Q,h)-\pi_{\theta}(\cdot\mid Q^{\prime},h^{\prime})|\leq 5\times 10^{-r-1}$ at every shared decision point in the future tree, so the rounding precision $r$ acts as the discrete counterpart of the threshold $\epsilon$ in Eq.
+(1) of the main text.
+The robustness check below uses $r\in\{1,2,3,4\}$.
+The set of active decision points is constructed by enumerating, for every $(Q,A)$ pair in the union of the training set and the validation set, the BFS subtree $\mathcal{T}_{Q}$ rooted at $Q$ to depth $L_{\max}$, retaining only those $(Q,h)$ pairs for which (i) $h$ has not yet hit the answer or the depth limit and (ii) every action in $h$ has conditional probability $\geq p_{\text{prune}}=0.05$.
+The size of the active set is then taken as the sum (not the deduplication) of node counts across all problems:
+$|\mathcal{S}_{\text{active}}|\;=\;\sum_{(Q,A)\in\,\text{train}\cup\text{val}}\big|\mathcal{T}_{Q}^{\text{active}}\big|.$ (5)
+Crucially, the signature $\sigma$ above does not include $Q$ as part of its hash — it depends only on the local future-tree structure of the policy.
+Consequently, two histories from different problems whose future subtrees are byte-identical contribute the same signature and are merged into a single PSV class.
+This implementation choice is what makes $R_{c}$ sensitive to the cross-problem state sharing emphasised in §EFFECTIVE MARKOV REPRESENTATION of the main text: as RLVR carves shared reasoning paths across problems, the count of distinct signatures $|\mathcal{S}_{\text{PSV}}|$ grows much more slowly than $|\mathcal{S}_{\text{active}}|$, driving $R_{c}$ upward in Stage 2.
+A.5 Compression Ratio $R_{c}$
+The compression ratio is defined as
+$R_{c}\;=\;1-\frac{|\mathcal{S}_{\text{PSV}}|}{|\mathcal{S}_{\text{active}}|},$ (6)
+[…]
+The trajectory of $R_{c}$ during training (Fig.
+1b of the main text) traces a non-monotonic V-shape: an initial drop driven by policy diversification, followed by a rebound and saturation in a high-compression regime.
+The V-shape is the direct empirical signature of PSV-class refinement followed by cross-problem state sharing, both of which are predictions of the effective Markov representation introduced in §EFFECTIVE MARKOV REPRESENTATION of the main text.
+[…]
+A.6 Robustness of $R_{c}$ to Rounding Precision
+The threshold $\epsilon$ in the PSV equivalence relation is implemented by the rounding precision $r$.
 Fig.
-44(c)4(d) provides a stark visualization of this: a unified web is instantly fragmented, offering a clear, structural cause for the catastrophic performance drop.
-However, this is where the sword’s second edge is revealed.
-The damage, while profound in its effect, is topologically localized.
-SFT does not erase the knowledge encoded in the now-disconnected subgraphs; it merely renders them unreachable.
-This localization explains the remarkable speed of recovery, which can be observed in Fig.
-44(a)4(b).
-When RLVR training resumes, it triggers a quick rebound because the task is simply to “re-solder” a few broken connections to restore access to the latent scaffold as shown in Fig.
-44(e).
-This two-step process of localized disruption followed by rapid repair is strongly analogous to simulated annealing Kirkpatrick et al.
-(1983).
-The brief SFT phase acts as a high-temperature “heating” shock, breaking connections and allowing the system to escape a potentially suboptimal state.
-The subsequent RLVR phase then acts as a controlled “cooling”, guiding the system as it re-forms connections, ideally settling into a more robust configuration (Lai et al., 2025).
-This insight—that a known vulnerability can be repurposed as a deliberate annealing step—forms the foundation for the Annealed-RLVR algorithm, which we introduce in Sec.
-5.
-4 Microscopic Mechanisms of RLVR: Forgetting by Frustration, Learning by Phase Transition
-Although the overall reward curve in RLVR training looks smooth, our sparse-web theory shows that much more is happening underneath.
-As shown in the individual trajectories in Fig.
-5, the network’s sparse structure creates two competing effects: (i) a forgetting process caused by frustration, which is strongest at the start of the slow-learning phase, and (ii) an opposite, phase-transition-like learning process that gradually builds and integrates knowledge during slow learning.
-(a) CoNet
-(b) LLM
-Figure 5: Microscopic Mechanisms of RLVR: Forgetting by Frustration, Learning by Phase Transition The learning trajectories for individual problems in CoNet (a) and the 1.5B LLM (b) reveal a fundamental duality.
-(i) Frustration-Induced Forgetting: At the onset of slow learning [see inset in (a)], the intense competition for connections on a sparse web manifests as volatile, non-monotonic accuracy curves, where some skills are competitively suppressed.
-(ii) Phase-Transition-Like Learning: Subsequently, the web’s sparse frontier enables new skills to be acquired in punctuated, accelerated jumps (orange and green).
-The smoother gradients in the LLM (vs.
-the clean steps in CoNet) can be attributed to finite-size effects, as explained in the main text.
-4.1 Frustration-Induced Forgetting: The Price of Integration
-The first dynamic is driven by frustration-induced forgetting: a competitive process where the advancement of some skills forces the deterministic regression of others as the model integrates its “skill islands”.
-This process manifests as volatile, non-monotonic learning curves in the learning dynamics (Fig.
-5) and represents an internal “push-and-pull” fundamentally different from the external shock of SFT-induced catastrophic forgetting (Shenfeld et al., 2025; Jin et al., 2025a).
-This process peaks in the maximally frustrated state at the onset of the slow-learning stage, a critical juncture where “skill islands” compete for limited connections under the $\left<k\right>\approx 2$ constraint.
-Paradoxically, this state maximizes exploration: the model’s ability to find diverse, correct out-of-distribution solutions peaks, far exceeding the base model or the same model after prolonged training (Fig.
-6(b), blue).
-This subsequent degradation in performance is consistent with the phenomenon of policy collapse, which has been recently reported in several studies to occur after extended RLVR, leading to a loss of exploratory power (Chen et al., 2025; Yue et al., 2025; Cui et al., 2025).
-4.2 Phase-Transition-Like Learning: The Engine of Growth
-The opposing dynamic, phase-transition-like learning, is an important engine of new knowledge acquisition (Power et al., 2022; Lin et al., 2025).
-The evidence for this is clearly shown in CoNet’s sharp, step-like accuracy jumps (Fig.
-5(a)), and the steep accelerated learning gradients observed in the 1.5B LLM (Fig.
-5(b), more details in Appendix G).
-This phenomenon is a multi-task extension of the “Learning at Criticality” (LaC) mechanism, first identified in a single-task context (Cai et al., 2025).
-The LaC framework models this process as a continuous phase transition (Stanley & Ahlers, 1972; Wilson, 1983) where the policy collapses from exploring diverse reasoning paths to exploiting a single optimal one.
-Crucially, analysis revealed that near this critical point, the system’s exploratory power is maximized, and the reasoning path lengths become scale-invariant—a hallmark of continuous phase transitions.
-Beyond this transition, the policy collapses onto fixed optimal paths; the associated generalization drop is possibly linked to overfitting within the RLVR process (Dong et al., 2025; Gui & Cheng, 2025; Shypula et al., 2025; Zheng et al., 2025).
-This context explains the smoother gradient in the LLM.
-Compared to the step-like jumps in CoNet, the LLM’s learning curve is a classic manifestation of finite-size effects (Fisher & Barber, 1972): it has a more constrained set of distinct reasoning paths for any given task, which rounds out the idealized sharp transition into a steep but continuous curve.
-The puzzle is why this delicate phenomenon survives in a chaotic multi-task setting.
-Our sparse-web model explains this through topological isolation: a node on the expanding frontier of a $\left<k\right>\approx 2$ network is statistically isolated, decoupling it from system-wide competition.
-Shielded from interference, a new skill can be efficiently integrated via a clean phase transition.
-This perspective also provides a microscopic explanation for aggregate policy collapse in the standard RLVR (Chen et al., 2025; Yue et al., 2025; Cui et al., 2025).
-We hypothesize this is not a monolithic failure but the statistical sum of asynchronous, microscopic collapses at the frontier, a hypothesis supported by the LaC finding that training just beyond criticality risks overfitting-induced collapse.
-5 Annealed-RLVR: An Algorithm Born from the Sparse-Web Theory
-(a) Training Sample (512 Problems)
-(b) Minerva
-(c) AIME 2024/2025
-Figure 6: Annealed-RLVR Outperforms Standard RLVR.
-This figure compares the best@k accuracy curves of our Annealed-RLVR ("Annealed") with the standard RLVR baseline ("RLVR").
-The evaluation is performed on (a) an in-distribution set (Randomly Selected 512 Training Problems) and two out-of-distribution (OOD) sets: (b) the Minerva dataset and (c) the AIME 2024/2025 datasets.
-The results show that Annealed-RLVR consistently achieves superior performance across all in-distribution and OOD benchmarks, confirming that the intervention fosters improved generalization.
-Our sparse-web theory reveals a core narrative of RLVR training: the system drives itself into a maximally frustrated state that is both a perilous bottleneck and a moment of peak exploratory potential (Fig.
-6(b), blue curve).
-This insight motivates Annealed-RLVR, a theory-driven intervention that leverages the high-entropy frustrated state to guide the model toward a more globally optimal configuration (Cheng et al., 2025; Chen et al., 2025).
-The algorithm operationalizes this concept with a precisely timed intervention.
-This intervention is triggered by monitoring the macroscopic signatures of the maximally frustrated state, which, as shown in Fig.
-1, correspond to the crossover regime characterized by the “knee” of the reward curve (when accuracy gains plateau) and the bottom of the V-shaped response length curve.
-At this juncture, a brief SFT “heating” phase begins.
-The timing is critical: because the “skill islands” have not yet solidified into a single web, the system is uniquely robust to a targeted shock that would otherwise cause catastrophic forgetting—a theoretical claim confirmed by our experiments (Fig.
-4).
-The SFT thus functions not as a disruptive overwrite but as a corrective jolt to break the policy’s premature commitment to suboptimal routes.
-We implement this by surgically applying SFT only to problems with low accuracy (e.g., <0.1<0.1) for which correct solutions exist.
-Immediately following this targeted heating, the “cooling” stage resumes standard RLVR, allowing the now more pliable policy to re-settle into a more robust and integrated final state.
-We validated this approach on the DeepSeek-R1-Distill-Qwen-1.5B model, trained on the DeepScaleR-Preview-Dataset .
-Our method, Annealed-RLVR, was compared against the standard RLVR baseline where the policy is optimized using a variant of the Group Relative Policy Optimization (GRPO) algorithm (Shao et al., 2024; Liu et al., 2025; Yu et al., 2025).
-The results (Fig.
-6) confirm our theory.
-Best@k accuracy curves demonstrate that Annealed-RLVR outperforms standard RLVR on both in-distribution evaluation set (512 randomly selected training problems) and on out-of-distribution (OOD) benchmarks, including the Minerva and AIME 2024/2025 dataset.
-The histograms (see Fig.
-13) reveal the underlying mechanism: the SFT “heating” provides a crucial performance boost to the most difficult problems, lifting their success rates from near-zero and making them “visible” to subsequent RLVR exploration.
-The RLVR “cooling” phase then efficiently guides these newly accessible problems toward mastery.
-This results in a systematic reduction of unsolved problems and a corresponding population shift towards solved ones, confirming that our intervention guides the model past the inherent bottlenecks of standard RLVR to achieve superior reasoning.
-More details are given in Appendix H and I.
-6 Conclusion and Outlook
-We have introduced a sparse-concept-net theory to explain the puzzling behaviors of the RLVR training, including the two-stage learning curve, V-shaped response curves, and vulnerability to catastrophic forgetting.
-Our theory proposes that these diverse phenomena are consequences of one underlying process: the formation of a sparse concept web, with the averaged degree pinned close to two.
-This physical model directly explains the V-shaped curve as a shift from local optimization to global integration and recasts catastrophic forgetting as the severing of the web’s critical, bridge-like connections.
-A natural consequence of this sparse structure is a fundamental RLVR mechanism for forgetting and learning within the microscopic training process.
-We find that the degree-2 constraint first forces the system into a maximally frustrated state at the onset of slow learning, a phase dominated by frustration-induced forgetting as skills are competitively suppressed.
-Subsequently, the same sparse structure enables a cleaner growth mechanism: phase-transition-like learning, where new knowledge is efficiently integrated at the web’s expanding frontier , yet this very per-skill collapse onto a single high-reward path also accumulates over time into a macroscopic policy collapse and a global loss of solution diversity when training proceeds too far beyond criticality.
-This framework is not just descriptive; it is prescriptive.
-It identifies the maximally frustrated state—the peak of competitive forgetting—as the ideal moment for SFT intervention.
-Our Annealed-RLVR algorithm is born from this insight.
-By applying a timed, corrective SFT shock, Annealed-RLVR mitigates the policy collapse and, as our experiments confirm, unlocks a superior final reasoning capability.
-Ultimately, this work provides a new lens for understanding RLVR: not as black-box optimization, but as a predictable process of structural self-organization, paving the way for more principled methods for engineering the advanced reasoning capabilities of future AI.
-Looking forward, a central long-term goal is the empirical mapping of the complete, microscopic reasoning graph from large-scale foundation models.
-This remains a significant undertaking, requiring substantial computational resources and novel algorithms for semantic chunking and clustering.
-The payoff for this effort, however, would be transformative, producing a powerful new analytical object: a detailed map of an AI’s internal world, composed of the reasoning graph and its emergent concept web.
-While the immediate application of this map would be to empirically test our sparse-web hypothesis, its long-term impact is far broader: charting this internal world is a critical step toward genuinely explainable LLMs and the robust safety of future artificial intelligence.
-7 Related Work
-Our work intersects with three domains: training paradigms for LLM reasoning, the modeling of reasoning as a complex network, and the study of RLVR’s failure modes.
-While building on the standard RLVR paradigm (Guo et al., 2025; Team et al., 2025; Yang et al., 2025), our goal is to provide a theoretical-physics-motivated theory for its perplexing macroscopic dynamics.
-To this end, we engage with the view of reasoning as a complex network; however, our approach differs from prior work that empirically extracts static reasoning graphs post-hoc (Xiang et al., 2025; Cabannes et al., 2024; Minegishi et al., 2025).
-Inspired by the core logic of the RG, we bypass the intractable graph-construction challenge and instead hypothesize a simple organizational principle for the emergent, coarse-grained concept web with $\left<k\right>\approx 2$, enabling us to study the universal behaviros of RLVR.
-This sparse-web perspective provides a unified origin for seemingly distinct failure modes like catastrophic forgetting (Li & Hoiem, 2017; Ding & Wang, 2025) and policy collapse (Yue et al., 2025; Cui et al., 2025), framing the former as the severing of the network’s fragile bridges, and the latter as a consequence of a phase-transition-like learning dynamics at its frontier (Power et al., 2022; Lin et al., 2025; Cai et al., 2025).
-An important line of research has made valuable progress on these failures using algorithmic solutions (Chen et al., 2025).
-Our work contributes a complementary, theory-driven perspective: we leverage the sparse-web model to design Annealed-RLVR, an intervention that targets the problem’s underlying structural root.
-See Appendix B for extended related work.
-8 Reproducibility Statement
-To ensure full reproducibility, we provide a complete experimental blueprint in the Appendix: Sec.
-I details the LLM training pipeline—models, hyperparameters, and SFT dataset construction, while Sec.
-J describes our evaluation protocol.
-Our code repository of CoNet can be available at: https://anonymous.4open.science/r/CoNet-83A4.
-The repository includes a README.md file that details the functionality of each key Python script.
-The training process is described in Appendix D and Appendix H.
-Appendix A The Use of Large Language Models (LLMs)
-This work utilized Large Language Models (LLMs) to assist in the preparation of the manuscript, primarily for improving language clarity and readability.
-For the accompanying CoNet repository, we also employed LLMs to generate Python codes, such as the training algorithm, measuring network properties and visualizing the concept web and its largest connected component.
-All LLM-generated outputs, including text and code, were meticulously reviewed, edited, and verified by the authors to ensure their accuracy and suitability for this work.
-Appendix B Extended Related Work
-System-2 Reasoning in LLMs.
-The dual-process theory, distinguishing between fast, intuitive System-1 thinking and slow, deliberate System-2 reasoning (Kahneman, 2011), has become a valuable framework for analyzing LLM behavior.
-While standard LLMs often exhibit System-1-like responses, significant research has focused on eliciting deliberate System-2 reasoning.
-Prompting strategies such as Chain-of-Thought (CoT) (Wei et al., 2022) and its variants like Self-Consistency (Wang et al., 2022) and Tree of Thoughts (ToT) (Yao et al., 2023) encourage models to generate explicit reasoning steps.
-Other works aim to build this deliberative capacity into the model’s architecture or training process through self-correction and reflection (Shinn et al., 2023; Madaan et al., 2023).
-Our work complements these efforts by shifting the focus from eliciting reasoning to understanding the emergent mechanism of how this capability develops internally during training.
-Complex Network Analysis of Neural Models.
-Applying network science to understand neural networks is a long-standing idea, previously used to study the topology and robustness of smaller models (Simard et al., 2005; Estrada & Hatano, 2008).
-With the advent of Transformers, researchers have begun analyzing attention patterns as graphs (Vig, 2019) to better understand information flow (Abnar & Zuidema, 2020; Chefer et al., 2021).
-More recently, studies have explored the topological properties of neural networks, identifying optimal graph structures (You et al., 2020) and the formation of essential sparse subnetworks (Frankle & Carbin, 2019; Filan et al., 2021).
-Complementing these static analyses, recent work has modeled the training process itself as the evolution of complex networks (Danovski et al., 2024; Rocchetti, 2025).
-A recent study by (Wang et al., 2024; Minegishi et al., 2025; Matsutani et al., 2025) shares our goal of analyzing the network structures that underpin System-2 reasoning, though they do so by empirically extracting a graph from hidden-state representations.
-In contrast, our CoNet framework utilizes a fixed, abstract concept space, which allows us to specifically isolate and analyze the dynamic mechanism of how this reasoning structure emerges during training.
-Appendix C The Reasoning Graph: From Empirical Extraction to a Unified Concept Web
-The idea of a reasoning graph has emerged in the literature as a powerful framework for understanding the chain-of-thought process in LLMs for a given task.
-This graph is not explicitly encoded in the Transformer’s static parameters but is an emergent property of its inference dynamics.
-The standard model for its construction is based on a key observation: an LLM’s generative process is guided by a few decision points of high Shannon entropy, or “forking tokens” (Fig.
-7).
-This allows for a deconstruction of a chain-of-thought: stable, low-entropy text chunks form the graph’s nodes (concepts), while the probabilistic choices at these high-entropy forks form the weighted, directed edges.
+7 reports the $R_{c}$ trajectories at $r\in\{1,2,3,4\}$ for both dist=4=4 and dist=6=6 training conditions.
+All four curves exhibit the same V-shaped dynamics and differ only by an overall vertical shift: the looser the rounding (rd=1=1), the easier the equivalence criterion is to satisfy and the higher $R_{c}$.
+At the strictest setting rd=4=4, the post-training compression ratio is $R_{c}\approx 0.93$ for dist=4=4 and $R_{c}\approx 0.84$ for dist=6=6 (final-epoch values $0.925$ and $0.842$, last-10-epoch means $0.94$ and $0.825$, peaks $0.964$ and $0.872$; data from rc_psv_rd4.json of the two experiments).
+Across the late training window the values fluctuate roughly between $78\%$ and $93\%$.
+The fact that $R_{c}$ remains in this range at the strictest threshold confirms that the compression is a genuine structural feature of the policy rather than an artefact of the rounding precision.
+Appendix B From Empirical Reasoning Graphs to Predictive State Variables
+A growing body of work represents LLM reasoning as a graph of effective states linked by transitions.
+Empirically, reasoning chains are decomposed into low-entropy “concept” chunks separated by high-entropy “forking tokens” at which the model commits to one of several semantically distinct continuations (Fig.
+8).
+Recent constructions take this view further: (Minegishi et al., 2025) clusters hidden-state representations of segmented reasoning steps to build task-specific graphs and reports that more capable models develop graphs of larger diameter; (Wang et al., 2024) analyses LLM reasoning through reasoning-path aggregation during pretraining.
 Figure 8: Illustration of the LLM reasoning process.
-A chain-of-thought is composed of stable, low-entropy text chunks (concepts), connected by high-entropy ’forking tokens’ that represent decision points.
-The LLM reasoning process can be deconstructed into stable “concepts” (e.g., step 3) linked by high-entropy decision points.
-The probabilities at these junctures (e.g., “next” vs.
-“tens”) represent transition weights on an underlying reasoning graph, a view supported by convergent findings on the role of high-entropy tokens.
-This figure is adapted from Ref.
-Cai et al.
-(2025).
-Recent work, notably by (Minegishi et al., 2025), has developed detailed methods to empirically construct these task-specific reasoning graphs.
-Their process involves generating a large corpus of reasoning chains, segmenting them into individual steps, and then computing a vector representation for each step by averaging the hidden states of its tokens at a specific Transformer layer.
-Finally, a clustering algorithm (e.g., K-means) is used on the full set of step vectors, with the resulting cluster centroids being defined as the graph’s nodes.
-By analyzing graphs constructed this way, they found that more advanced models exhibit distinct topological properties, such as larger diameters, which correlate with improved performance.
-However, the cluster centroids derived from K-means offer limited interpretability as stable reasoning concepts, making their correspondence to the model’s internal processes unclear.
-Moreover, the induced graph’s structure is unstable, varying significantly with the choice of model, dataset, segmentation, and representation layer.
-Hence, this graph cannot track evolving representational dynamics during RLVR training process.
-How to construct a principled model of an LLM’s internal reasoning process remains a formidable open problem.
-While analyzing these static, task-specific graphs is useful, it leads to a deeper question: what is the underlying, universal knowledge structure that allows an LLM to generate so many different reasoning graphs across a multitude of tasks?
-It follows that these empirically observed reasoning graphs are best understood as small projections of a much vaster, latent concept web.
-This web represents the entirety of the model’s abstract knowledge, and the ultimate goal is to understand its structure.
-Figure 8: From Intractable LLM Dynamics to a Minimal CoNet Model.
-The reasoning process in an LLM (left) can be viewed as a traversal on a high-dimensional latent graph, where path probabilities are determined by the complex, autoregressive Transformer.
-Directly analyzing this graph’s evolution is intractable.
-CoNet (right) provides a minimal abstraction by replacing the intractable latent graph with a fixed K-regular random graph and simplifying the generation policy to a learnable Markov chain.
-This creates a tractable “computational microscope” to precisely analyze how the transition policy evolves under reinforcement, isolating the core dynamics of reasoning path formation.
-This figure is adapted from Ref.
-Cai et al.
-(2025).
-However, attempting to reverse-engineer this universal concept web from its sparse, task-specific projections presents formidable obstacles.
-The difficulties are multifold.
-First, there is an issue of scale: the true concept web is of astronomical size, and any extracted graph is a small sample.
-Second, the mapping itself is ambiguous: the mapping from a low-level token sequence to a high-level semantic concept is inherently many-to-one.
-Third, this already-ambiguous mapping is itself dynamic, as it can shift during RLVR training.
-These challenges collectively lead to the observable node stability problem in practice, making it nearly impossible with current methods to track the web’s true topological evolution.
-It is precisely this observational challenge that our paper bypasses with a two-pronged, indirect approach.
-First, we form a theoretical hypothesis about the consequences of a specific topological structure (i.e., sparsity).
-Second, we use a minimal model (CoNet) that captures the essential physics of the system to study these consequences in a controlled environment.
-Thus, while the direct construction of the concept web remains an open problem, our work proceeds from a central hypothesis about its fundamental structure.
-This hypothesis—that the web is profoundly sparse with an average degree near two—serves as the theoretical foundation for our paper.
-From it, we first demonstrate how this assumption of sparsity explains the known macroscopic dynamics of RLVR.
-We then use it to deduce previously overlooked microscopic mechanisms, namely frustration-induced forgetting and phase-transition-like learning.
-Finally, armed with these new insights, we explore how this deeper understanding can be leveraged to design a novel, improved RLVR algorithm.
-Appendix D CoNet: A Minimal Model for RLVR dynamics
-To bypass the intractable node stability problem, we leverage the Concept Network Model (CoNet).
-CoNet was originally introduced in (Cai et al., 2025) to study the dynamics of single-task RLVR, but its core design makes it perfectly suited for our investigation.
-By positing a fixed space of concept nodes (a static graph structure) and isolating all learning dynamics to the evolving policy that navigates it, CoNet allows us to sidestep the challenges of direct graph extraction and precisely measure how the reinforcement signal sculpts reasoning paths.
-The model represents the LLM’s concept space as an abstract K-regular random graph, with the reasoning process simplified to a policy-guided Markovian random walk, as shown in Fig.
-8.
-The transition policy, $\pi_{\theta}(j|i)\propto\theta_{ij}$, is governed by a set of learnable parameters $\theta_{ij}$ that represent the strength of the connection from node i to node j.
-This policy is updated via RLVR on successful paths from a question (Q) to an answer (A).
-This design offers a clear cognitive analogy.
-A single-step hop on the graph represents “System 1” thinking—fast, intuitive associations.
-A multi-step traversal, guided by the learned policy, represents “System 2” reasoning—a deliberate, compositional chain of thought.
-It is crucial to note that CoNet is not a model of the Transformer’s parameters themselves; rather, it models the emergent semantic network induced by inference—the structure that governs how questions are connected to answers through chains of concepts.
-CoNet was originally introduced to model the phenomenon of Learning at Criticality (LaC), which was identified in single-task RLVR training (Cai et al., 2025).
-The central finding of LaC is that for a model trained on extremely sparse data (e.g., a single exemplar), generalization performance on unseen problems peaks precisely at a sharp, sigmoidal learning transition, before declining due to overfitting.
-The original work used CoNet to establish that this dynamic is a learning phase transition, providing a physical model for this peak in generalization.
-The central methodological advance of this paper is to deploy this existing model in a richer, multi-task setting with numerous concurrent Q-A pairs.
-This extension is crucial: it transforms CoNet from a model of isolated skill acquisition into an ideal theoretical laboratory for studying skill integration.
-In this regime, the model is compelled by competing reinforcement signals to discover a shared, reusable structure.
-As Fig.
-1 shows, this multi-task CoNet stunningly reproduces the key macroscopic signatures of the 1.5B LLM.
-This correspondence validates our use of CoNet as a minimal model for exploring the network-level mechanisms of frustration and web formation.
-Appendix E CoNet Implementation Details
+A chain-of-thought is composed of stable, low-entropy text chunks (concepts), connected by high-entropy “forking tokens” that represent decision points (e.g., “next” vs.
+“tens”).
+The probabilities at these junctures represent transition weights on an underlying reasoning graph.
+Figure adapted from Ref.
+(Cai et al., 2025).
+Useful as these constructions are, post-hoc graph extraction faces a fundamental node-stability problem.
+K-means cluster centroids over hidden states do not have an intrinsic dynamical meaning: the resulting graph shifts when the representation layer, segmentation granularity, clustering algorithm, or training checkpoint changes.
+The mapping from low-level token sequences to high-level semantic concepts is many-to-one and itself evolves during RLVR training, so an extracted graph cannot reliably track the same nodes across training time.
+The effective Markov representation introduced in the main text resolves this difficulty in a principled way.
+Predictive state variables (PSVs) define nodes as equivalence classes of histories whose future conditional distributions agree to a fixed precision (Eq.
+1 of the main text).
+This definition is layer-agnostic and clustering-free: two histories belong to the same PSV class if and only if they generate indistinguishable futures, an intrinsic property of the model’s policy.
+Finite parametric capacity bounds the number of distinguishable conditional distributions by $\exp(O(d))$, so a great many distinct histories must necessarily collapse onto a small set of effective states.
+The sliding-puzzle experiment in §EFFECTIVE MARKOV REPRESENTATION of the main text directly verifies this compression: a 14K-parameter Transformer trained with GRPO achieves a compression ratio of 78–93% across rounding precisions, confirming that the PSV graph is dynamics-defined and observable rather than hypothesised.
+The CoNet picture used in §THE CONCEPT NETWORK is the abstraction obtained by replacing the implicit, dynamics-defined PSV graph with an explicit, parametrically-controlled random graph (Fig.
+9).
+It is a minimal computational platform on which the same merging+frustration mechanisms can be studied with full microscopic observability, complementing the indirect PSV measurement on the sliding-puzzle Transformer.
+Figure 9: From an implicit PSV graph to a minimal CoNet model.
+The reasoning process in an LLM (left) is a traversal on the implicit PSV equivalence-class graph; transition probabilities are determined by the autoregressive Transformer policy.
+CoNet (right) is the minimal explicit counterpart: a fixed $K$-regular random graph with a learnable Markov-chain policy.
+Both share the structure “random walk on a sparse graph”; CoNet exposes the policy parameters directly so that merging and frustration can be observed.
+Figure adapted from Ref.
+(Cai et al., 2025).
+Appendix C CoNet: Origin and Multi-Task Extension
+CoNet was introduced by (Cai et al., 2025) as a minimal abstraction of LLM reasoning dynamics.
+The model represents the concept space as a $K$-regular random graph, and a reasoning chain as a policy-guided random walk; the transition policy $\pi_{\theta}(j\mid i)\propto\theta_{ij}$ is governed by learnable parameters $\theta_{ij}$ representing the connection strength from node $i$ to node $j$.
+This policy is updated by GRPO on the success or failure of a trajectory from question $Q$ to answer $A$.
+In the original single-task setting, CoNet captured the Learning-at-Criticality (LaC) phenomenon—a sharp, sigmoidal learning transition accompanying generalisation peaks under extremely sparse training data—and identified it with a continuous learning phase transition.
+The methodological contribution of the present work is to deploy CoNet in a multi-task regime, with many concurrent question–answer pairs sharing a single set of transition parameters.
+The shift is conceptually decisive.
+With one task, the policy only needs to find an efficient path; with many tasks sharing parameters, distinct reasoning paths inevitably collide on shared nodes and place mutually exclusive demands on the same outgoing-edge probabilities, which is the microscopic origin of the frustration mechanism analysed in §TWO CORE MECHANISMS of the main text.
+The single-task LaC sigmoid (monotonic length decrease + sigmoidal reward) is recovered as a limiting case in §SK.
+In the cross-scale comparison of the main text (Fig.
+3 of the main text), CoNet is the smallest of three systems—CoNet, a 14K-parameter sliding-puzzle Transformer, and the 1.5-billion-parameter DSR-1.5B—that all display the same two-stage dynamics and V-shaped path-length curve.
+CoNet’s role is therefore not to “bypass” a problem unique to LLMs but to provide an analytically transparent counterpart: every transition probability is logged in full and every connected component of high-probability edges is directly visualisable, enabling the microscopic tests of merging and frustration reported in the main text.
+A single-step hop on CoNet’s graph represents fast, intuitive “System 1” association; a multi-step traversal guided by the learned policy represents deliberate, compositional “System 2” reasoning.
+Importantly, CoNet does not model the Transformer’s parameters themselves: it models the emergent semantic network induced by inference—the structure that connects questions to answers through chains of effective states.
+The hyperparameters of the multi-task CoNet experiments reported here are detailed in §SD.
+Appendix D CoNet Implementation Details
 We configured CoNet with a specific set of structural and learning parameters.
@@
 This microscopic conflict is the origin of the macroscopic “frustration” discussed in the main text, compelling the system to find a shared, globally coherent structure that can arbitrate these conflicts, rather than settling for 128 disconnected “skill islands”.
-The policy was optimized using a variant of the Group Relative Policy Optimization (GRPO) algorithm ((Shao et al., 2024; Liu et al., 2025; Yu et al., 2025)).
+The policy was optimized using a variant of the GRPO algorithm ((Shao et al., 2024; Liu et al., 2025; Yu et al., 2025)).
 For a given Q-A pair, $n_{\text{rollout}}=128$ reasoning paths (indexed by $m$) were sampled, with each path’s exploration capped at a maximum length of 20 steps.
@@
 The robust emergence of these LLM-like signatures in this intermediate regime confirms that our CoNet setup successfully reproduces the essential dynamics of structural self-organization under multi-task reinforcement.
-Appendix F Structural Evolution of the Concept Web in CoNet
+Appendix E Structural Evolution of the Concept Web in CoNet
 This section provides direct, quantitative evidence from the CoNet model for the two-stage learning dynamic—a fast acquisition process followed by a slow integration phase—that underlies our theory.
 Fig.
-9 serves as the central exhibit for this analysis, tracking two key structural properties over the course of training: the total number of disconnected solution clusters (the “skill islands”) and the size of the single largest cluster, namely the concept web.
+10 serves as the central exhibit for this analysis, tracking three key structural properties over the course of training: the total number of disconnected solution clusters (the “skill islands”), the size of the single largest cluster, namely the concept web and the average degree of the concept web.
 Figure 10: The Formation of the Concept Web in CoNet.
-This figure illustrates the evolution of network clusters during training, showing a structural reorganization from isolated skills to a unified conceptual web.
+This figure illustrates the evolution of network clusters and topology during training, showing a structural reorganization from isolated skills to a unified conceptual web.
 In the initial phase, the Cluster Number (orange curve) spikes, indicating the rapid discovery of numerous disconnected “skill islands”.
 The peak marks a critical structural juncture.
-Following this peak, the Cluster Number steadily declines as these islands begin to coalesce.
+Following this peak, the Cluster Number steadily declines as these islands begin to integrate.
 This merging process is confirmed by the corresponding monotonic growth in the Max Cluster Size (blue curve), which represents the formation of a single giant component.
-By the end of training, the network is dominated by this large, interconnected structure.
+Crucially, the Average Degree of the conept web (green curve) rapidly converges to and remains pinned at $\approx 2$ throughout the slow-learning phase.
+This persistent sparsity confirms that the global integration occurs within a strictly constrained, tree-like topology, validating our central hypothesis.
 The first stage of training ( $\sim$ 0-50 steps), corresponding to the fast-learning phase, is characterized by the rapid proliferation of “skill islands”.
 The quantitative evidence for this is the steep, initial rise of the “Cluster Number” (orange curve) in Fig.
-9.
+10.
 Each new cluster represents the discovery of a distinct, high-confidence reasoning path for a specific problem.
@@
 The causal link is direct: the largest component grows precisely by absorbing smaller, independent clusters, which necessarily reduces the total cluster count.
+Moreover, the average degree of the concept web remains pinned at $\approx 2$ throughout the expansion, indicating that the concept web maintains a strictly sparse, tree-like topology.
 This structural evolution culminates in a state where a single, giant component—the functional concept web—dominates the network.
 This process provides a direct, mechanical explanation for the V-shaped response length signature: Stage 1 shortens paths by finding local solutions, while Stage 2 lengthens them by forcing the policy to traverse longer, connective paths between previously distant concepts.
+(a) Step 20
+(b) Step 50
+(c) Step 800
+Figure 11: Direct visualization of inverse-tree freezing in CoNet.
+Network snapshots at the three training steps marked on Fig.
+10, retaining only edges with transition probability $p>0.95$.
+Green = question nodes, red = answer nodes, blue = intermediate nodes; edge colour encodes the transition probability (right colour bar).
+Unlike Fig.
+4 of the main text, which extracts only the largest inverse tree at each step, this figure shows all surviving high-probability edges, exposing the full landscape of connected components throughout training.
+(a) Step 20: multiple short, disjoint subtrees nucleate independently across the graph—the nascent “skill islands”—coinciding with the peak of the cluster-number curve in Fig.
+10 and the onset of the maximally frustrated state.
+(b) Step 50: a dominant component has emerged through the merging of compatible paths at shared nodes, while a handful of small subtrees have yet to be absorbed; this is the regime in which frustration-induced forgetting is most active.
+(c) Step 800: the surviving high-probability edges have coalesced into a single connected inverse tree with average degree $\langle k\rangle\approx 2$, the frozen-in concept web that supports the trained policy.
+These three macroscopic indicators (Fig.
+10) are made microscopically concrete by the network snapshots in Fig.
+11, which visualize the same three training steps directly.
+The Step 20 snapshot makes the “proliferation of skill islands” literal: many short, disjoint paths populate the graph in parallel, exactly as expected when nucleation dominates and merging has not yet set in.
+The Step 50 snapshot shows the structural fingerprint of the maximally frustrated state: a dominant connected component has begun to grow by welding compatible paths at shared nodes, but several subordinate islands persist, sustaining the inter-task competition that drives frustration-induced forgetting.
+By Step 800, the high-probability edge set has frozen into a single, sparse inverse tree—the concept web—whose multi-input, single-output topology and average degree $\langle k\rangle\approx 2$ are precisely the structural ingredients that the main text identifies as responsible for both the V-shaped response-length signature and the bridge-node fragility under SFT.
+Appendix F Geometric Origin of the Rising Branch of Response Length
+The main text argues that the lengthening of correct responses during Stage 2 is a direct geometric consequence of the inverse tree: with average degree $\langle k\rangle\approx 2$ and out-degree close to $1$ at almost every node, the network has no short-range shortcuts, so as the inverse tree extends to cover more question–answer pairs the typical question-to-answer geodesic grows monotonically.
+This appendix provides the corresponding microscopic evidence in CoNet: tracking the largest inverse tree alongside the distribution of correct path lengths during Stage 2 shows that the two evolve together, exactly as the inverse-tree picture predicts.
+(a) Correct-response length distribution
+(b) Step 50
+(c) Step 800
+Figure 12: Inverse-tree growth and the rising branch of response length in CoNet.
+(b, c) Largest connected component at $p>0.95$ at Step 50 and Step 800, drawn with the same colour and marker scheme as Fig.
+11.
+The inverse tree expands substantially between the two steps—absorbing additional question–answer pairs through merging at shared nodes—while keeping its sparse, multi-input/single-output character: average degree stays $\langle k\rangle\approx 2$ (cf.
+Fig.
+10).
+(a) Distribution of correct solution path lengths over the same training window (raw counts, with the probability density shown in the inset).
+As the inverse tree grows, the length distribution shifts decisively to the right; both the mean and the upper tail move outward, while the support remains bounded above by the maximum episode length.
+The rightward shift of the length distribution in Fig.
+12a is not an incidental feature of CoNet’s hyperparameters but the direct macroscopic signature of the geometric constraint imposed by the frozen inverse tree.
+Once Stage 2 begins, new question–answer pairs are no longer mastered by discovering fresh short-cut paths in the underlying random graph; they are mastered by being grafted, through merging, onto the existing inverse tree at one of its branching nodes.
+Each such grafting attaches an additional segment of length set by the graph distance from the new question node to the nearest node already on the tree, and this distance is itself bounded below by the tree’s growing radius.
+As a result, the typical solution path length must increase together with the size of the inverse tree, exactly as observed when the largest component expands from Step 50 (Fig.
+12b) to Step 800 (Fig.
+12c).
+This sparse-tree-geodesic argument also clarifies what the rising branch is not.
+It is not driven by the model adopting longer reasoning styles, nor by a switch to harder problems, nor by stochastic noise on already-mastered paths: in CoNet none of these can vary, since the graph, the problem set, and the policy parameterisation are all fixed by construction.
+The only quantity that changes between Step 50 and Step 800 is the topology of the high-probability sub-network, and yet the response-length histogram shifts decisively.
+The rising branch of the V is therefore a structural readout of inverse-tree growth, in line with the corresponding mechanism stated in the main text.
+Self-correction or reflection-style cycles are fully compatible with this picture: they are sparse, local loops on the same backbone, occasional rather than dense, and do not appreciably change the global average degree from $\langle k\rangle\approx 2$.
 Appendix G Learning by Phase Transition in the Slow-Learning Stage
@@
 The success of this SFT-RLVR cycle is also confirmed by the results in Fig.
-12.
+15.
 The line plot on the right shows the immediate effect of the intervention: best@1 accuracy (a proxy for exploitation) temporarily dips, while best@16 accuracy (a proxy for exploration) surges.
@@
 I.1 Base Model and GRPO Setup
-We used the DeepSeek-R1-Distill-Qwen-1.5B model (Guo et al., 2025), trained on the DeepScaleR-Preview-Dataset (Luo et al., 2025a) with their modified version of the veRL open-source engine (Sheng et al., 2024; ByteDance Seed Team and verl community, ).
-Our Group Relative Policy Optimization (GRPO) hyperparameters followed the first-stage settings, which specified a response length of 8192, as outlined in the DeepScaleR paper (Luo et al., 2025a).
+We used the DeepSeek-R1-Distill-Qwen-1.5B model (Guo et al., 2025), trained on the DeepScaleR-Preview-Dataset (Luo et al., 2025) with their modified version of the veRL open-source engine (Sheng et al., 2025; ByteDance Seed Team and verl community, ).
+Our Group Relative Policy Optimization (GRPO) hyperparameters followed the first-stage settings, which specified a response length of 8192, as outlined in the DeepScaleR paper (Luo et al., 2025).
 I.2 SFT Dataset for Annealed-RLVR
 The SFT dataset for the annealed-RLVR experiment (labeled “Annealed” in Fig.
-6) was constructed to focus on difficult problems.
+6 of the main text) was constructed to focus on difficult problems.
 First, we trained the base model using GRPO for 100 steps to create a checkpoint (CKPT100), which is in the vicinity of the maximally frustrated state.
@@
 The SFT dataset for the catastrophic forgetting experiment (labeled “SFT+RLVR” in Fig.
-4(b)) was designed to specifically target problems forgotten by a more saturated model, using a data generation process that involved multiple checkpoints.
+5b of the main text) was designed to specifically target problems forgotten by a more saturated model, using a data generation process that involved multiple checkpoints.
 The process began with problem filtering: we used a checkpoint from step 600 (CKPT600, deep within the slow-learning stage) to generate 8 rollouts for all training problems.
@@
 This $O(\eta N)$ SFT cost is significantly smaller than the evaluation cost.
-Both of these marginal, one-time costs are much smaller than the total iterative cost of the full GRPO training run (e.g., 1,700 steps).
+Both of these marginal, one-time costs are much smaller than the total iterative cost of the full GRPO training experiment (e.g., 1,700 steps).
 Appendix J LLM Evaluation Details
-We evaluated checkpoints from our RLVR (“GRPO”) and annealed-RLVR (“Annealed”) training runs on two datasets: 512 randomly selected training problems and the Minerva math dataset (Hendrycks et al., 2021; Lewkowycz et al., 2022).
+We evaluated checkpoints from our RLVR (“GRPO”) and annealed-RLVR (“Annealed”) training experiments on two datasets: 512 randomly selected training problems and the Minerva math dataset (Hendrycks et al., 2021; Lewkowycz et al., 2022).
 The evaluation parameters were kept consistent with the training configuration (temperature: $0.6$, top-p: $1.0$, response length: $8192$, etc.).
@@
 Fig.
-6 illustrates the full performance curves, while Table 1 highlights the results from representative checkpoints.
+6 of the main text illustrates the full performance curves, while Table 3 highlights the results from representative checkpoints.
 (a) Training Sample (512 Problems)
@@
 Annealed 63.8(0.5) 64.2(0.5) 67.0(0.5) 66.7(0.5) 67.5(0.5) 67.7(0.5)
+Appendix K Supplementary Evidence: Single-Task Dynamics
+(a) CoNet
+(b) LLM
+Figure 17: Single-Task Dynamics We compare the training dynamics of CoNet (a) and the DeepScaleR-1.5B LLM (b) when trained on a single problem.
+The single-task experiment exhibits: (i) The reward curves (top panels) follow a sigmoidal, phase-transition-like trajectory; (ii) The correct response length (bottom panels) decreases monotonically, converging to the efficient paths.
+Lines in (a) and (b) are smoothed with a 5-step and 4-step moving average respectively.
+We reproduced the single-problem training experiment studied by (Cai et al., 2025) using the DeepScaleR protocol (Luo et al., 2025).
+Specifically, we selected a single problem from the DeepScaleR-Preview-Dataset (Luo et al., 2025) and repeated the single-task training experiment with their modified version of the veRL (Volcano Engine Reinforcement Learning).
+As shown in Fig.
+17, the single-task dynamics exhibit a sigmoidal learning curve and a monotonically decreasing response length, standing in sharp contrast to the two-stage reward dynamic and V-shaped response length observed during multi-task training.
+CoNet faithfully reproduces these distinct behaviors when restricted to the same single-task condition.
+This single-task training more effectively highlights the phase-transition-like learning mechanism described in the main text.
+By restricting the system to a single task, we isolate the fundamental phase-transition dynamics of skill acquisition: the model focuses solely on identifying an optimal trajectory among multiple feasible paths.
+Without the structural requirement to integrate conflicting skills into a shared, sparse backbone, the system avoids the “maximally frustrated state.” Consequently, the response length decreases monotonically as the policy converges to a fixed optimal path, and the reward curve follows a smooth, sigmoidal trajectory characteristic of a continuous phase transition, free from competitive frustration.
+We also note that the reward curve is influenced by finite-size effects imposed by the maximum response length constraint, this phenomenon remains a key subject for further theoretical investigation.
+This capacity to replicate the distinct phenomenologies of both conditions—from the phase transition of isolated learning to the competitive frustration inherent in multi-trajectory integration—provides strong evidence that CoNet serves as a minimal model for investigating the emergent mechanisms of RLVR.